### v1

# Data Contamination Issues in Brain-to-Text Decoding

Abstract
Decoding non-invasive cognitive signals to natural language has long been the goal of building practical brain-computer interfaces (BCIs). Recent major milestones have successfully decoded cognitive signals like functional Magnetic Resonance Imaging (fMRI) and electroencephalogram (EEG) into text under open vocabulary setting. However, how to split the datasets for training, validating, and testing in cognitive signal decoding task still remains controversial. In this paper, we conduct systematic analysis on current dataset splitting methods and find the existence of data contamination largely exaggerates model performance. Specifically, first we find the leakage of test subjects’ cognitive signals corrupts the training of a robust encoder. Second, we prove the leakage of text stimuli causes the auto-regressive decoder to memorize information in test set. The decoder generates highly accurate text not because it truly understands cognitive signals. To eliminate the influence of data contamination and fairly evaluate different models’ generalization ability, we propose a new splitting method for different types of cognitive datasets (e.g. fMRI, EEG). We also test the performance of SOTA Brain-to-Text decoding models under the proposed dataset splitting paradigm as baselines for further research.
11footnotetext: Work was done when the author was an intern in JD.com.22footnotetext: Corresponding author.
1 Introduction
Brain-computer interface (BCI) builds connections between cognitive signals of human brain and external devices (e.g. computer). It has been widely researched in the field of neuroscience and has gained remarkable success like repairing damaged sight or restoring movement of disabled people (Polikov et al., 2005; Hochberg et al., 2012; Bouton et al., 2016). However, it’s still challenging in restoring subjects’ cognitive signals when they read or hear text stimuli back to corresponding natural language, especially when it comes to non-invasive cognitive signals like functional Magnetic Resonance Imaging (fMRI) or electroencephalogram (EEG), which are noisy and of low resolution (Mridha et al., 2021).
Figure 1: General frameworks of decoding cognitive signals to text and possible ways of data contamination.
Previous work (Herff et al., 2015; Anumanchipalli et al., 2019; Zou et al., 2021; Moses et al., 2021; Défossez et al., 2023) mainly focused on word-level cognitive signal decoding with restricted vocabulary size (hundreds of words). Although these models achieve high accuracy, they struggle to generate unseen words, thus leading to poor generalization. Recent methods (Wang and Ji, 2022; Xi et al., 2023) leveraged powerful pre-trained language models like BART (Lewis et al., 2020) in decoding cognitive signals to text and achieved open vocabulary sentence-level generation. These methods typically view the cognitive signal decoding task as machine translation (Sutskever et al., 2014; Bahdanau et al., 2015; Makin et al., 2020) and adopt an encoder-decoder framework, where the encoder is responsible for converting cognitive signals into low-dimensional representations and the decoder learns to map the representations to natural language. Specifically, as shown in Figure1, the encoder usually consists of a spatial and time series feature extractor. It can be trained either in an end-to-end manner with decoder (Figure 1 (a)) or first pre-trained through a signal reconstruction task and then applied in decoder training (Figure 1 (b)). Despite recent success in model design, it still remains controversial in how to split the dataset for training, validating, and testing (Xi et al., 2023). Addressing this issue is urgent and meaningful, because fair evaluations and comparisons of different models are impossible without a widely recognized dataset splitting paradigm.
Datasets are usually formatted as signal-sentence pairs in decoding cognitive signals to text task. In most cases, each sentence belongs to a certain task, so signal-sentence pairs can be further divided into signal-task and task-sentence pairs. Current dataset splitting methods (Wang and Ji, 2022; Xi et al., 2023) can be summarized into five categories: (1) split by subjects (2) split by tasks (3) split by randomly picking signal frames (4) split by randomly picking signal frames under certain task (5) split by randomly picking consecutive signal frames under certain task. However, all these methods suffer from data contamination on encoder side, decoder side, or both. As to encoder, if subjects’ cognitive signals in test set are mixed to training set, encoder will become overfitted and fail to well represent unseen subjects’ cognitive signals. As to decoder, situation gets worse if text stimuli is leaked. Since the decoder generates token by token in an auto-regressive manner, data contamination will cause the decoder to memorize seen paragraphs and probability distribution, which means given the first few tokens the decoder can predict next tokens without encoded cognitive signal representations.
To address the problems, we propose a new dataset splitting method that eradicates data contamination from both encoder and decoder sides. We focus on fMRI and EEG signals in experiments, although the proposed splitting method could be applied to any cognitive signals satisfying the above-mentioned format. In our method of dataset partition, dataset is split according to subject-stimuli pairs with the following rules: (1) Cognitive signals collected from subjects in validation and test set won’t appear in training set, which means the trained encoder can’t access any brain information belonging to the subjects in test set. (2) Text stimuli in validation and test set won’t appear in training set. The decoder learns the mapping between cognitive signal representation and token embedding instead of memorizing seen segment.
Our contributions can be summarized as follows:
  • We investigate current dataset splitting methods and analyze the effect on existing frameworks in decoding cognitive signals to text.
  • We prove data contamination exists in current dataset splitting methods through analysis and experiments, which seriously exaggerates model performance.
  • We propose the first dataset splitting method without data contamination. We also release a fair benchmark to evaluate different models’ generalization performance for further research in this domain.
Figure 2: Different cognitive dataset splitting methods.
2 Related Work
Cognitive Signals Cognitive signals can be classified into three categories: invasive (microelectrode array), partially invasive (ECoG) and non-invasive (EEG, MEG, fMRI) according to how close electrodes get to brain tissue. Due to the high cost and complexity of invasive and partially invasive methods, it’s hard to apply them in building generic and practical BCIs. Here we mainly focus on non-invasive signals EEG and fMRI. EEG signal is electrogram of the spontaneous electrical activity of the brain. Its frequencies usually range from 1 to 30 Hz, divided into several groups like alpha (4-13 Hz), beta (13-30 Hz), delta (0.5-4 Hz), theta (4-7 Hz). EEG is of high temporal resolution and relatively tolerant of subject movement, but its spatial resolution is low and it can’t display active areas of the brain directly. fMRI measures brain activity by detecting changes of blood flow. Blood flow of a specific region increases when this brain area is in use. The spatial resolution of fMRI is measured by the size of voxel, which is a three-dimensional rectangular cuboid ranging from 3mm to 5mm (Vouloumanos et al., 2001; Noppeney and Price, 2004). Unlike EEG which samples brain signals continuously, fMRI samples according to the smallest time period named TR, usually at second level.
Non-invasive Brain Decoding Previous researches (Herff et al., 2015; Anumanchipalli et al., 2019; Zou et al., 2021; Moses et al., 2021; Défossez et al., 2023) on brain-to-text decoding mainly focused on word-level decoding in a restricted vocabulary with hundreds of words (Panachakel and Ramakrishnan, 2021). These models typically apply recurrent neural network or long short-term memory (Hochreiter and Schmidhuber, 1997) network to build mapping between cognitive signals and words in vocabulary. Despite relatively good accuracy, these methods fail to generalize to unseen words. Some progress (Sun et al., 2019) has been made by expanding word-level decoding to sentence-level through encoder-decoder framework, or use less noisy ECoG data (Burle et al., 2015; Anumanchipalli et al., 2019). However, these models struggle to generate accurate and fluent sentences limited by decoder ability. Wang and Ji (2022) introduced the first open vocabulary EEG-to-text decoding model by leveraging the power of pre-trained language models. Xi et al. (2023) improved the model design and proposed a unified framework for decoding both fMRI and EEG signals.
3 Methodology
In this section, we will first introduce the definition of decoding cognitive signals to text task and the general description of dataset format. Then we systematically analyze current dataset splitting methods and point out that all existing methods suffer from two kinds of data contamination: cognitive signal leakage and text stimuli leakage. Finally, a new dataset splitting method is proposed to avoid the above-mentioned data contamination.
3.1 Task Definition
Given the cognitive signal $F_{ij}$ stimulated by $i$-th subject $S_{i}$ hearing or reading certain text $T_{j}$, the goal is to decode $F_{ij}$ back to text $T_{j}^{\prime}$ and make $T_{j}^{\prime}$ as similar as possible to $T_{j}$. The composition of $F_{ij}$ and $T_{j}$ is different as to fMRI and EEG signals. The former samples brain information discretely with a fixed time interval TR, while the latter samples continuously. To fMRI, $L_{window}$ consistent sentence segments $s_{j}$ with fMRI frames $f_{ij}$ are concatenated to form a sample pair $\left\langle F_{ij},T_{j}\right\rangle$. Here $T_{j}=concat(s_{j},\ldots,s_{j+L_{window}-1})$ and $F_{ij}=concat(f_{ij},\ldots,f_{ij+L_{window}-1})$. To EEG, since signals corresponding to a complete sentence are available, we bond sentence (aka text stimuli) $T_{j}$ and EEG signal $F_{ij}$ together to form a sample pair $\left\langle F_{ij},T_{j}\right\rangle$. Under most circumstances, each sentence $T_{j}$ belongs to a certain task $M_{k}$. So the signal-sentence pair $\left\langle F_{ij},T_{j}\right\rangle$ can be further split into $\left\langle F_{ij},M_{k}\right\rangle$ and $\left\langle M_{k},T_{kj}\right\rangle$.
In decoding cognitive signals to text, the ultimate goal of trained BCI models is to generalize to unseen subjects with unseen text stimuli (Huang et al., 2010; Handiru and Prasad, 2016; Gao et al., 2021). If cognitive signal $F_{ij}$ appears in test set $S_{test}$, then any subject $i$’s signals $F_{i}$ should not appear in training set $S_{train}$. Similarly, text stimuli $T_{kj}$ in $S_{test}$ should not appear in $S_{train}$. So the dataset splitting rules can be formally defined as:
$S_{train}=\{\left\langle F_{ij},T_{kj}\right\rangle|F_{i},T_{kj}\notin S_{val},S_{test}\}$ (1)
$S_{val}=\{\left\langle F_{ij},T_{kj}\right\rangle|F_{i},T_{kj}\notin S_{train},S_{test}\}$ (2)
$S_{test}=\{\left\langle F_{ij},T_{kj}\right\rangle|F_{i},T_{kj}\notin S_{train},S_{val}\}$ (3)
3.2 Dataset Splitting Methods
Current dataset splitting methods can be summarized as five categories according to classifying objective $S_{i},M_{k},T_{kj},F_{ij}$. More specifically, five dataset splitting methods are named as (1) split by subjects (2) split by tasks (3) split by randomly picking signal frames (4) split by randomly picking signal frames under certain task (5) split by randomly picking consecutive signal frames under certain task, corresponding to image (a), (b), (c), (d), (e) in Figure 2. Figure 2 is plotted to vividly display the difference between current dataset splitting methods. For simplicity of expression, we choose four subjects with three tasks each containing four, three, four sentences in Figure 2. The line connecting two symbols indicates they are related to a sample. Take path $S_{1},M_{1},T_{11},F_{11}$ for example, it’s a sample where $S_{1}$ listens to stimuli $T_{11}$ belonging to $M_{1}$ and $S_{1}$’s brain signal is recorded as $F_{11}$. Some symbols are connected with several lines. For example, the four lines between $S_{1}$ and $M_{1}$ correspond to $\left\langle M_{1},T_{11}\right\rangle$, $\left\langle M_{1},T_{12}\right\rangle$, $\left\langle M_{1},T_{13}\right\rangle$, $\left\langle M_{1},T_{14}\right\rangle$ counting from left to right. Similarly, the three lines between $M_{1}$ and $T_{11}$ correspond to $\left\langle S_{1},M_{1}\right\rangle$, $\left\langle S_{2},M_{1}\right\rangle$, $\left\langle S_{3},M_{1}\right\rangle$ respectively. The same rules can be extended to other lines and symbols. The green lines and orange lines in Figure 2 stand for training samples and testing samples. The grey dotted line means the sample is abandoned, which will be introduced in our splitting method. As the splitting of validation set is similar to test set, we only consider training set and test set in this section.
We will use method (a), (b), (c), (d), (e) to represent five current dataset splitting methods in the rest of the paper. Method (a) splits the dataset according to subjects, which can be described as
$S_{train}=\{\left\langle F_{ij},T_{kj}\right\rangle|S_{i}\notin S_{val},S_{test}\}$ (4)
for training set. Method (b) splits the dataset according to tasks, which can be described as
$S_{train}=\{\left\langle F_{ij},T_{kj}\right\rangle|M_{k}\notin S_{val},S_{test}\}$ (5)
for training set. Method (c), (d), and (e) all split the dataset according to cognitive signal frames, which can be described as
$S_{train}=\{\left\langle F_{ij},T_{kj}\right\rangle|F_{ij}\notin S_{val},S_{test}\}$ (6)
for training set. However, there are slight differences between these methods. Method (c) views all the cognitive signal frames in dataset as a whole and splits according to the default proportion (e.g. 8:1:1). Method (d) views signal frames under certain task $M_{k}$ as a whole and splits proportionally. Then union all training sets under different tasks to form a complete set for training. Method (e) is similar to method (d). They both first split training, validation, and test set under certain task proportionally and then union them. The difference lies in that method (d) randomly picks signal frames while method (e) consecutively picks signal frames.
Type Method Narratives / ZuCo Average
seed1111 seed2222 seed3333 seed4444
[…]
(b) 6.73 / - 6.32 / - 7.7 / - 17.93 / - 9.67 / -
(c) 12.55 / 12.52 12.52 / 12.55 12.48 / 12.48 12.44 /12.46 12.50 / 12.50
(d) 12.81 / 12.60 12.8 / 12.58 12.78 / 12.56 12.79 / 12.61 12.795 / 12.59
(e) 12.28 / - 12.27 / - 12.26 / - 12.27 / - 12.27 / -
(f) 0.00 / 0.00 0.00 / 0.00 0.00 / 0.00 0.00 / 0.00 0.00 / 0.00
[…]
(b) 0.00 / - 0.00 / - 0.00 / - 0.00 / - 0.00 / -
(c) 100.00 / 13.21 100.00 / 13.06 100.00 / 12.91 100.00 / 13.1 100.00 / 13.07
(d) 99.93 / 0.00 99.81 / 0.00 99.54 / 0.00 99.99 / 0.00 99.82 / 0.00
(e) 9.19 / - 9.31 / - 9.36 / - 9.29 / - 9.29 / -
(f) 0.00 / 0.00 0.00 / 0.00 0.00 / 0.00 0.00 / 0.00 0.00 / 0.00
Table 1: Results of Cognitive Signal Leakage Rate (CSLR) and Text Stimuli Leakage Rate (TSLR)
The goal of decoding cognitive signals to text models is to generalize to unseen subjects with unseen text stimuli, which means both subject’s brain information and received stimuli are new to the trained model. In this sense, we define two kinds of data contamination: cognitive signal leakage and text stimuli leakage. The former will harm encoder training, while the latter will pollute the decoder. The data contamination condition of different methods is reflected in Figure 2. If lines connecting two symbols are of different colours, it means data in test set leaks into training set. Lines between $S_{i}$ and $M_{k}$ indicate cognitive signal leakage condition and lines between $T_{kj}$ and $M_{k}$ indicate text stimuli leakage condition. As shown in the figure, cognitive signal leakage doesn’t exist in method (a), but method (a) suffers from text stimuli leakage. The situation of method (b) is opposite to that of method (a). There’s no text stimuli leakage but cognitive signal leakage. Method (c) and method (d) are similar. They suffer from both cognitive signal leakage and text stimuli leakage. Method (e) is in the same situation as method (b). The composition of samples differs between fMRI signals and EEG signals, so the dataset splitting methods for two cognitive signals are different too. Since fMRI signals need to be sampled continuously with a certain length $L_{window}$, the path of a sample shown in Figure 2 is actually the first part of one fMRI sample, with $L_{window}-1$ continuous part following. In this sense, method (c), (d), and (e) which seem the same for EEG are different for fMRI. Besides, in method (e) there still exists slight text stimuli leakage in the overlap between training samples and test samples for fMRI signal.
3.3 Our Method
To eliminate data contamination from both cognitive signal leakage and text stimuli leakage, we split the dataset by $\left\langle S_{i},T_{kj}\right\rangle$ pairs as shown in (f) of Figure 2. Since EEG and fMRI are different in the composition of dataset, we treat them separately and propose two dataset splitting methods. As to EEG dataset where $F_{ij}$ and $T_{kj}$ form a sample, we consider a bipartite graph $\mathcal{G}_{1}=(\mathcal{U},\mathcal{V},\mathcal{E})$, $\mathcal{U}=\{S_{i}\}_{i=1}^{N}$, $\mathcal{V}=\{T_{j}\}_{j=1}^{M}$. Here $\mathcal{E}$ is connection between node in $\mathcal{U}$ and node in $\mathcal{V}$, indicating $\left\langle S_{i},T_{j}\right\rangle$ pairs in the dataset. $N$ is the total number of subjects and $M$ is the total number of unique text stimuli. We assert $M>N$, so $e=(u,v)\in\mathcal{E}$ exists for every $v\in\mathcal{V}$, as every text stimuli is listened or read by at least one subject. We pick one edge for each node $v\in\mathcal{V}$ and build a new bipartite graph $\mathcal{G}_{2}=(\mathcal{U},\mathcal{V},\mathcal{E}^{\prime})$. Then we split graph $\mathcal{G}_{2}$ by subject $\mathcal{U}$ with the given splitting ratio and form three disjoint graphs $\mathcal{G}_{train},\mathcal{G}_{val},\mathcal{G}_{test}$. In the final step, some edges satisfy no data contamination but are not included in the graph. We add these edges, extending each graph $\mathcal{G}_{train},\mathcal{G}_{val},\mathcal{G}_{test}$ to its maximally scalable state.
As to fMRI dataset, $F_{ij},\ldots,F_{i{j+L_{window}}-1}$ and $T_{kj},\ldots,T_{k{j+L_{window}}-1}$ form a sample. If we follow the same method as EEG, text stimuli leakage will occur in the overlapping part when one sample is assigned to the training set and the other is assigned to the validation or test set. We propose a simple solution that achieves the balance between abandoning as little data as possible and ensuring zero data contamination. Instead of $\left\langle S_{i},T_{kj}\right\rangle$ pair, we consider $\left\langle S_{i},M_{k}\right\rangle$ pair and apply the same algorithm as to EEG dataset. More details and pseudo-code are available in Appendix B.
[…]
$N=1$ $N=2$ $N=3$ $N=4$ $F$ $P$ $R$
UniCoRN 10+1e-3+(a) 49.56 30.49 21.07 15.49 44.83 50.41 40.65
10+1e-3+(b) 26.37 7.50 2.48 0.99 22.28 25.99 19.62
10+1e-3+(c) 50.24 30.83 21.23 15.60 44.68 49.44 41.01
10+1e-3+(d) 49.63 30.29 20.85 15.32 45.06 50.47 41.03
10+1e-3+(e) 28.94 9.39 4.07 1.53 21.68 24.64 19.49
$*{}^{*}$ 20+1e-4+(a) 50.19 34.25 25.98 21.00 46.59 50.36 43.62
30+1e-4+(a) 55.46 40.99 32.85 27.56 52.08 55.02 49.68
20+1e-4+(b) 25.91 8.80 3.84 1.66 20.65 27.74 16.57
30+1e-4+(b) 25.91 8.80 3.84 1.66 20.65 27.74 16.57
20+1e-4+(c) 72.44 60.84 53.35 47.88 70.52 74.10 67.53
30+1e-4+(c) 72.82 61.42 53.95 48.44 71.24 74.41 68.57
20+1e-4+(d) 65.31 51.02 42.54 36.72 62.76 67.09 59.29
30+1e-4+(d) 66.56 53.00 44.75 39.02 63.89 67.51 60.95
20+1e-4+(e) 32.15 12.34 5.57 2.45 24.28 30.43 20.35
30+1e-4+(e) 32.15 12.34 5.57 2.45 24.28 30.43 20.35
Table 2: Generation quality of UniCoRN model for fMRI under different training settings. Here $*{}^{*}$ indicates the encoder of UniCoRN is randomly initialized instead of pre-trained through signal reconstruction task.
4 Experiments and Analysis
We test state-of-the-art cognitive signal to text models on two popular datasets. Comprehensive experiments are conducted to prove the existence of the following phenomenon: (1) Cognitive signals and text stimuli in test set leak into training set in all current dataset splitting methods. (2) The model’s generalization ability, particularly that of the auto-regressive decoder, has been overestimated due to data contamination.
4.1 Datasets
[…]
4.2 Implementation
We follow the same settings of UniCoRN (Xi et al., 2023) and EEG-To-Text (Wang and Ji, 2022), except the datasets are split to the ratio of 8:1:1 for fair comparison. All experiments are conducted on NVIDIA A100-SXM4-40GB GPUs. More details are shown in Appendix A.
[…]
$N=1$ $N=2$ $N=3$ $N=4$ $F$ $P$ $R$
UniCoRN 50+1e-4+(a) 58.09 49.23 43.23 38.43 63.88 61.12 67.50
80+1e-4+(a) 60.88 50.52 43.42 37.84 65.17 61.16 70.72
50+1e-4+(c) 52.30 42.89 36.80 32.17 57.39 51.09 67.29
80+1e-4+(c) 60.78 55.92 53.18 51.10 84.64 63.16 71.50
50+1e-4+(d) 22.90 7.36 2.71 0.95 17.73 19.90 17.33
80+1e-4+(d) 22.90 7.36 2.71 0.95 17.73 19.90 17.33
EEG2Text 50+1e-4+(a) 51.22 33.83 22.99 16.05 46.40 46.85 46.58
80+1e-4+(a) 63.32 52.52 45.19 39.50 65.96 64.74 68.01
50+1e-4+(c) 53.83 38.99 29.57 23.01 53.64 54.19 53.56
80+1e-4+(c) 65.42 57.56 52.56 48.60 73.00 69.99 77.01
50+1e-4+(d) 23.92 8.16 3.21 1.20 20.78 19.96 23.89
80+1e-4+(d) 23.92 8.16 3.21 1.20 20.78 19.96 23.89
Table 3: Generation quality of UniCoRN and EEG2Text model for EEG under different training settings.
4.3 Data Contamination
We have analyzed two kinds of data contamination, cognitive signal leakage and text stimuli leakage in Methodology section. In this part, we will quantify data contamination situation through experiments.
To better illustrate the extent of data contamination across different dataset splitting methods, we design two novel evaluation metrics named Cognitive Signal Leakage Rate (CSLR) and Text Stimuli Leakage Rate (TSLR) for detecting cognitive signal leakage and text stimuli leakage. Note that the situation for validation set is similar as test set, we only consider test set in experiments. CSLR indicates the average percentage of each subject’s cognitive signals in test set appearing in training set, which could be formulated as
$\frac{1}{N}\sum_{i=1}^{N}\min(1,\frac{|\{F_{ij}|F_{ij}\in S_{test}\cap S_{train}\}|}{|\{F_{ij}|F_{ij}\in S_{train}\}|})$ (7)
where $N$ stands for the total number of subjects in test set. ||$\cdot$|| ⋅ | stands for the cardinality of a set. Function $\min(\cdot,\cdot)$ is applied to make sure for each subject the data leakage rate is less than 1.
The definition of TSLR is somewhat different for EEG data and fMRI data. As to EEG data where cognitive signals are sampled continuously, it’s easy to match certain sentence stimuli with corresponding signals. Its TSLR is similar to CSLR, which indicates the average percentage of certain text in test set appearing in training set. TSLR for EEG data can be calculated through
$\frac{1}{M}\sum_{j=1}^{M}\min(1,\frac{|\{T_{ij}|T_{ij}\in S_{test}\cap S_{train}\}|}{|\{T_{ij}|T_{ij}\in S_{train}\}|})$ (8)
where $M$ stands for the total number of unique text periods in test set and $T_{ij}$ stands for $j$-th period of text stimuli received by $i$-th subject, $T_{\cdot j}=T_{j}$, $i$ ranges from 1 to $N$.
Figure 3: Validation loss of different dataset splitting methods in two datasets.
fMRI data is sampled discretely with a deterministic interval, making it hard to acquire signals corresponding to sentences. Previous method instead concatenated continuous fMRI frames of certain length with their corresponding sentence segments as training sample. As a result, we consider the average percentage of the same sentence segments in test set appearing in training set as TSLR of fMRI data. It can be formulated as
$\frac{1}{M}\sum_{j=1}^{M}\tau\frac{|\{T_{ij}|T_{ij}\in S_{test}\cap S_{train}\}|}{|S_{test}|\times L_{window}}$ (9)
where $\tau=0$ if $\{T_{ij}|T_{ij}\in S_{test}\cap S_{train}\}=\emptyset$ else
$\tau=\min(1,\frac{|\{T_{ij}|T_{ij}\in S_{train}\}|}{|\{T_{ij}|T_{ij}\in S_{test}\cap S_{train}\}|}).$ (10)
We test current dataset splitting methods and our method on two cognitive datasets. Considering the influence of randomness in splitting, we select four seeds for experiments. The results are shown in Table 1 and are aligned with theoretical analysis. For fMRI, current methods apart from method (a) suffer from cognitive signal data leakage, while method (a) has serious text stimuli data leakage. Method (b) gets no text stimuli data leakage but has slight cognitive signal data leakage. The situation for EEG is similar to that of fMRI. Apart from our proposed method (f), there is no way to achieve zero cognitive signal leakage and text stimuli leakage at the same time.
[…]
$N=1$ $N=2$ $N=3$ $N=4$ $F$ $P$ $R$
Narratives UniCoRN 22.83 5.69 1.43 0.48 15.55 24.80 19.04
ZuCo UniCoRN 23.32 7.78 3.01 1.09 18.47 20.00 17.92
EEG2Text 24.49 7.49 2.28 0.62 23.98 23.95 25.74
Table 4: A fair benchmark for evaluating cognitive signals to text decoding.
4.4 Overestimated Performance
We discover that cognitive signal leakage and text stimuli leakage affect model training from encoder side and decoder side respectively.
4.4.1 Encoder
As shown in Figure 1, encoder in current models are trained in two different ways: either jointly trained with decoder or solely trained through a reconstruction task. In the former end-to-end training scenario, it’s hard to evaluate encoder performance separately. So we mainly focus on the latter, in which case the encoder is trained through an encoder-decoder framework to reconstruct input cognitive signals. The decoder here doesn’t refer to the decoder for text generation. It’s similar to the structure of the encoder and will be abandoned once the encoder is trained. Although a proper evaluation index of the encoder’s representation ability is unavailable, validation loss is used to measure the influence of data contamination.
We test splitting methods on two datasets and the validation loss of encoder is shown in Figure 3. For fMRI, influenced by leakage of cognitive signals, the validation loss of method (b), (c), (d), (e) keeps dropping even with long training epochs. The encoder is actually overfitting and degrading. For method (a) and (f) without cognitive signal leakage, the validation loss quickly rises after reaching the lowest point with a few epochs, satisfying the basic rule of machine learning. For EEG, we find validation loss keeps dropping for all methods even with very long training epochs, regardless of cognitive signal leakage or not. We think the poor spatial resolution of EEG signal might lead to this phenomenon.
4.4.2 Decoder
All state-of-the-art models choose pre-trained language model BART (Lewis et al., 2020) as decoder. On one hand, the powerful auto-regressive decoder is able to achieve fluent sentence-level open vocabulary text generation. On the other hand, if data contamination occurs, due to the feature of auto-regressive generation, the decoder will generate memorized text given the first few words, which is obviously an act of cheating.
The influence of text stimuli leakage on decoder is detected through BLEU scores (Papineni et al., 2002) and ROUGE-1 scores (Lin, 2004), which measure text similarity between generated text and ground truth. If evaluation indicators keep improving as training epochs increase, we believe part of the test set is leaked into training set and the model is overfitting. For fMRI signal, we test five current dataset splitting methods under different training settings. For fMRI, as shown in Table 2, we test two kinds of UniCoRN models. One is UniCoRN with finely tuned hyper-parameters claimed in the original paper, and the other is $*{}^{*}$ with randomly initialized encoder. Empirically, the former will perform much better than the latter. However, in method (a), (c), (d), due to text stimuli leakage, if we reduce the learning rate and extend training epochs, $*{}^{*}$ performs much better than UniCoRN and its performance keeps rising with longer training epochs. As to method (b) and (e) with no text stimuli leakage, changing training epochs or learning rates makes no obvious difference to model performance. For EEG signal, the conclusion is similar as shown in Table 3. For method (a) and (c) with text stimuli leakage, model performance keeps rising with longer training epochs. For method (d) without text stimuli leakage, both models reach optimal performance after the first few rounds of training epochs.
4.5 Fair Benchmark
We evaluate current state-of-the-art models for decoding cognitive signals to text under our proposed dataset splitting method and release a fair benchmark for further research. UniCoRN is tested for both fMRI and EEG decoding, EEG2Text model is tested for EEG decoding. The results are listed in Table 4. For EEG dataset, UniCoRN achieves higher results in BLEU-2,3,4 while EEG2Text is better in BLEU-1 and ROUGE-1.
5 Conclusion
In this paper, we explore a controversial topic: Due to the complexity of cognitive datasets, no consensus has been reached on how to split the dataset for training, validating, and testing in decoding cognitive signals to text task. We analyze current dataset splitting methods and find serious data contamination largely exaggerates model performance and leads to poor generalization. Sufficient experiments and analysis are conducted to prove the existence of the above phenomena. Moreover, we propose a new dataset splitting method to fix this problem, which can avoid both cognitive signal and text stimuli leakage. State-of-the-art models are reevaluated in this setting and a fair benchmark is released for further research in this domain.
Limitations
The "Narratives" dataset and the ZuCo dataset provide researchers with precise cognitive signal resources stimulated by text or voice. However, in decoding cognitive signals to text task, both subject’s brain signals and text stimuli in validation and test set need to be invisible to training set, which makes splitting these datasets difficult. Our proposed dataset splitting method meets the above requirements at the expense of discarding some data in the dataset. We recommend future datasets in this domain follow these guidelines. The division of the training set, validation set, and test set should be provided when the dataset is released. Besides, we suggest hiring new subjects with unique stimuli for validation set and test set, which is good for testing the generalization ability of models without loss of data. What’s more, we find existing models rely more on the strong auto-regressive decoder to achieve good generation quality. The encoder is of limited use in all SOTA models, which might become a research point in the future.
Ethics Statement
In this paper, we introduce a new dataset splitting method to avoid data contamination for decoding cognitive signals to text task. Experiments are conducted on public accessible cognitive datasets Narratives and ZuCo1.0 with the authorization from their respective maintainers. Both datasets have been de-identified by dataset providers and used for researches only.
Acknowledgements
This research is supported by the National Natural Science Foundation of China (No.62106105), the CCF-Baidu Open Fund (No.CCF-Baidu202307), the CCF-Zhipu AI Large Model Fund (No.CCF-Zhipu202315), the Scientific Research Starting Foundation of Nanjing University of Aeronautics and Astronautics (No.YQR21022), and the High Performance Computing Platform of Nanjing University of Aeronautics and Astronautics. We appreciate JD.com for providing powerful GPUs for experiments.
Appendix A Implementation Details
More details in experiments are supplemented in this section. We perform the same filtering steps to Narratives dataset as UniCoRN paper (Xi et al., 2023) and the same filtering steps to ZuCo1.0 as EEG2Text paper (Wang and Ji, 2022). In signal reconstruction task for encoder of UniCoRN, the batch size of EEG and fMRI data is 512 and 320 respectively. The learning rate is set as 1e-4 and 1e-3 separately as the author claimed in the original paper. In the fair benchmark, for fMRI data, encoder of UniCoRN is trained through 1e-4 learning rate and decaying to 1e-6 finally for 30 training epochs. Decoder is trained through 1e-4 learning rate and decaying to 1e-6 finally for 10 training epochs with 90 batch size. Sample length $L_{window}$ is set as 10 for all experiments related to fMRI. For EEG data, EEG2Text model is trained with 1e-6 learning rate for 80 epochs. UniCoRN model is trained with the same settings as fMRI data.
Figure 4: The process of our proposed dataset splitting method.
Appendix B Our Dataset Splitting Method
In this part, we plot a sketch map of the dataset splitting process and release the pseudo-code of two dataset splitting methods for EEG and fMRI signal. As shown in Figure 4, our proposed dataset splitting method consists of four steps. The blue lines stand for the situation of original dataset. The main difference between two methods lies in the how $\mathcal{G}_{2}$ is generated. We always choose the side with fewer nodes in bipartite graph $\mathcal{G}_{1}$ to perform $\mathcal{G}_{2}$ generation. For example, in Algorithm 1 where we assert $|\mathcal{U}|<|\mathcal{V}|$, the adjacency matrix is initialized as $M\times N$. In Algorithm 2 where $|\mathcal{V}|<|\mathcal{U}|$, the adjacency matrix is initialized as $N\times K$. All hypotheses are based on analysis of cognitive datasets.
One more thing to notice is that in Line 14 of both pseudo-code, the loop indicates extending training set, validation set, and test set respectively. So the names of variable should be alternated in the repeat loop and the displayed part in pseudo-cod is a case example of extending training set. We write it in this way for simplicity of expression.
1 Initialize: Bipartite graph $\mathcal{G}_{1}=(\mathcal{U},\mathcal{V},\mathcal{E})$, $\mathcal{G}_{2}=(\mathcal{U},\mathcal{V},\mathcal{E}^{\prime})$ where $\mathcal{U}=\{S_{i}\}_{i=1}^{N}$ and $\mathcal{V}=\{T_{j}\}_{j=1}^{M}$, Adjacency matrix $A_{1}$ of $\mathcal{G}_{1}$ where $A_{1}[i][j]=1$ if node $i$ and node $j$ is connected else $A_{1}[i][j]=0$, Adjacency matrix $A_{2}$ of $\mathcal{G}_{2}$ where $A_{2}[i][j]=0$, Array $C$ where $len(C)=len(\mathcal{U})$ and $C[i]=0$;
2 for $u\leftarrow U_{1}$ to $U_{N}$ do
3       $C_{copy}\leftarrow C$;
4       for $v\leftarrow A_{1}[u][0]$ to $A_{1}[u][M]$ do
5             if $v=0$ then
6                  $C_{copy}[v.index]\leftarrow\infty$;
7
8      Minimum $=\min(C_{copy})$;
[…]
11Split by subjects $\mathcal{U}$ according to default ratio;
12 $\mathcal{G}_{2}=\mathcal{G}_{train}\cup\mathcal{G}_{val}\cup\mathcal{G}_{test},\,\mathcal{U}_{train}\cap\mathcal{U}_{val}\cap\mathcal{U}_{test}=\emptyset,\,\mathcal{V}_{train}\cap\mathcal{V}_{val}\cap\mathcal{V}_{test}=\emptyset$;
13 repeat // To three sets respectively, below is for training set
14       for $u$ in $\mathcal{U}$  do
15             for $v$ in $\mathcal{V}$ do
16                   if $e=(u,v)\in\mathcal{E}$ and $e=(u,v)\notin\mathcal{E}^{\prime}_{train}$ and $u\notin\mathcal{U}_{val}\cup\mathcal{U}_{test}$ then
17                        $\mathcal{E}_{train}^{\prime}\leftarrow\mathcal{E}_{train}^{\prime}\cup\{e\}$;
18
19
20
21until $\mathcal{G}_{train},\mathcal{G}_{val},\mathcal{G}_{test}$ are all extended;
22return $\mathcal{G}_{train},\mathcal{G}_{val},\mathcal{G}_{test}$;
Algorithm 1 Dataset splitting method for EEG signal
23 Initialize: Bipartite graph $\mathcal{G}_{1}=(\mathcal{U},\mathcal{V},\mathcal{E})$, $\mathcal{G}_{2}=(\mathcal{U},\mathcal{V},\mathcal{E}^{\prime})$ where $\mathcal{U}=\{S_{i}\}_{i=1}^{N}$, $\mathcal{V}=\{M_{k}\}_{k=1}^{K}$, Adjacency matrix $A_{1}$ of $\mathcal{G}_{1}$ where $A_{1}[i][j]=1$ if node $i$ and node $j$ is connected else $A_{1}[i][j]=0$, Adjacency matrix $A_{2}$ of $\mathcal{G}_{2}$ where $A_{2}[i][j]=0$, Array $C$ where $len(C)=len(\mathcal{V})$ and $C[i]=0$;
24 for $v\leftarrow V_{1}$ to $V_{K}$ do
25       $C_{copy}\leftarrow C$;
26       for $u\leftarrow A_{1}[v][0]$ to $A_{1}[v][K]$ do
27             if $u=0$ then
28                  $C_{copy}[u.index]\leftarrow\infty$;
29
30      Minimum $=\min(C_{copy})$;
[…]
33Split by tasks $\mathcal{V}$ according to default ratio;
34 $\mathcal{G}_{2}=\mathcal{G}_{train}\cup\mathcal{G}_{val}\cup\mathcal{G}_{test},\,\mathcal{U}_{train}\cap\mathcal{U}_{val}\cap\mathcal{U}_{test}=\emptyset,\,\mathcal{V}_{train}\cap\mathcal{V}_{val}\cap\mathcal{V}_{test}=\emptyset$;
35 repeat // To three sets respectively, below is for training set
36       for $v$ in $\mathcal{V}$  do
37             for $u$ in $\mathcal{U}$ do
38                   if $e=(u,v)\in\mathcal{E}$ and $e=(u,v)\notin\mathcal{E}^{\prime}_{train}$ and $v\notin\mathcal{V}_{val}\cup\mathcal{V}_{test}$ then
39                        $\mathcal{E}_{train}^{\prime}\leftarrow\mathcal{E}_{train}^{\prime}\cup\{e\}$;
40
41
42
43until $\mathcal{G}_{train},\mathcal{G}_{val},\mathcal{G}_{test}$ are all extended;
44return $\mathcal{G}_{train},\mathcal{G}_{val},\mathcal{G}_{test}$;
Algorithm 2 Dataset splitting method for fMRI signal

### v2

@@
 Recent major milestones have successfully decoded cognitive signals like functional Magnetic Resonance Imaging (fMRI) and electroencephalogram (EEG) into text under open vocabulary setting.
-However, how to split the datasets for training, validating, and testing in cognitive signal decoding task still remains controversial.
-In this paper, we conduct systematic analysis on current dataset splitting methods and find the existence of data contamination largely exaggerates model performance.
+However, how to split the datasets for training, validating, and testing in brain-to-text decoding still remains controversial.
+In this paper, we conduct systematic analysis on current dataset splitting methods and find data contamination largely exaggerates model performance.
 Specifically, first we find the leakage of test subjects’ cognitive signals corrupts the training of a robust encoder.
-Second, we prove the leakage of text stimuli causes the auto-regressive decoder to memorize information in test set.
-The decoder generates highly accurate text not because it truly understands cognitive signals.
-To eliminate the influence of data contamination and fairly evaluate different models’ generalization ability, we propose a new splitting method for different types of cognitive datasets (e.g.
+Second, we prove the leakage of text stimuli causes the auto-regressive decoder to memorize seen information in test set.
+To eliminate the influence of data contamination and fairly evaluate different models’ generalization ability, we propose a new splitting method for different types of cognitive dataset (e.g.
 fMRI, EEG).
@@
 1 Introduction
-Brain-computer interface (BCI) builds connections between cognitive signals of human brain and external devices (e.g.
+Brain-computer interface (BCI) builds connections between human brain and external devices (e.g.
 computer).
 It has been widely researched in the field of neuroscience and has gained remarkable success like repairing damaged sight or restoring movement of disabled people (Polikov et al., 2005; Hochberg et al., 2012; Bouton et al., 2016).
-However, it’s still challenging in restoring subjects’ cognitive signals when they read or hear text stimuli back to corresponding natural language, especially when it comes to non-invasive cognitive signals like functional Magnetic Resonance Imaging (fMRI) or electroencephalogram (EEG), which are noisy and of low resolution (Mridha et al., 2021).
-Figure 1: General frameworks of decoding cognitive signals to text and possible ways of data contamination.
-Previous work (Herff et al., 2015; Anumanchipalli et al., 2019; Zou et al., 2021; Moses et al., 2021; Défossez et al., 2023) mainly focused on word-level cognitive signal decoding with restricted vocabulary size (hundreds of words).
-Although these models achieve high accuracy, they struggle to generate unseen words, thus leading to poor generalization.
-Recent methods (Wang and Ji, 2022; Xi et al., 2023) leveraged powerful pre-trained language models like BART (Lewis et al., 2020) in decoding cognitive signals to text and achieved open vocabulary sentence-level generation.
-These methods typically view the cognitive signal decoding task as machine translation (Sutskever et al., 2014; Bahdanau et al., 2015; Makin et al., 2020) and adopt an encoder-decoder framework, where the encoder is responsible for converting cognitive signals into low-dimensional representations and the decoder learns to map the representations to natural language.
+However, it’s still challenging in decoding cognitive signals to corresponding natural language when subjects read or hear text stimuli, especially for non-invasive cognitive signals like functional Magnetic Resonance Imaging (fMRI) or electroencephalogram (EEG) that are noisy and of low resolution (Mridha et al., 2021).
+Figure 1: General frameworks of brain-to-text decoding and possible situations of data contamination.
+Recent methods (Makin et al., 2020; Wang and Ji, 2022; Xi et al., 2023; Tang et al., 2023) typically viewed brain-to-text decoding as machine translation (Sutskever et al., 2014; Bahdanau et al., 2015) and adopted an encoder-decoder framework, where the encoder is responsible for converting cognitive signals into low-dimensional representations and the decoder learns to map the representations to natural language.
 Specifically, as shown in Figure1, the encoder usually consists of a spatial and time series feature extractor.
-It can be trained either in an end-to-end manner with decoder (Figure 1 (a)) or first pre-trained through a signal reconstruction task and then applied in decoder training (Figure 1 (b)).
+It can be trained either in an end-to-end manner with decoder (Figure 1 a) or first pre-trained through a signal reconstruction task and then applied in decoder training (Figure 1 b).
 Despite recent success in model design, it still remains controversial in how to split the dataset for training, validating, and testing (Xi et al., 2023).
-Addressing this issue is urgent and meaningful, because fair evaluations and comparisons of different models are impossible without a widely recognized dataset splitting paradigm.
-Datasets are usually formatted as signal-sentence pairs in decoding cognitive signals to text task.
-In most cases, each sentence belongs to a certain task, so signal-sentence pairs can be further divided into signal-task and task-sentence pairs.
-Current dataset splitting methods (Wang and Ji, 2022; Xi et al., 2023) can be summarized into five categories: (1) split by subjects (2) split by tasks (3) split by randomly picking signal frames (4) split by randomly picking signal frames under certain task (5) split by randomly picking consecutive signal frames under certain task.
+Addressing this issue is urgent and meaningful, as fair evaluation of models is impossible without a widely recognized dataset splitting paradigm.
+A cognitive dataset is usually formatted in signal-sentence pair.
+In most cases for brain-to-text decoding task, each sentence belongs to a certain task, so signal-sentence pair can be further divided into signal-task and task-sentence pair.
+Current dataset splitting methods (Wang and Ji, 2022; Xi et al., 2023) can be summarized into five categories: (1) split by subjects, (2) split by tasks, (3) split by randomly picking signal frames, (4) split by randomly picking signal frames under certain task, (5) split by randomly picking consecutive signal frames under certain task.
 However, all these methods suffer from data contamination on encoder side, decoder side, or both.
-As to encoder, if subjects’ cognitive signals in test set are mixed to training set, encoder will become overfitted and fail to well represent unseen subjects’ cognitive signals.
-As to decoder, situation gets worse if text stimuli is leaked.
-Since the decoder generates token by token in an auto-regressive manner, data contamination will cause the decoder to memorize seen paragraphs and probability distribution, which means given the first few tokens the decoder can predict next tokens without encoded cognitive signal representations.
-To address the problems, we propose a new dataset splitting method that eradicates data contamination from both encoder and decoder sides.
-We focus on fMRI and EEG signals in experiments, although the proposed splitting method could be applied to any cognitive signals satisfying the above-mentioned format.
-In our method of dataset partition, dataset is split according to subject-stimuli pairs with the following rules: (1) Cognitive signals collected from subjects in validation and test set won’t appear in training set, which means the trained encoder can’t access any brain information belonging to the subjects in test set.
-(2) Text stimuli in validation and test set won’t appear in training set.
-The decoder learns the mapping between cognitive signal representation and token embedding instead of memorizing seen segment.
+As shown in Figure 1, to encoder if subjects’ cognitive signals in test set are mixed into training set, encoder will become overfitted and fail to well represent unseen subjects’ cognitive signals.
+As to decoder, situation gets worse if text stimuli are leaked.
+Since the decoder generates token by token in an auto-regressive manner, data contamination will cause the decoder to memorize seen paragraphs and probability distribution, which means given the first few tokens the decoder is able to predict next token regardless of encoded cognitive signal representations.
+To address the above-mentioned problems, we propose a new dataset splitting method that eradicates data contamination from both encoder and decoder sides.
+We focus on fMRI and EEG signals in experiments, although the proposed splitting method could be applied to any cognitive signals satisfying the given format.
+In our method, the dataset is split according to subject-stimuli pairs with the following rules: (1) Cognitive signals collected from subjects in validation set and test set won’t appear in training set, which means the trained encoder can’t get access to any brain information belonging to subjects in test set.
+(2) Text stimuli in validation set and test set won’t appear in training set.
+The decoder learns the mapping between cognitive signal representation and token embedding instead of memorizing seen text.
 Our contributions can be summarized as follows:
-  • We investigate current dataset splitting methods and analyze the effect on existing frameworks in decoding cognitive signals to text.
-  • We prove data contamination exists in current dataset splitting methods through analysis and experiments, which seriously exaggerates model performance.
-  • We propose the first dataset splitting method without data contamination.
+  • We investigate current dataset splitting methods and analyze their influence on popular frameworks in brain-to-text decoding.
+  • We prove the existence of data contamination in current dataset splitting methods through analysis and experiments, which seriously exaggerates model performance.
+  • We propose the first splitting method without data contamination on public cognitive datasets.
 We also release a fair benchmark to evaluate different models’ generalization performance for further research in this domain.
@@
 2 Related Work
-Cognitive Signals Cognitive signals can be classified into three categories: invasive (microelectrode array), partially invasive (ECoG) and non-invasive (EEG, MEG, fMRI) according to how close electrodes get to brain tissue.
+Cognitive Signal Cognitive signals can be classified into three categories: invasive, partially invasive, and non-invasive according to how close electrodes get to brain tissue.
 Due to the high cost and complexity of invasive and partially invasive methods, it’s hard to apply them in building generic and practical BCIs.
-Here we mainly focus on non-invasive signals EEG and fMRI.
+In this paper, we mainly focus on non-invasive signals EEG and fMRI.
 EEG signal is electrogram of the spontaneous electrical activity of the brain.
@@
 Unlike EEG which samples brain signals continuously, fMRI samples according to the smallest time period named TR, usually at second level.
-Non-invasive Brain Decoding Previous researches (Herff et al., 2015; Anumanchipalli et al., 2019; Zou et al., 2021; Moses et al., 2021; Défossez et al., 2023) on brain-to-text decoding mainly focused on word-level decoding in a restricted vocabulary with hundreds of words (Panachakel and Ramakrishnan, 2021).
+Brain-to-text Decoding Previous research on brain-to-text decoding (Herff et al., 2015; Anumanchipalli et al., 2019; Zou et al., 2021; Moses et al., 2021; Défossez et al., 2023) mainly focused on word-level decoding in a restricted vocabulary with hundreds of words (Panachakel and Ramakrishnan, 2021).
 These models typically apply recurrent neural network or long short-term memory (Hochreiter and Schmidhuber, 1997) network to build mapping between cognitive signals and words in vocabulary.
@@
 3 Methodology
-In this section, we will first introduce the definition of decoding cognitive signals to text task and the general description of dataset format.
+In this section, we will first introduce the definition of decoding cognitive signals to text and the general description of dataset format.
 Then we systematically analyze current dataset splitting methods and point out that all existing methods suffer from two kinds of data contamination: cognitive signal leakage and text stimuli leakage.
-Finally, a new dataset splitting method is proposed to avoid the above-mentioned data contamination.
+Finally, a new dataset splitting method is proposed to avoid the above-mentioned two kinds of data contamination.
 3.1 Task Definition
-Given the cognitive signal $F_{ij}$ stimulated by $i$-th subject $S_{i}$ hearing or reading certain text $T_{j}$, the goal is to decode $F_{ij}$ back to text $T_{j}^{\prime}$ and make $T_{j}^{\prime}$ as similar as possible to $T_{j}$.
-The composition of $F_{ij}$ and $T_{j}$ is different as to fMRI and EEG signals.
+Given the cognitive signal $F_{ij}$ stimulated by $i$-th subject $S_{i}$ hearing or reading certain text $T_{j}$, brain-to-text decoding aims to decode $F_{ij}$ back to text $T_{j}^{\prime}$ and make $T_{j}^{\prime}$ as similar as possible to $T_{j}$.
+The composition of $F_{ij}$ and $T_{j}$ is different as to fMRI and EEG.
 The former samples brain information discretely with a fixed time interval TR, while the latter samples continuously.
-To fMRI, $L_{window}$ consistent sentence segments $s_{j}$ with fMRI frames $f_{ij}$ are concatenated to form a sample pair $\left\langle F_{ij},T_{j}\right\rangle$.
-Here $T_{j}=concat(s_{j},\ldots,s_{j+L_{window}-1})$ and $F_{ij}=concat(f_{ij},\ldots,f_{ij+L_{window}-1})$.
+To fMRI, $L$ consistent sentence segments $s_{j}$ with corresponding fMRI frames $f_{ij}$ are concatenated to form a sample pair $\left\langle F_{ij},T_{j}\right\rangle$.
+Here $T_{j}=concat(s_{j},s_{j+1},\ldots,s_{j+L-1})$ and $F_{ij}=concat(f_{ij},f_{i,j+1},\ldots,f_{i,j+L-1})$.
 To EEG, since signals corresponding to a complete sentence are available, we bond sentence (aka text stimuli) $T_{j}$ and EEG signal $F_{ij}$ together to form a sample pair $\left\langle F_{ij},T_{j}\right\rangle$.
-Under most circumstances, each sentence $T_{j}$ belongs to a certain task $M_{k}$.
+Under most circumstances, each sentence $T_{j}$ belongs to one certain task $M_{k}$.
 So the signal-sentence pair $\left\langle F_{ij},T_{j}\right\rangle$ can be further split into $\left\langle F_{ij},M_{k}\right\rangle$ and $\left\langle M_{k},T_{kj}\right\rangle$.
-In decoding cognitive signals to text, the ultimate goal of trained BCI models is to generalize to unseen subjects with unseen text stimuli (Huang et al., 2010; Handiru and Prasad, 2016; Gao et al., 2021).
-If cognitive signal $F_{ij}$ appears in test set $S_{test}$, then any subject $i$’s signals $F_{i}$ should not appear in training set $S_{train}$.
+Figure 3: The process of our proposed dataset splitting method.
+In brain-to-text decoding, the ultimate goal of trained BCI models is to generalize to unseen subjects with unseen text stimuli (Huang et al., 2010; Handiru and Prasad, 2016; Gao et al., 2021).
+As a result, if cognitive signal $F_{ij}$ appears in test set $S_{test}$, then any subject $i$’s signal $F_{i}$ should not appear in training set $S_{train}$.
 Similarly, text stimuli $T_{kj}$ in $S_{test}$ should not appear in $S_{train}$.
@@
 3.2 Dataset Splitting Methods
-Current dataset splitting methods can be summarized as five categories according to classifying objective $S_{i},M_{k},T_{kj},F_{ij}$.
-More specifically, five dataset splitting methods are named as (1) split by subjects (2) split by tasks (3) split by randomly picking signal frames (4) split by randomly picking signal frames under certain task (5) split by randomly picking consecutive signal frames under certain task, corresponding to image (a), (b), (c), (d), (e) in Figure 2.
-Figure 2 is plotted to vividly display the difference between current dataset splitting methods.
-For simplicity of expression, we choose four subjects with three tasks each containing four, three, four sentences in Figure 2.
-The line connecting two symbols indicates they are related to a sample.
-Take path $S_{1},M_{1},T_{11},F_{11}$ for example, it’s a sample where $S_{1}$ listens to stimuli $T_{11}$ belonging to $M_{1}$ and $S_{1}$’s brain signal is recorded as $F_{11}$.
+Current dataset splitting methods can be summarized as five categories according to classifying objectives $S_{i},M_{k},T_{kj},F_{ij}$.
+More specifically, the five dataset splitting methods are characterised as (1) split by subjects, (2) split by tasks, (3) split by randomly picking signal frames, (4) split by randomly picking signal frames under certain task, (5) split by randomly picking consecutive signal frames under certain task, corresponding to image (a), (b), (c), (d), (e) in Figure 2.
+Figure 2 vividly displays the differences between current dataset splitting methods.
+For simplicity of expression, we choose four subjects with three tasks each containing four, three, four sentences.
+The line connecting two symbols indicates they are related to one sample.
+Take path $S_{1},M_{1},T_{11},F_{11}$ for example, it’s one sample where subject $S_{1}$ listens to text stimuli $T_{11}$ belonging to task $M_{1}$ and $S_{1}$’s corresponding brain signal is recorded as $F_{11}$.
 Some symbols are connected with several lines.
@@
 The same rules can be extended to other lines and symbols.
-The green lines and orange lines in Figure 2 stand for training samples and testing samples.
-The grey dotted line means the sample is abandoned, which will be introduced in our splitting method.
+The green lines and orange lines stand for training samples and testing samples.
+The grey dotted line means the sample is abandoned, which will be introduced in our dataset splitting method.
 As the splitting of validation set is similar to test set, we only consider training set and test set in this section.
@@
 $S_{train}=\{\left\langle F_{ij},T_{kj}\right\rangle|M_{k}\notin S_{val},S_{test}\}$ (5)
-for training set. Method (c), (d), and (e) all split the dataset according to cognitive signal frames, which can be described as
-$S_{train}=\{\left\langle F_{ij},T_{kj}\right\rangle|F_{ij}\notin S_{val},S_{test}\}$ (6)
-for training set.
+for training set. Method (c), (d), and (e) all split the dataset according to cognitive signal frames
+$S_{train}=\{\left\langle F_{ij},T_{kj}\right\rangle|F_{ij}\notin S_{val},S_{test}\}.$ (6)
 However, there are slight differences between these methods.
@@
 8:1:1).
-Method (d) views signal frames under certain task $M_{k}$ as a whole and splits proportionally.
-Then union all training sets under different tasks to form a complete set for training.
+Method (d) views signal frames under certain task $M_{k}$ as a whole and splits proportionally, and then union all training sets under different tasks to form a complete set for training.
 Method (e) is similar to method (d).
 They both first split training, validation, and test set under certain task proportionally and then union them.
-The difference lies in that method (d) randomly picks signal frames while method (e) consecutively picks signal frames.
+The difference lies in that method (d) randomly picks signal frames while method (e) picks consecutive signal frames.
 Type Method Narratives / ZuCo Average
@@
 (f) 0.00 / 0.00 0.00 / 0.00 0.00 / 0.00 0.00 / 0.00 0.00 / 0.00
-Table 1: Results of Cognitive Signal Leakage Rate (CSLR) and Text Stimuli Leakage Rate (TSLR)
-The goal of decoding cognitive signals to text models is to generalize to unseen subjects with unseen text stimuli, which means both subject’s brain information and received stimuli are new to the trained model.
+Table 1: Results of Cognitive Signal Leakage Rate (CSLR) and Text Stimuli Leakage Rate (TSLR).
+The goal of brain-to-text decoding models is to generalize to unseen subjects with unseen text stimuli, which means both subject’s brain information and received text stimuli are new to the trained model.
 In this sense, we define two kinds of data contamination: cognitive signal leakage and text stimuli leakage.
-The former will harm encoder training, while the latter will pollute the decoder.
-The data contamination condition of different methods is reflected in Figure 2.
-If lines connecting two symbols are of different colours, it means data in test set leaks into training set.
-Lines between $S_{i}$ and $M_{k}$ indicate cognitive signal leakage condition and lines between $T_{kj}$ and $M_{k}$ indicate text stimuli leakage condition.
-As shown in the figure, cognitive signal leakage doesn’t exist in method (a), but method (a) suffers from text stimuli leakage.
-The situation of method (b) is opposite to that of method (a).
-There’s no text stimuli leakage but cognitive signal leakage.
+The data contamination situation of different methods is reflected in Figure 2.
+If lines associated with $S_{i}$ or $T_{kj}$ are of different colours, data in test set leaks into training set.
+Lines between $S_{i}$ and $M_{k}$ indicate cognitive signal leakage situation and lines between $T_{kj}$ and $M_{k}$ indicate text stimuli leakage situation.
+Remind the composition of samples differs as to fMRI signal and EEG signal, so the dataset splitting methods are different for two cognitive signals too.
+Since fMRI signals need to be sampled continuously with a certain length $L$, the path of a sample shown in Figure 2 is actually the first part of one fMRI sample, with $L-1$ continuous part following.
+In this sense, for EEG cognitive signal leakage doesn’t exist in method (a), but method (a) suffers from text stimuli leakage.
+The situation of method (b) is opposite to that of method (a), where there’s no text stimuli leakage but cognitive signal leakage.
 Method (c) and method (d) are similar.
@@
 Method (e) is in the same situation as method (b).
-The composition of samples differs between fMRI signals and EEG signals, so the dataset splitting methods for two cognitive signals are different too.
-Since fMRI signals need to be sampled continuously with a certain length $L_{window}$, the path of a sample shown in Figure 2 is actually the first part of one fMRI sample, with $L_{window}-1$ continuous part following.
-In this sense, method (c), (d), and (e) which seem the same for EEG are different for fMRI.
-Besides, in method (e) there still exists slight text stimuli leakage in the overlap between training samples and test samples for fMRI signal.
+For fMRI, method (c), (d), and (e) which seem the same for EEG are actually different splitting ways.
+The situation of data contamination for different methods is similar to EEG, except for method (e) there still exists slight text stimuli leakage in the overlap between training samples and test samples.
 3.3 Our Method
-To eliminate data contamination from both cognitive signal leakage and text stimuli leakage, we split the dataset by $\left\langle S_{i},T_{kj}\right\rangle$ pairs as shown in (f) of Figure 2.
+To eliminate data contamination from both cognitive signal leakage and text stimuli leakage, we split the dataset by $\left\langle S_{i},T_{j}\right\rangle$ pairs as shown in (f) of Figure 2.
 Since EEG and fMRI are different in the composition of dataset, we treat them separately and propose two dataset splitting methods.
-As to EEG dataset where $F_{ij}$ and $T_{kj}$ form a sample, we consider a bipartite graph $\mathcal{G}_{1}=(\mathcal{U},\mathcal{V},\mathcal{E})$, $\mathcal{U}=\{S_{i}\}_{i=1}^{N}$, $\mathcal{V}=\{T_{j}\}_{j=1}^{M}$.
-Here $\mathcal{E}$ is connection between node in $\mathcal{U}$ and node in $\mathcal{V}$, indicating $\left\langle S_{i},T_{j}\right\rangle$ pairs in the dataset.
+As to EEG dataset where $F_{ij}$ and $T_{j}$ form a sample, we consider a bipartite graph $\mathcal{G}_{1}=(\mathcal{U},\mathcal{V},\mathcal{E})$ where $\mathcal{U}=\{S_{i}\}_{i=1}^{N}$, $\mathcal{V}=\{T_{j}\}_{j=1}^{M}$.
+$\mathcal{E}$ is the edge between node in $\mathcal{U}$ and node in $\mathcal{V}$, indicating $\left\langle S_{i},T_{j}\right\rangle$ pair in the dataset.
 $N$ is the total number of subjects and $M$ is the total number of unique text stimuli.
-We assert $M>N$, so $e=(u,v)\in\mathcal{E}$ exists for every $v\in\mathcal{V}$, as every text stimuli is listened or read by at least one subject.
-We pick one edge for each node $v\in\mathcal{V}$ and build a new bipartite graph $\mathcal{G}_{2}=(\mathcal{U},\mathcal{V},\mathcal{E}^{\prime})$.
+We assert $M>N$, so $e=(u,v)\in\mathcal{E}$ exists for every $v\in\mathcal{V}$, as each text stimuli is listened or read by at least one subject.
+As shown in step 2 of Figure 3, first we pick one edge for each node $v\in\mathcal{V}$ and build a new bipartite graph $\mathcal{G}_{2}=(\mathcal{U},\mathcal{V},\mathcal{E}^{\prime})$.
 Then we split graph $\mathcal{G}_{2}$ by subject $\mathcal{U}$ with the given splitting ratio and form three disjoint graphs $\mathcal{G}_{train},\mathcal{G}_{val},\mathcal{G}_{test}$.
-In the final step, some edges satisfy no data contamination but are not included in the graph.
-We add these edges, extending each graph $\mathcal{G}_{train},\mathcal{G}_{val},\mathcal{G}_{test}$ to its maximally scalable state.
-As to fMRI dataset, $F_{ij},\ldots,F_{i{j+L_{window}}-1}$ and $T_{kj},\ldots,T_{k{j+L_{window}}-1}$ form a sample.
-If we follow the same method as EEG, text stimuli leakage will occur in the overlapping part when one sample is assigned to the training set and the other is assigned to the validation or test set.
+In step 4, some edges satisfying zero data contamination condition are not included in the graph.
+We add these edges to corresponding graphs, extending each graph $\mathcal{G}_{train},\mathcal{G}_{val},\mathcal{G}_{test}$ to its maximally scalable state and finishing the dataset splitting process.
+$F_{ij}=concat(f_{ij},f_{i,j+1},\ldots,f_{i,j+L-1})$ and $T_{j}=concat(s_{j},s_{j+1},\ldots,s_{j+L-1})$ form a sample pair in fMRI dataset.
+If we follow the same process as EEG, text stimuli leakage will occur in the overlapping part of two samples, when one sample is assigned to training set and the other is assigned to validation or test set.
 We propose a simple solution that achieves the balance between abandoning as little data as possible and ensuring zero data contamination.
-Instead of $\left\langle S_{i},T_{kj}\right\rangle$ pair, we consider $\left\langle S_{i},M_{k}\right\rangle$ pair and apply the same algorithm as to EEG dataset.
+Instead of $\left\langle S_{i},T_{j}\right\rangle$ pair, we consider $\left\langle S_{i},M_{k}\right\rangle$ pair and apply the above-mentioned algorithm.
 More details and pseudo-code are available in Appendix B.
@@
 4 Experiments and Analysis
-We test state-of-the-art cognitive signal to text models on two popular datasets.
-Comprehensive experiments are conducted to prove the existence of the following phenomenon: (1) Cognitive signals and text stimuli in test set leak into training set in all current dataset splitting methods.
+We test state-of-the-art brain-to-text decoding models on two popular cognitive datasets.
+Comprehensive experiments are conducted to prove the existence of the following phenomena: (1) Cognitive signals and text stimuli in test set leak into training set in all current dataset splitting methods.
 (2) The model’s generalization ability, particularly that of the auto-regressive decoder, has been overestimated due to data contamination.
+Because the number of tasks in EEG dataset is too small and method (e) makes no difference to EEG as method (d), we only consider method (a), (c), (d).
 4.1 Datasets
@@
 4.2 Implementation
-We follow the same settings of UniCoRN (Xi et al., 2023) and EEG-To-Text (Wang and Ji, 2022), except the datasets are split to the ratio of 8:1:1 for fair comparison.
+We follow the same settings of UniCoRN (Xi et al., 2023) and EEG2Text (Wang and Ji, 2022), except the datasets are split to the ratio of 8:1:1 for fair comparison.
 All experiments are conducted on NVIDIA A100-SXM4-40GB GPUs.
@@
 Function $\min(\cdot,\cdot)$ is applied to make sure for each subject the data leakage rate is less than 1.
-The definition of TSLR is somewhat different for EEG data and fMRI data.
-As to EEG data where cognitive signals are sampled continuously, it’s easy to match certain sentence stimuli with corresponding signals.
+The definition of TSLR is somewhat different for EEG signal and fMRI signal.
+As to EEG signal where cognitive signals are sampled continuously, it’s easy to match certain sentence stimuli with corresponding signals.
 Its TSLR is similar to CSLR, which indicates the average percentage of certain text in test set appearing in training set.
@@
 $\frac{1}{M}\sum_{j=1}^{M}\min(1,\frac{|\{T_{ij}|T_{ij}\in S_{test}\cap S_{train}\}|}{|\{T_{ij}|T_{ij}\in S_{train}\}|})$ (8)
-where $M$ stands for the total number of unique text periods in test set and $T_{ij}$ stands for $j$-th period of text stimuli received by $i$-th subject, $T_{\cdot j}=T_{j}$, $i$ ranges from 1 to $N$.
-Figure 3: Validation loss of different dataset splitting methods in two datasets.
-fMRI data is sampled discretely with a deterministic interval, making it hard to acquire signals corresponding to sentences.
-Previous method instead concatenated continuous fMRI frames of certain length with their corresponding sentence segments as training sample.
-As a result, we consider the average percentage of the same sentence segments in test set appearing in training set as TSLR of fMRI data.
+where $M$ stands for the total number of unique text periods in test set and $T_{ij}$ stands for $j$-th period of text stimuli received by $i$-th subject.
+Figure 4: Validation loss of encoder under different dataset splitting methods in two datasets.
+The fMRI signal is sampled discretely with a deterministic interval TR, making it hard to acquire signals corresponding to sentences.
+Previous methods instead concatenated continuous fMRI frames of certain length with their corresponding sentence segments as training samples.
+As a result, we consider the average percentage of the same sentence segments in test set appearing in training set as TSLR of fMRI signal.
 It can be formulated as
-$\frac{1}{M}\sum_{j=1}^{M}\tau\frac{|\{T_{ij}|T_{ij}\in S_{test}\cap S_{train}\}|}{|S_{test}|\times L_{window}}$ (9)
+$\frac{1}{M}\sum_{j=1}^{M}\tau\frac{|\{T_{ij}|T_{ij}\in S_{test}\cap S_{train}\}|}{|S_{test}|\times L}$ (9)
 where $\tau=0$ if $\{T_{ij}|T_{ij}\in S_{test}\cap S_{train}\}=\emptyset$ else
 $\tau=\min(1,\frac{|\{T_{ij}|T_{ij}\in S_{train}\}|}{|\{T_{ij}|T_{ij}\in S_{test}\cap S_{train}\}|}).$ (10)
-We test current dataset splitting methods and our method on two cognitive datasets.
+We test current dataset splitting methods and our method on fMRI dataset Narratives and EEG dataset ZuCo.
 Considering the influence of randomness in splitting, we select four seeds for experiments.
-The results are shown in Table 1 and are aligned with theoretical analysis.
-For fMRI, current methods apart from method (a) suffer from cognitive signal data leakage, while method (a) has serious text stimuli data leakage.
-Method (b) gets no text stimuli data leakage but has slight cognitive signal data leakage.
+The results are shown in Table 1 and are consistent with theoretical analysis.
+For fMRI, current methods apart from method (a) suffer from cognitive signal leakage, while method (a) has serious text stimuli leakage.
+Method (b) gets no text stimuli leakage but has slight cognitive signal leakage.
 The situation for EEG is similar to that of fMRI.
@@
 EEG2Text 24.49 7.49 2.28 0.62 23.98 23.95 25.74
-Table 4: A fair benchmark for evaluating cognitive signals to text decoding.
+Table 4: A fair benchmark for evaluating brain-to-text decoding.
 4.4 Overestimated Performance
-We discover that cognitive signal leakage and text stimuli leakage affect model training from encoder side and decoder side respectively.
+Cognitive signal leakage and text stimuli leakage affect models from encoder side and decoder side.
 4.4.1 Encoder
-As shown in Figure 1, encoder in current models are trained in two different ways: either jointly trained with decoder or solely trained through a reconstruction task.
+As shown in Figure 1, encoder in current models is trained in two different ways: either jointly trained with decoder or solely trained through a reconstruction task.
 In the former end-to-end training scenario, it’s hard to evaluate encoder performance separately.
@@
 It’s similar to the structure of the encoder and will be abandoned once the encoder is trained.
-Although a proper evaluation index of the encoder’s representation ability is unavailable, validation loss is used to measure the influence of data contamination.
-We test splitting methods on two datasets and the validation loss of encoder is shown in Figure 3.
+Although a proper evaluation index of the encoder’s representation ability is missing, validation loss is used to measure the influence of data contamination.
+We test different splitting methods on two cognitive datasets.
+The validation loss of encoder is shown in Figure 4.
 For fMRI, influenced by leakage of cognitive signals, the validation loss of method (b), (c), (d), (e) keeps dropping even with long training epochs.
@@
 For fMRI signal, we test five current dataset splitting methods under different training settings.
-For fMRI, as shown in Table 2, we test two kinds of UniCoRN models.
-One is UniCoRN with finely tuned hyper-parameters claimed in the original paper, and the other is $*{}^{*}$ with randomly initialized encoder.
+As shown in Table 2, we test two kinds of UniCoRN models.
+One is UniCoRN with finely tuned hyper-parameters claimed in the original paper, and the other is $*{}^{*}$ with a randomly initialized encoder.
 Empirically, the former will perform much better than the latter.
@@
 4.5 Fair Benchmark
-We evaluate current state-of-the-art models for decoding cognitive signals to text under our proposed dataset splitting method and release a fair benchmark for further research.
+We evaluate current SOTA models for brain-to-text decoding under our dataset splitting method and release a fair benchmark.
 UniCoRN is tested for both fMRI and EEG decoding, EEG2Text model is tested for EEG decoding.
@@
 5 Conclusion
-In this paper, we explore a controversial topic: Due to the complexity of cognitive datasets, no consensus has been reached on how to split the dataset for training, validating, and testing in decoding cognitive signals to text task.
+In this paper, we explore a controversial topic: Due to the complexity of cognitive datasets, no consensus has been reached on how to split the dataset for training, validating, and testing in brain-to-text decoding.
 We analyze current dataset splitting methods and find serious data contamination largely exaggerates model performance and leads to poor generalization.
 Sufficient experiments and analysis are conducted to prove the existence of the above phenomena.
-Moreover, we propose a new dataset splitting method to fix this problem, which can avoid both cognitive signal and text stimuli leakage.
-State-of-the-art models are reevaluated in this setting and a fair benchmark is released for further research in this domain.
+Moreover, we propose a new dataset splitting method which can avoid both cognitive signal and text stimuli leakage to fix this problem.
+Current state-of-the-art models are reevaluated under this setting and a fair benchmark is released for further research in the domain.
 Limitations
 The "Narratives" dataset and the ZuCo dataset provide researchers with precise cognitive signal resources stimulated by text or voice.
-However, in decoding cognitive signals to text task, both subject’s brain signals and text stimuli in validation and test set need to be invisible to training set, which makes splitting these datasets difficult.
+However, in brain-to-text decoding task, both subject’s cognitive signals and text stimuli in validation and test set need to be invisible to training set, which makes splitting these public datasets difficult.
 Our proposed dataset splitting method meets the above requirements at the expense of discarding some data in the dataset.
@@
 The division of the training set, validation set, and test set should be provided when the dataset is released.
-Besides, we suggest hiring new subjects with unique stimuli for validation set and test set, which is good for testing the generalization ability of models without loss of data.
+Besides, we suggest hiring new subjects with unique stimuli for validation set and test set, which is good for testing the generalization ability of models without loss of data (Tang et al., 2023).
 What’s more, we find existing models rely more on the strong auto-regressive decoder to achieve good generation quality.
@@
 We perform the same filtering steps to Narratives dataset as UniCoRN paper (Xi et al., 2023) and the same filtering steps to ZuCo1.0 as EEG2Text paper (Wang and Ji, 2022).
+In CSLR and TSLR calculation, the number of four different seeds are set as $1,2,3,4$ respectively.
 In signal reconstruction task for encoder of UniCoRN, the batch size of EEG and fMRI data is 512 and 320 respectively.
@@
 Decoder is trained through 1e-4 learning rate and decaying to 1e-6 finally for 10 training epochs with 90 batch size.
-Sample length $L_{window}$ is set as 10 for all experiments related to fMRI.
+Sample length $L$ is set as 10 for all experiments related to fMRI.
 For EEG data, EEG2Text model is trained with 1e-6 learning rate for 80 epochs.
 UniCoRN model is trained with the same settings as fMRI data.
-Figure 4: The process of our proposed dataset splitting method.
 Appendix B Our Dataset Splitting Method
-In this part, we plot a sketch map of the dataset splitting process and release the pseudo-code of two dataset splitting methods for EEG and fMRI signal.
+In this part, we release the pseudo-code of two dataset splitting methods for EEG and fMRI signal.
 As shown in Figure 3, our proposed dataset splitting method consists of four steps.

### v3

@@
-Data Contamination Issues in Brain-to-Text Decoding
+Cross-Subject Data Splitting for Brain-to-Text Decoding
 Abstract
-Decoding non-invasive cognitive signals to natural language has long been the goal of building practical brain-computer interfaces (BCIs).
-Recent major milestones have successfully decoded cognitive signals like functional Magnetic Resonance Imaging (fMRI) and electroencephalogram (EEG) into text under open vocabulary setting.
-However, how to split the datasets for training, validating, and testing in brain-to-text decoding still remains controversial.
-In this paper, we conduct systematic analysis on current dataset splitting methods and find data contamination largely exaggerates model performance.
-Specifically, first we find the leakage of test subjects’ cognitive signals corrupts the training of a robust encoder.
-Second, we prove the leakage of text stimuli causes the auto-regressive decoder to memorize seen information in test set.
-To eliminate the influence of data contamination and fairly evaluate different models’ generalization ability, we propose a new splitting method for different types of cognitive dataset (e.g.
-fMRI, EEG).
-We also test the performance of SOTA brain-to-text decoding models under the proposed dataset splitting paradigm as baselines for further research.
-11footnotetext: Work was done when the author was an intern in JD.com.22footnotetext: Corresponding author.
+Recent major milestones have successfully decoded non-invasive brain signals (e.g.
+functional Magnetic Resonance Imaging (fMRI) and electroencephalogram (EEG)) into natural language.
+Despite the progress in model design, how to split the datasets for training, validating, and testing still remains a matter of debate.
+Most of the prior researches applied subject-specific data splitting, where the decoding model is trained and evaluated per subject.
+Such splitting method poses challenges to the utilization efficiency of dataset as well as the generalization of models.
+In this study, we propose a cross-subject data splitting criterion for brain-to-text decoding on various types of cognitive dataset (fMRI, EEG), aiming to maximize dataset utilization and improve model generalization.
+We undertake a comprehensive analysis on existing cross-subject data splitting strategies and prove that all these methods suffer from data leakage, namely the leakage of test data to training set, which significantly leads to overfitting and overestimation of decoding models.
+The proposed cross-subject splitting method successfully addresses the data leakage problem and we re-evaluate some SOTA brain-to-text decoding models as baselines for further research.
 1 Introduction
-Brain-computer interface (BCI) builds connections between human brain and external devices (e.g.
-computer).
-It has been widely researched in the field of neuroscience and has gained remarkable success like repairing damaged sight or restoring movement of disabled people (Polikov et al., 2005; Hochberg et al., 2012; Bouton et al., 2016).
-However, it’s still challenging in decoding cognitive signals to corresponding natural language when subjects read or hear text stimuli, especially for non-invasive cognitive signals like functional Magnetic Resonance Imaging (fMRI) or electroencephalogram (EEG) that are noisy and of low resolution (Mridha et al., 2021).
-Figure 1: General frameworks of brain-to-text decoding and possible situations of data contamination.
-Recent methods (Makin et al., 2020; Wang and Ji, 2022; Xi et al., 2023; Tang et al., 2023) typically viewed brain-to-text decoding as machine translation (Sutskever et al., 2014; Bahdanau et al., 2015) and adopted an encoder-decoder framework, where the encoder is responsible for converting cognitive signals into low-dimensional representations and the decoder learns to map the representations to natural language.
-Specifically, as shown in Figure1, the encoder usually consists of a spatial and time series feature extractor.
-It can be trained either in an end-to-end manner with decoder (Figure 1 a) or first pre-trained through a signal reconstruction task and then applied in decoder training (Figure 1 b).
-Despite recent success in model design, it still remains controversial in how to split the dataset for training, validating, and testing (Xi et al., 2023).
-Addressing this issue is urgent and meaningful, as fair evaluation of models is impossible without a widely recognized dataset splitting paradigm.
-A cognitive dataset is usually formatted in signal-sentence pair.
-In most cases for brain-to-text decoding task, each sentence belongs to a certain task, so signal-sentence pair can be further divided into signal-task and task-sentence pair.
-Current dataset splitting methods (Wang and Ji, 2022; Xi et al., 2023) can be summarized into five categories: (1) split by subjects, (2) split by tasks, (3) split by randomly picking signal frames, (4) split by randomly picking signal frames under certain task, (5) split by randomly picking consecutive signal frames under certain task.
-However, all these methods suffer from data contamination on encoder side, decoder side, or both.
-As shown in Figure 1, to encoder if subjects’ cognitive signals in test set are mixed into training set, encoder will become overfitted and fail to well represent unseen subjects’ cognitive signals.
-As to decoder, situation gets worse if text stimuli are leaked.
-Since the decoder generates token by token in an auto-regressive manner, data contamination will cause the decoder to memorize seen paragraphs and probability distribution, which means given the first few tokens the decoder is able to predict next token regardless of encoded cognitive signal representations.
-To address the above-mentioned problems, we propose a new dataset splitting method that eradicates data contamination from both encoder and decoder sides.
-We focus on fMRI and EEG signals in experiments, although the proposed splitting method could be applied to any cognitive signals satisfying the given format.
-In our method, the dataset is split according to subject-stimuli pairs with the following rules: (1) Cognitive signals collected from subjects in validation set and test set won’t appear in training set, which means the trained encoder can’t get access to any brain information belonging to subjects in test set.
-(2) Text stimuli in validation set and test set won’t appear in training set.
-The decoder learns the mapping between cognitive signal representation and token embedding instead of memorizing seen text.
+Brain-to-text decoding aims to recover natural language from brain signals stimulated by corresponding speech.
+Recent studies (Makin et al., 2020; Wang and Ji, 2022; Xi et al., 2023; Tang et al., 2023; Duan et al., 2024) have successfully decoded non-invasive brain signals (e.g.
+fMRI, EEG) to text by applying deep computational neural networks.
+However, no consensus has reached on how to split the cognitive dataset for training, validating, and testing.
+Most of the prior work (Ye et al., 2023; Tang et al., 2023) performed subject-specific data splitting for training and evaluating decoding models.
+Under this splitting rule, data for training, validating, and testing all comes from one specific subject from the cognitive dataset.
+For example, Tang et al.
+(2023) picked three subjects out of seven and conducted model training and evaluation on the three subjects respectively.
+This subject-specific splitting method causes two main problems.
+First, it only utilizes a tiny part of the whole dataset.
+Since the collection of brain signals is costly and time-consuming, such splitting method results in significant waste of data resources.
+Second, it leads to the poor generalization of decoding models.
+As every subject’s brain has unique functional and anatomical structures, subject-specific models may exhibit considerable variability across individuals and fail to generalize to other subjects.
+Moreover, decoding models trained from scratch on limited data are prone to facing the overfitting problem.
+Figure 1: Illustration of naturalistic language comprehension dataset for brain-to-text decoding. Path with the same color indicates one sample for training/validating/testing.
+Some studies (Wang and Ji, 2022; Xi et al., 2023) began to shed light on cross-subject data splitting, which views all the subjects’ data as a whole and performs splitting according to a given ratio (e.g.
+8:1:1 for the training set, validation set, and test set).
+Cross-subject data splitting effectively compensates for the shortcomings of subject-specific splitting and has been widely applied in brain-to-image decoding (Wang et al., 2024; Liu et al., 2024).
+However, unlike datasets for brain-to-image decoding, where subjects are guided to see different and unrepeated pictures, different subjects will hear the same story in the naturalistic language comprehension dataset for brain-to-text decoding, which challenges cross-subject data splitting.
+As shown in Figure 1, such dataset is usually formatted in subject-task-text-signal ($S$-$M$-$T$-$F$) pair, indicating the brain signal $F$ of subject $S$ stimulated by hearing text $T$ from task $M$.
+Current cross-subject data splitting methods (Wang and Ji, 2022; Xi et al., 2023) can be summarized as five categories: (1) split by subjects $S$, (2) split by tasks $M$, (3) split by randomly picking signal frames $F$, (4) split by randomly picking signal frames under certain task $F$-$M$, (5) split by randomly picking consecutive signal frames under certain task $F$-$M$.
+However, based on our observations, all these splitting methods suffer from data leakage problem, namely part of the test data is mixed into the training set, which leads to overfitting in model training and overestimation in model evaluation.
+Specifically, modern brain-to-text decoding models follow an encoder-decoder manner.
+We pick two representative models: EEG2Text (Wang and Ji, 2022) and UniCoRN (Xi et al., 2023) for investigating the damage of data leakage.
+The former is an end-to-end encoder-decoder framework, while the latter first pre-trains the encoder and then applies it in the decoder training.
+Experiments support that data leakage affects model training on both the encoder side and decoder side.
+For the encoder component, if subjects’ brain signals in the test set are mixed into the training set, the encoder will become overfitting and fail to well represent unseen subjects’ brain signals.
+As to the decoder, the situation gets worse if text stimuli are leaked.
+Since the decoder follows an auto-regressive manner and generates token by token, data leakage will cause the decoder to memorize seen paragraphs during the teacher-forcing training stage, which leads to poor generalization to unseen text.
+To avoid data leakage and fairly evaluate the performance of brain-to-text decoding models, we propose a cross-subject data splitting criterion.
+We focus on fMRI and EEG signals in this study, although the proposed criterion could be applied to any datasets satisfying the prescribed format.
+In the proposed method, the dataset is split according to subject-text ($S$-$T$) pairs with the following rules: (1) Brain signals collected from specific subject in validation set and test set will not appear in training set, which means the trained encoder cannot get access to any brain information belonging to subjects in test set.
+(2) Text stimuli in validation set and test set will not appear in training set.
+The decoder learns to reconstruct language with brain signals instead of memorizing seen text.
 Our contributions can be summarized as follows:
-  • We investigate current dataset splitting methods and analyze their influence on popular frameworks in brain-to-text decoding.
-  • We prove the existence of data contamination in current dataset splitting methods through analysis and experiments, which seriously exaggerates model performance.
-  • We propose the first splitting method without data contamination on public cognitive datasets.
-We also release a fair benchmark to evaluate different models’ generalization performance for further research in this domain.
-Figure 2: Different cognitive dataset splitting methods.
+  • To the best of our knowledge, we propose the first cross-subject data splitting criterion for brain-to-text decoding.
+  • We comprehensively analyze current cross-subject data splitting methods and find that all existing methods suffer from data leakage problem, which severely affects the training and evaluation of decoding models.
+  • Some SOTA brain-to-text decoding models are re-evaluated under the proposed cross-subject data splitting method as baselines for further research.
+Figure 2: Different splitting methods for cognitive dataset. (Color printing is preferred.)
 2 Related Work
-Cognitive Signal Cognitive signals can be classified into three categories: invasive, partially invasive, and non-invasive according to how close electrodes get to brain tissue.
-Due to the high cost and complexity of invasive and partially invasive methods, it’s hard to apply them in building generic and practical BCIs.
+Brain Signal
+Brain signals can be classified into three categories: invasive, partially invasive, and non-invasive according to how close electrodes get to brain tissue.
 In this paper, we mainly focus on non-invasive signals EEG and fMRI.
-EEG signal is electrogram of the spontaneous electrical activity of the brain.
-Its frequencies usually range from 1 to 30 Hz, divided into several groups like alpha (4-13 Hz), beta (13-30 Hz), delta (0.5-4 Hz), theta (4-7 Hz).
+EEG signal is electrogram of the spontaneous electrical activity of the brain, with frequencies ranging from 1 Hz to 30 Hz.
 EEG is of high temporal resolution and relatively tolerant of subject movement, but its spatial resolution is low and it can’t display active areas of the brain directly.
@@
 The spatial resolution of fMRI is measured by the size of voxel, which is a three-dimensional rectangular cuboid ranging from 3mm to 5mm (Vouloumanos et al., 2001; Noppeney and Price, 2004).
-Unlike EEG which samples brain signals continuously, fMRI samples according to the smallest time period named TR, usually at second level.
-Brain-to-text Decoding Previous research on brain-to-text decoding (Herff et al., 2015; Anumanchipalli et al., 2019; Zou et al., 2021; Moses et al., 2021; Défossez et al., 2023) mainly focused on word-level decoding in a restricted vocabulary with hundreds of words (Panachakel and Ramakrishnan, 2021).
-These models typically apply recurrent neural network or long short-term memory (Hochreiter and Schmidhuber, 1997) network to build mapping between cognitive signals and words in vocabulary.
+Unlike EEG which samples brain signals continuously, fMRI samples based on a fixed time interval named TR, usually at second level.
+Brain-to-text Decoding
+Previous research on brain-to-text decoding (Herff et al., 2015; Anumanchipalli et al., 2019; Zou et al., 2021; Moses et al., 2021; Défossez et al., 2023) mainly focused on word-level decoding in a restricted vocabulary with hundreds of words (Panachakel and Ramakrishnan, 2021).
+These models typically apply recurrent neural network or long short-term memory (Hochreiter and Schmidhuber, 1997) network to build mapping between brain signals and words in vocabulary.
 Despite relatively good accuracy, these methods fail to generalize to unseen words.
-Some progress (Sun et al., 2019) has been made by expanding word-level decoding to sentence-level through encoder-decoder framework, or use less noisy ECoG data (Burle et al., 2015; Anumanchipalli et al., 2019).
+Some progress (Sun et al., 2019) has been made by expanding word-level decoding to sentence-level through encoder-decoder framework or using less noisy ECoG data (Burle et al., 2015; Anumanchipalli et al., 2019).
 However, these models struggle to generate accurate and fluent sentences limited by decoder ability.
@@
 3 Methodology
-In this section, we will first introduce the definition of decoding cognitive signals to text and the general description of dataset format.
-Then we systematically analyze current dataset splitting methods and point out that all existing methods suffer from two kinds of data contamination: cognitive signal leakage and text stimuli leakage.
-Finally, a new dataset splitting method is proposed to avoid the above-mentioned two kinds of data contamination.
+In this section, we will first introduce the formal definition of brain-to-text decoding and the general description of dataset format.
+Then we systematically analyze current cross-subject data splitting methods and point out that all existing methods suffer from two kinds of data leakage issues: brain signal leakage and text stimuli leakage.
+Finally, a cross-subject splitting criterion is proposed which avoids the above-mentioned data leakage problems.
 3.1 Task Definition
-Given the cognitive signal $F_{ij}$ stimulated by $i$-th subject $S_{i}$ hearing or reading certain text $T_{j}$, brain-to-text decoding aims to decode $F_{ij}$ back to text $T_{j}^{\prime}$ and make $T_{j}^{\prime}$ as similar as possible to $T_{j}$.
+Given the brain signal $F_{ij}$ stimulated by $i$-th subject $S_{i}$ hearing or reading certain text $T_{j}$, brain-to-text decoding aims to decode $F_{ij}$ back to text $T_{j}^{\prime}$ and make $T_{j}^{\prime}$ as similar as possible to $T_{j}$.
 The composition of $F_{ij}$ and $T_{j}$ is different as to fMRI and EEG.
 The former samples brain information discretely with a fixed time interval TR, while the latter samples continuously.
-To fMRI, $L$ consistent sentence segments $s_{j}$ with corresponding fMRI frames $f_{ij}$ are concatenated to form a sample pair $\left\langle F_{ij},T_{j}\right\rangle$.
-Here $T_{j}=concat(s_{j},s_{j+1},\ldots,s_{j+L-1})$ and $F_{ij}=concat(f_{ij},f_{i,j+1},\ldots,f_{i,j+L-1})$.
-To EEG, since signals corresponding to a complete sentence are available, we bond sentence (aka text stimuli) $T_{j}$ and EEG signal $F_{ij}$ together to form a sample pair $\left\langle F_{ij},T_{j}\right\rangle$.
-Under most circumstances, each sentence $T_{j}$ belongs to one certain task $M_{k}$.
-So the signal-sentence pair $\left\langle F_{ij},T_{j}\right\rangle$ can be further split into $\left\langle F_{ij},M_{k}\right\rangle$ and $\left\langle M_{k},T_{kj}\right\rangle$.
-Figure 3: The process of our proposed dataset splitting method.
-In brain-to-text decoding, the ultimate goal of trained BCI models is to generalize to unseen subjects with unseen text stimuli (Huang et al., 2010; Handiru and Prasad, 2016; Gao et al., 2021).
-As a result, if cognitive signal $F_{ij}$ appears in test set $S_{test}$, then any subject $i$’s signal $F_{i}$ should not appear in training set $S_{train}$.
+To fMRI, consistent text segments $s_{j}$ with corresponding fMRI frames $f_{ij}$ are concatenated to form a sample pair $\left\langle F_{ij},T_{j}\right\rangle$, where $T_{j}=concat(s_{j},s_{j+1},\ldots,s_{j+L-1})$ and $F_{ij}=concat(f_{ij},f_{i,j+1},\ldots,f_{i,j+L-1})$, and $|T_{j}|=|F_{ij}|=L$.
+To EEG, since signals corresponding to complete text stimuli are available and they are continuous, we bond text $T_{j}$ (i.e.
+text stimuli) and EEG signal $F_{ij}$ together to form a sample pair $\left\langle F_{ij},T_{j}\right\rangle$.
+Under most scenarios, each text stimulus $T_{j}$ belongs to one certain task $M_{k}$.
+So the signal-text pair $\left\langle F_{ij},T_{j}\right\rangle$ can be further split into $\left\langle F_{ij},M_{k}\right\rangle$ and $\left\langle M_{k},T_{kj}\right\rangle$.
+Figure 3: The process of our proposed cross-subject data splitting method. (Color printing is preferred.)
+One of the purposes of cross-subject splitting is to endow models the ability to decode unseen subject’s brain signal.
+As a result, if brain signal $F_{ij}$ appears in test set $S_{test}$, any signal $F_{i*}$ belonging to subject $i$ should not appear in training set $S_{train}$.
 Similarly, text stimuli $T_{kj}$ in $S_{test}$ should not appear in $S_{train}$.
-So the dataset splitting rules can be formally defined as:
-$S_{train}=\{\left\langle F_{ij},T_{kj}\right\rangle|F_{i},T_{kj}\notin S_{val},S_{test}\}$ (1)
-$S_{val}=\{\left\langle F_{ij},T_{kj}\right\rangle|F_{i},T_{kj}\notin S_{train},S_{test}\}$ (2)
-$S_{test}=\{\left\langle F_{ij},T_{kj}\right\rangle|F_{i},T_{kj}\notin S_{train},S_{val}\}$ (3)
-3.2 Dataset Splitting Methods
-Current dataset splitting methods can be summarized as five categories according to classifying objectives $S_{i},M_{k},T_{kj},F_{ij}$.
-More specifically, the five dataset splitting methods are characterised as (1) split by subjects, (2) split by tasks, (3) split by randomly picking signal frames, (4) split by randomly picking signal frames under certain task, (5) split by randomly picking consecutive signal frames under certain task, corresponding to image (a), (b), (c), (d), (e) in Figure 2.
+The above splitting rules for training set can be formulated by Cartesian product:
+$S_{train}=F_{train}\times T_{train},$ (1)
+where
+$F_{train}=\{F_{ij}|i\in I\},$ (2)
+$I=\{i|F_{ij}\notin S_{val},S_{test},\forall j\},$ (3)
+and
+$T_{train}=\{T_{kj}|T_{kj}\notin S_{val},S_{test}\}.$ (4)
+Such rules are also applicable to validation and test set. We omit their displays here for simplicity.
+3.2 Current Splitting Methods
+Current cross-subject data splitting methods can be summarized as five categories according to classifying objectives $S_{i},M_{k},T_{kj},F_{ij}$.
+More specifically, the five dataset splitting methods are characterised as (1) split by subjects $S_{i}$, (2) split by tasks $M_{k}$, (3) split by randomly picking signal frames $F_{ij}$, (4) split by randomly picking signal frames under certain task $F_{ij}$-$M_{k}$, (5) split by randomly picking consecutive signal frames under certain task $F_{ij}$-$M_{k}$, corresponding to (a), (b), (c), (d), (e) in Figure 2.
 Figure 2 vividly displays the differences between current dataset splitting methods.
-For simplicity of expression, we choose four subjects with three tasks each containing four, three, four sentences.
-The line connecting two symbols indicates they are related to one sample.
-Take path $S_{1},M_{1},T_{11},F_{11}$ for example, it’s one sample where subject $S_{1}$ listens to text stimuli $T_{11}$ belonging to task $M_{1}$ and $S_{1}$’s corresponding brain signal is recorded as $F_{11}$.
+In this example, we choose 4 subjects ($S_{1}$ to $S_{4}$) with 3 tasks ($M_{1}$ to $M_{3}$) each containing 4 ($T_{11}$ to $T_{14}$), 3 ($T_{21}$ to $T_{23}$), 4 ($T_{31}$ to $T_{34}$) text stimuli respectively.
+$F_{ij}$ indicates the brain signal of $i$-th subject stimulated by $j$-th text under task $M_{k}$, e.g.
+$F_{21}$ means brain signal of $S_{2}$ hearing $T_{11}$.
+The line connecting two symbols indicates they are related to one sample in dataset.
+Take path $S_{1},M_{1},T_{11},F_{11}$ for example, it indicates that subject $S_{1}$ listens to text stimuli $T_{11}$ belonging to task $M_{1}$ and $S_{1}$’s corresponding brain signal is recorded as $F_{11}$.
 Some symbols are connected with several lines.
@@
 The green lines and orange lines stand for training samples and testing samples.
-The grey dotted line means the sample is abandoned, which will be introduced in our dataset splitting method.
-As the splitting of validation set is similar to test set, we only consider training set and test set in this section.
+The grey dotted line means the sample is abandoned, which will be introduced in our data splitting method.
+As the splitting of validation set is the same as test set, we only consider training set and test set in this section for simplicity.
 We will use method (a), (b), (c), (d), (e) to represent five current dataset splitting methods in the rest of the paper. Method (a) splits the dataset according to subjects, which can be described as
 $S_{train}=\{\left\langle F_{ij},T_{kj}\right\rangle|S_{i}\notin S_{val},S_{test}\}$ (5)
-for training set. Method (b) splits the dataset according to tasks, which can be described as
+for training set. Method (b) splits the dataset according to tasks, which is described as
 $S_{train}=\{\left\langle F_{ij},T_{kj}\right\rangle|M_{k}\notin S_{val},S_{test}\}$ (6)
-for training set. Method (c), (d), and (e) all split the dataset according to cognitive signal frames
+for training set. Method (c), (d), and (e) all split the dataset according to brain signal frames
 $S_{train}=\{\left\langle F_{ij},T_{kj}\right\rangle|F_{ij}\notin S_{val},S_{test}\}.$ (7)
-However, there are slight differences between these methods.
-Method (c) views all the cognitive signal frames in dataset as a whole and splits according to the default proportion (e.g.
+However, there are slight differences between these three methods.
+Method (c) views all the brain signal frames in dataset as a whole and splits according to the default proportion (e.g.
 8:1:1).
-Method (d) views signal frames under certain task $M_{k}$ as a whole and splits proportionally, and then union all training sets under different tasks to form a complete set for training.
+Method (d) views signal frames under certain task $M_{k}$ as a whole and splits proportionally, and then unions all training sets under different tasks to form a complete set for training.
 Method (e) is similar to method (d).
 They both first split training, validation, and test set under certain task proportionally and then union them.
-The difference lies in that method (d) randomly picks signal frames while method (e) picks consecutive signal frames.
+The difference lies in that method (d) randomly picks signal frames, while method (e) picks consecutive signal frames.
 Type Method Narratives / ZuCo Average
@@
 (f) 0.00 / 0.00 0.00 / 0.00 0.00 / 0.00 0.00 / 0.00 0.00 / 0.00
-Table 1: Results of Cognitive Signal Leakage Rate (CSLR) and Text Stimuli Leakage Rate (TSLR).
-The goal of brain-to-text decoding models is to generalize to unseen subjects with unseen text stimuli, which means both subject’s brain information and received text stimuli are new to the trained model.
-In this sense, we define two kinds of data contamination: cognitive signal leakage and text stimuli leakage.
-The data contamination situation of different methods is reflected in Figure 2.
+Table 1: Results of Brain Signal Leakage Rate (BSLR) and Text Stimuli Leakage Rate (TSLR).
+We first point out the data leakage problems in current splitting methods through the analysis of training set and test set composition.
+Specifically, following the definition in subsection 3.1, two kinds of data leakage, brain signal leakage and text stimuli leakage, are defined.
+The data leakage situation of different methods can be reflected through visualization in Figure 2.
+Lines between $S_{i}$ and $M_{k}$ indicate brain signal leakage situation and lines between $T_{kj}$ and $M_{k}$ indicate text stimuli leakage situation.
 If lines associated with $S_{i}$ or $T_{kj}$ are of different colours, data in test set leaks into training set.
-Lines between $S_{i}$ and $M_{k}$ indicate cognitive signal leakage situation and lines between $T_{kj}$ and $M_{k}$ indicate text stimuli leakage situation.
-Remind the composition of samples differs as to fMRI signal and EEG signal, so the dataset splitting methods are different for two cognitive signals too.
-Since fMRI signals need to be sampled continuously with a certain length $L$, the path of a sample shown in Figure 2 is actually the first part of one fMRI sample, with $L-1$ continuous part following.
-In this sense, for EEG cognitive signal leakage doesn’t exist in method (a), but method (a) suffers from text stimuli leakage.
-The situation of method (b) is opposite to that of method (a), where there’s no text stimuli leakage but cognitive signal leakage.
-Method (c) and method (d) are similar.
-They suffer from both cognitive signal leakage and text stimuli leakage.
-Method (e) is in the same situation as method (b).
-For fMRI, method (c), (d), and (e) which seem the same for EEG are actually different splitting ways.
-The situation of data contamination for different methods is similar to EEG, except for method (e) there still exists slight text stimuli leakage in the overlap between training samples and test samples.
-3.3 Our Method
-To eliminate data contamination from both cognitive signal leakage and text stimuli leakage, we split the dataset by $\left\langle S_{i},T_{j}\right\rangle$ pairs as shown in (f) of Figure 2.
-Since EEG and fMRI are different in the composition of dataset, we treat them separately and propose two dataset splitting methods.
+Remind the composition of samples differs as to fMRI signal and EEG signal, so the dataset splitting methods are different for two kinds of brain signal too.
+Since fMRI signals need to be sampled continuously with a certain length $L$, one sample shown in Figure 2 (e.g.
+$S_{1}$-$M_{1}$-$T_{11}$-$F_{11}$) is actually the first part of one fMRI sample, with $L-1$ continuous brain signal frames following (e.g.
+$F_{12},F_{13},\ldots$).
+In this sense, brain signal leakage doesn’t exist in method (a) for EEG, but method (a) suffers from text stimuli leakage.
+Text stimuli leakage does not exist in method (b) but brain signal leakage exists.
+Method (c) suffers from both brain signal leakage and text stimuli leakage.
+Method (d) and (e) are the same to EEG, with brain signal leakage.
+For fMRI, the situation of data leakage for different methods is similar to EEG, except for method (d) and (e), which are the same for EEG but actually different for fMRI.
+Method (d) suffers from both brain signal leakage and text stimuli leakage while in method (e) text stimuli leakage happens in the overlap between training samples and test samples.
+3.3 Cross-Subject Splitting Criterion
+To eliminate data leakage from both brain signal leakage and text stimuli leakage, we split the dataset by $\left\langle S_{i},T_{j}\right\rangle$ pairs as shown in (f) of Figure 2.
+Since EEG and fMRI are different in the composition of dataset, we treat them separately and propose two data splitting methods.
 As to EEG dataset where $F_{ij}$ and $T_{j}$ form a sample, we consider a bipartite graph $\mathcal{G}_{1}=(\mathcal{U},\mathcal{V},\mathcal{E})$ where $\mathcal{U}=\{S_{i}\}_{i=1}^{N}$, $\mathcal{V}=\{T_{j}\}_{j=1}^{M}$.
@@
 As shown in step 2 of Figure 3, first we pick one edge for each node $v\in\mathcal{V}$ and build a new bipartite graph $\mathcal{G}_{2}=(\mathcal{U},\mathcal{V},\mathcal{E}^{\prime})$.
-Then we split graph $\mathcal{G}_{2}$ by subject $\mathcal{U}$ with the given splitting ratio and form three disjoint graphs $\mathcal{G}_{train},\mathcal{G}_{val},\mathcal{G}_{test}$.
-In step 4, some edges satisfying zero data contamination condition are not included in the graph.
+Then following step 3, we split graph $\mathcal{G}_{2}$ by subject $\mathcal{U}$ with the given splitting ratio and form three disjoint graphs $\mathcal{G}_{train},\mathcal{G}_{val},\mathcal{G}_{test}$.
+In step 4, some edges satisfying zero data leakage condition are not included in the graph.
 We add these edges to corresponding graphs, extending each graph $\mathcal{G}_{train},\mathcal{G}_{val},\mathcal{G}_{test}$ to its maximally scalable state and finishing the dataset splitting process.
-$F_{ij}=concat(f_{ij},f_{i,j+1},\ldots,f_{i,j+L-1})$ and $T_{j}=concat(s_{j},s_{j+1},\ldots,s_{j+L-1})$ form a sample pair in fMRI dataset.
+The sample in fMRI dataset is formatted in $F_{ij}=concat(f_{ij},f_{i,j+1},\ldots,f_{i,j+L-1})$ and $T_{j}=concat(s_{j},s_{j+1},\ldots,s_{j+L-1})$.
 If we follow the same process as EEG, text stimuli leakage will occur in the overlapping part of two samples, when one sample is assigned to training set and the other is assigned to validation or test set.
-We propose a simple solution that achieves the balance between abandoning as little data as possible and ensuring zero data contamination.
+We propose a simple solution that achieves the balance between abandoning as little data as possible and ensuring zero data leakage.
 Instead of $\left\langle S_{i},T_{j}\right\rangle$ pair, we consider $\left\langle S_{i},M_{k}\right\rangle$ pair and apply the above-mentioned algorithm.
@@
 […]
-$N=1$ $N=2$ $N=3$ $N=4$ $F$ $P$ $R$
+$N=1$ $N=2$ $N=3$ $N=4$ $R$ $P$ $F$
 UniCoRN 10+1e-3+(a) 49.56 30.49 21.07 15.49 44.83 50.41 40.65
@@
 10+1e-3+(e) 28.94 9.39 4.07 1.53 21.68 24.64 19.49
-$*{}^{*}$ 20+1e-4+(a) 50.19 34.25 25.98 21.00 46.59 50.36 43.62
+UniCoRN 20+1e-4+(a) 50.19 34.25 25.98 21.00 46.59 50.36 43.62
 30+1e-4+(a) 55.46 40.99 32.85 27.56 52.08 55.02 49.68
@@
 Table 2: Generation quality of UniCoRN model for fMRI under different training settings.
-Here $*{}^{*}$ indicates the encoder of UniCoRN is randomly initialized instead of pre-trained through signal reconstruction task.
-4 Experiments and Analysis
-We test state-of-the-art brain-to-text decoding models on two popular cognitive datasets.
-Comprehensive experiments are conducted to prove the existence of the following phenomena: (1) Cognitive signals and text stimuli in test set leak into training set in all current dataset splitting methods.
-(2) The model’s generalization ability, particularly that of the auto-regressive decoder, has been overestimated due to data contamination.
-Because the number of tasks in EEG dataset is too small and method (e) makes no difference to EEG as method (d), we only consider method (a), (c), (d).
-4.1 Datasets
-[…]
+Here UniCoRN indicates the encoder of UniCoRN is randomly initialized instead of pre-trained through signal reconstruction task.
+4 Experimental Settings
+We test some SOTA brain-to-text decoding models on two popular cognitive datasets Narratives (Nastase et al., 2021) and ZuCo (Hollenstein et al., 2018).
+Dataset details are introduced in Appendix A.
+Comprehensive experiments are conducted to evidence the existence of the following phenomena: (1) Brain signals and text stimuli in test set leak into training set in all current dataset splitting methods.
+(2) The model’s generalization ability, particularly that of the auto-regressive decoder, has been overestimated due to data leakage.
+Because the number of tasks in EEG dataset is too small and method (e) makes no difference to EEG as method (d), we only consider method (a), (c), (d) for EEG.
 4.1 Implementation
-We follow the same settings of UniCoRN (Xi et al., 2023) and EEG2Text (Wang and Ji, 2022), except the datasets are split to the ratio of 8:1:1 for fair comparison.
+We follow the same settings of UniCoRN (Xi et al., 2023) and EEG2Text (Wang and Ji, 2022), except all the datasets are split to the ratio of 8:1:1 for fair comparison.
 All experiments are conducted on NVIDIA A100-SXM4-40GB GPUs.
@@
 […]
-$N=1$ $N=2$ $N=3$ $N=4$ $F$ $P$ $R$
+$N=1$ $N=2$ $N=3$ $N=4$ $R$ $P$ $F$
 UniCoRN 50+1e-4+(a) 58.09 49.23 43.23 38.43 63.88 61.12 67.50
@@
 Table 3: Generation quality of UniCoRN and EEG2Text model for EEG under different training settings.
-4.3 Data Contamination
-We have analyzed two kinds of data contamination, cognitive signal leakage and text stimuli leakage in Methodology section.
-In this part, we will quantify data contamination situation through experiments.
-To better illustrate the extent of data contamination across different dataset splitting methods, we design two novel evaluation metrics named Cognitive Signal Leakage Rate (CSLR) and Text Stimuli Leakage Rate (TSLR) for detecting cognitive signal leakage and text stimuli leakage.
-Note that the situation for validation set is similar as test set, we only consider test set in experiments.
-CSLR indicates the average percentage of each subject’s cognitive signals in test set appearing in training set, which could be formulated as
+4.2 Data Leakage Metrics
+We have analyzed two kinds of data leakage: brain signal leakage and text stimuli leakage in section 3. In this part, we will quantify two kinds of data leakage through experiments.
+To better illustrate the extent of data leakage of different data splitting methods, we design two novel evaluation metrics named Brain Signal Leakage Rate (BSLR) and Text Stimuli Leakage Rate (TSLR) for detecting brain signal leakage and text stimuli leakage.
+Note that the situation for validation set is the same as test set, so we only consider test set in experiments.
+BSLR indicates the average percentage of each subject’s brain signals in test set appearing in training set, which could be formulated as
 $\frac{1}{N}\sum_{i=1}^{N}\min(1,\frac{|\{F_{ij}|F_{ij}\in S_{test}\cap S_{train}\}|}{|\{F_{ij}|F_{ij}\in S_{train}\}|})$ (8)
@@
 ||$\cdot$|| ⋅ | stands for the cardinality of a set.
-Function $\min(\cdot,\cdot)$ is applied to make sure for each subject the data leakage rate is less than 1.
+Function $\min(\cdot,\cdot)$ is applied to make sure for each subject the data leakage rate is less than one.
 The definition of TSLR is somewhat different for EEG signal and fMRI signal.
-As to EEG signal where cognitive signals are sampled continuously, it’s easy to match certain sentence stimuli with corresponding signals.
-Its TSLR is similar to CSLR, which indicates the average percentage of certain text in test set appearing in training set.
+As to EEG signal where brain signals are sampled continuously, it’s easy to match certain text stimuli with corresponding signals.
+Its TSLR is similar to BSLR, which indicates the average percentage of certain text in test set appearing in training set.
 TSLR for EEG data can be calculated through
@@
 where $M$ stands for the total number of unique text periods in test set and $T_{ij}$ stands for $j$-th period of text stimuli received by $i$-th subject.
-Figure 4: Validation loss of encoder under different dataset splitting methods in two datasets.
-The fMRI signal is sampled discretely with a deterministic interval TR, making it hard to acquire signals corresponding to sentences.
-Previous methods instead concatenated continuous fMRI frames of certain length with their corresponding sentence segments as training samples.
-As a result, we consider the average percentage of the same sentence segments in test set appearing in training set as TSLR of fMRI signal.
+The fMRI signal is sampled discretely with a deterministic interval TR, making it hard to acquire signals corresponding to text.
+Previous methods instead concatenated continuous fMRI frames of certain length with their corresponding text segments as training samples.
+As a result, we consider the average percentage of the same text segments in test set appearing in training set as TSLR for fMRI signal.
 It can be formulated as
@@
 $\tau=\min(1,\frac{|\{T_{ij}|T_{ij}\in S_{train}\}|}{|\{T_{ij}|T_{ij}\in S_{test}\cap S_{train}\}|}).$ (11)
-We test current dataset splitting methods and our method on fMRI dataset Narratives and EEG dataset ZuCo.
-Considering the influence of randomness in splitting, we select four seeds for experiments.
-The results are shown in Table 1 and are consistent with theoretical analysis.
-For fMRI, current methods apart from method (a) suffer from cognitive signal leakage, while method (a) has serious text stimuli leakage.
-Method (b) gets no text stimuli leakage but has slight cognitive signal leakage.
-The situation for EEG is similar to that of fMRI.
-Apart from our proposed method (f), there is no way to achieve zero cognitive signal leakage and text stimuli leakage at the same time.
 […]
-$N=1$ $N=2$ $N=3$ $N=4$ $F$ $P$ $R$
+$N=1$ $N=2$ $N=3$ $N=4$ $R$ $P$ $F$
 Narratives UniCoRN 22.83 5.69 1.43 0.48 15.55 24.80 19.04
@@
 Table 4: A fair benchmark for evaluating brain-to-text decoding.
-4.4 Overestimated Performance
-Cognitive signal leakage and text stimuli leakage affect models from encoder side and decoder side.
-4.4.1 Encoder
-As shown in Figure 1, encoder in current models is trained in two different ways: either jointly trained with decoder or solely trained through a reconstruction task.
-In the former end-to-end training scenario, it’s hard to evaluate encoder performance separately.
-So we mainly focus on the latter, in which case the encoder is trained through an encoder-decoder framework to reconstruct input cognitive signals.
-The decoder here doesn’t refer to the decoder for text generation.
-It’s similar to the structure of the encoder and will be abandoned once the encoder is trained.
-Although a proper evaluation index of the encoder’s representation ability is missing, validation loss is used to measure the influence of data contamination.
+5 Results and Analysis
+5.1 Verification for Data Leakage
+We test current data splitting methods and our method on fMRI dataset “Narratives” and EEG dataset ZuCo.
+Considering the influence of randomness in splitting, we randomly select four seeds for experiments.
+The results are shown in Table 1, and are consistent with theoretical analysis.
+For fMRI, current methods apart from method (a) suffer from brain signal leakage, while method (a) has serious text stimuli leakage.
+Method (b) gets no text stimuli leakage but has slight brain signal leakage.
+The situation for EEG is similar to that of fMRI.
+Apart from our proposed method (f), there is no way to achieve zero brain signal leakage and text stimuli leakage at the same time.
+5.2 Damage of Data Leakage
+Brain signal leakage and text stimuli leakage will damage brain-to-text decoding models from both the encoder side and decoder side.
+Effect on Encoder
+The encoder of current brain-to-text decoding models can be trained in two ways: either jointly with the decoder or solely pre-trained through a reconstruction task.
+In the former end-to-end training scenario, it is hard to evaluate encoder performance separately.
+So we mainly focus on the latter, in which case the encoder is pre-trained through an encoder-decoder framework to reconstruct input brain signals.
+The decoder here does not refer to the decoder for text generation.
+It is similar to the structure of the encoder and will be abandoned once the encoder is pre-trained.
+Since a proper evaluation index of the encoder’s representation ability is missing, validation loss is applied to measure the effect of data leakage.
 We test different splitting methods on two cognitive datasets.
 The validation loss of encoder is shown in Figure 4.
-For fMRI, influenced by leakage of cognitive signals, the validation loss of method (b), (c), (d), (e) keeps dropping even with long training epochs.
+For fMRI, influenced by leakage of brain signals, the validation loss of method (b), (c), (d), (e) keeps dropping even with long training epochs.
 The encoder is actually overfitting and degrading.
-For method (a) and (f) without cognitive signal leakage, the validation loss quickly rises after reaching the lowest point with a few epochs, satisfying the basic rule of machine learning.
-For EEG, we find validation loss keeps dropping for all methods even with very long training epochs, regardless of cognitive signal leakage or not.
+For method (a) and (f) without brain signal leakage, the validation loss quickly rises after reaching the lowest point with a few epochs, satisfying the basic rule of machine learning.
+For EEG, we find validation loss keeps dropping for all methods even with very long training epochs, regardless of brain signal leakage or not.
 We think the poor spatial resolution of EEG signal might lead to this phenomenon.
-4.4.2 Decoder
-All state-of-the-art models choose pre-trained language model BART (Lewis et al., 2020) as decoder.
-On one hand, the powerful auto-regressive decoder is able to achieve fluent sentence-level open vocabulary text generation.
-On the other hand, if data contamination occurs, due to the feature of auto-regressive generation, the decoder will generate memorized text given the first few words, which is obviously an act of cheating.
+Figure 4: Validation loss of encoder under different dataset splitting methods in two datasets.
+Effect on Decoder
+All SOTA models choose the pre-trained language model BART (Lewis et al., 2020) as decoder.
+The powerful model is able to achieve fluent open vocabulary text generation.
+However, if data leakage occurs, due to the auto-regressive generation that calculates the probability of current token based on all previous tokens, the decoder will generate seen text given the first few words, and fail to generalize to unseen text.
 The influence of text stimuli leakage on decoder is detected through BLEU scores (Papineni et al., 2002) and ROUGE-1 scores (Lin, 2004), which measure text similarity between generated text and ground truth.
@@
 As shown in Table 2, we test two kinds of UniCoRN models.
-One is UniCoRN with finely tuned hyper-parameters claimed in the original paper, and the other is $*{}^{*}$ with a randomly initialized encoder.
+One is UniCoRN with finely tuned hyper-parameters claimed in the original paper, and the other is UniCoRN with a randomly initialized encoder.
 Empirically, the former will perform much better than the latter.
-However, in method (a), (c), (d), due to text stimuli leakage, if we reduce the learning rate and extend training epochs, $*{}^{*}$ performs much better than UniCoRN and its performance keeps rising with longer training epochs.
+However, in method (a), (c), (d), due to text stimuli leakage, if we reduce the learning rate and extend training epochs, UniCoRN performs much better than UniCoRN and its performance keeps rising with longer training epochs.
 As to method (b) and (e) with no text stimuli leakage, changing training epochs or learning rates makes no obvious difference to model performance.
@@
 For method (d) without text stimuli leakage, both models reach optimal performance after the first few rounds of training epochs.
-4.5 Fair Benchmark
-We evaluate current SOTA models for brain-to-text decoding under our dataset splitting method and release a fair benchmark.
+5.3 A Fair Benchmark
+We re-evaluate current SOTA models for brain-to-text decoding under our cross-subject data splitting method and release a fair benchmark.
 UniCoRN is tested for both fMRI and EEG decoding, EEG2Text model is tested for EEG decoding.
@@
 6 Conclusion
-In this paper, we explore a controversial topic: Due to the complexity of cognitive datasets, no consensus has been reached on how to split the dataset for training, validating, and testing in brain-to-text decoding.
-We analyze current dataset splitting methods and find serious data contamination largely exaggerates model performance and leads to poor generalization.
-Sufficient experiments and analysis are conducted to prove the existence of the above phenomena.
-Moreover, we propose a new dataset splitting method which can avoid both cognitive signal and text stimuli leakage to fix this problem.
-Current state-of-the-art models are reevaluated under this setting and a fair benchmark is released for further research in the domain.
+In this paper, we conduct a comprehensive study on existing cross-subject data splitting methods, and evidence that all these methods suffer from data leakage problem.
+Such data leakage largely exaggerates model performance and leads to poor generalization.
+To fix this issue, we propose a cross-subject data splitting criterion for brain-to-text decoding, aiming to improve the utilization efficiency of cognitive dataset and the generalization ability of decoding models.
+Experiments are conducted on fMRI and EEG dataset respectively.
+Current SOTA models are re-evaluated under this proposed splitting method and a fair benchmark is released for further research in this domain.
 Limitations
-The "Narratives" dataset and the ZuCo dataset provide researchers with precise cognitive signal resources stimulated by text or voice.
-However, in brain-to-text decoding task, both subject’s cognitive signals and text stimuli in validation and test set need to be invisible to training set, which makes splitting these public datasets difficult.
+The “Narratives” dataset and the ZuCo dataset provide researchers with precise brain signal resources stimulated by text or voice.
+However, in the brain-to-text decoding task, both subject’s brain signals and text stimuli in the validation and test set need to be invisible to the training set, which makes splitting these public datasets difficult.
 Our proposed dataset splitting method meets the above requirements at the expense of discarding some data in the dataset.
@@
 The division of the training set, validation set, and test set should be provided when the dataset is released.
-Besides, we suggest hiring new subjects with unique stimuli for validation set and test set, which is good for testing the generalization ability of models without loss of data (Tang et al., 2023).
-What’s more, we find existing models rely more on the strong auto-regressive decoder to achieve good generation quality.
-The encoder is of limited use in all SOTA models, which might become a research point in the future.
+Besides, we suggest hiring new subjects with unique stimuli for the validation set and test set, which is good for testing the generalization ability of models without loss of data.
+What’s more, we find existing models rely more on a strong auto-regressive decoder to achieve good generation quality.
+The encoder is of limited use in all SOTA models.
+And we also notice in experiments that the encoder of EEG2Text keeps overfitting whether with or without brain signal leakage.
+We leave it to research in the future.
 Ethics Statement
-In this paper, we introduce a new dataset splitting method to avoid data contamination for decoding cognitive signals to text task.
-Experiments are conducted on public accessible cognitive datasets Narratives and ZuCo1.0 with the authorization from their respective maintainers.
+In this paper, we introduce a new dataset splitting method to avoid data leakage for decoding brain signals to text task.
+Experiments are conducted on the publicly accessible cognitive datasets “Narratives” and ZuCo1.0 with the authorization from their respective maintainers.
 Both datasets have been de-identified by dataset providers and used for researches only.
-Acknowledgements
-This research is supported by the National Natural Science Foundation of China (No.62106105), the CCF-Baidu Open Fund (No.CCF-Baidu202307), the CCF-Zhipu AI Large Model Fund (No.CCF-Zhipu202315), the Scientific Research Starting Foundation of Nanjing University of Aeronautics and Astronautics (No.YQR21022), and the High Performance Computing Platform of Nanjing University of Aeronautics and Astronautics.
-We appreciate JD.com for providing powerful GPUs for experiments.
 Appendix A Implementation Details
+[…]
 More details in experiments are supplemented in this section.
-We perform the same filtering steps to Narratives dataset as UniCoRN paper (Xi et al., 2023) and the same filtering steps to ZuCo1.0 as EEG2Text paper (Wang and Ji, 2022).
-In CSLR and TSLR calculation, the number of four different seeds are set as $1,2,3,4$ respectively.
+We perform the same filtering steps to “Narratives” dataset as UniCoRN paper (Xi et al., 2023) and the same filtering steps to ZuCo1.0 as EEG2Text paper (Wang and Ji, 2022).
+In BSLR and TSLR calculation, the number of four different seeds are set as $1,2,3,4$ respectively.
 In signal reconstruction task for encoder of UniCoRN, the batch size of EEG and fMRI data is 512 and 320 respectively.
@@
 UniCoRN model is trained with the same settings as fMRI data.
-Appendix B Our Dataset Splitting Method
-In this part, we release the pseudo-code of two dataset splitting methods for EEG and fMRI signal.
+Appendix B Details of Cross-Subject Splitting
+For fMRI dataset, we consider a bipartite graph $\mathcal{G}_{1}=(\mathcal{U},\mathcal{V},\mathcal{E})$ where $\mathcal{U}=\{S_{i}\}_{i=1}^{N}$, $\mathcal{V}=\{M_{k}\}_{k=1}^{M}$.
+$\mathcal{E}$ is the edge between node in $\mathcal{U}$ and node in $\mathcal{V}$, indicating $\left\langle S_{i},M_{k}\right\rangle$ pair in the dataset.
+$N$ is the total number of subjects and $M$ is the total number of tasks.
+We assert $M<N$, so $e=(u,v)\in\mathcal{E}$ exists for every $v\in\mathcal{V}$, as each text stimuli is listened or read by at least one subject.
+As shown in step 2 of Figure 3, first we pick one edge for each node $v\in\mathcal{V}$ and build a new bipartite graph $\mathcal{G}_{2}=(\mathcal{U},\mathcal{V},\mathcal{E}^{\prime})$.
+Then following step 3, we split graph $\mathcal{G}_{2}$ by subject $\mathcal{U}$ with the given splitting ratio and form three disjoint graphs $\mathcal{G}_{train},\mathcal{G}_{val},\mathcal{G}_{test}$.
+In step 4, some edges satisfying zero data leakage condition are not included in the graph.
+We add these edges to corresponding graphs, extending each graph $\mathcal{G}_{train},\mathcal{G}_{val},\mathcal{G}_{test}$ to its maximally scalable state and finishing the dataset splitting process.
+We also release the pseudo-code of two dataset splitting methods for EEG and fMRI signal.
 As shown in Figure 3, our proposed dataset splitting method consists of four steps.
@@
 In Algorithm 2 where $|\mathcal{V}|<|\mathcal{U}|$, the adjacency matrix is initialized as $N\times K$.
-All hypotheses are based on analysis of cognitive datasets.
+All assertions are based on analysis of cognitive datasets.
 One more thing to notice is that in Line 14 of both pseudo-code, the loop indicates extending training set, validation set, and test set respectively.

### v4

@@
-Cross-Subject Data Splitting for Brain-to-Text Decoding
+Rethinking Cross-Subject Data Splitting for Brain-to-Text Decoding
 Abstract
-Recent major milestones have successfully decoded non-invasive brain signals (e.g.
-functional Magnetic Resonance Imaging (fMRI) and electroencephalogram (EEG)) into natural language.
-Despite the progress in model design, how to split the datasets for training, validating, and testing still remains a matter of debate.
-Most of the prior researches applied subject-specific data splitting, where the decoding model is trained and evaluated per subject.
-Such splitting method poses challenges to the utilization efficiency of dataset as well as the generalization of models.
-In this study, we propose a cross-subject data splitting criterion for brain-to-text decoding on various types of cognitive dataset (fMRI, EEG), aiming to maximize dataset utilization and improve model generalization.
-We undertake a comprehensive analysis on existing cross-subject data splitting strategies and prove that all these methods suffer from data leakage, namely the leakage of test data to training set, which significantly leads to overfitting and overestimation of decoding models.
-The proposed cross-subject splitting method successfully addresses the data leakage problem and we re-evaluate some SOTA brain-to-text decoding models as baselines for further research.
+Recent major milestones have successfully reconstructed natural language from non-invasive brain signals (e.g.
+functional Magnetic Resonance Imaging (fMRI) and Electroencephalogram (EEG)) across subjects.
+However, we find current dataset splitting strategies for cross-subject brain-to-text decoding are wrong.
+Specifically, we first demonstrate that all current splitting methods suffer from data leakage problem, which refers to the leakage of validation and test data into training set, resulting in significant overfitting and overestimation of decoding models.
+In this study, we develop a right cross-subject data splitting criterion without data leakage for decoding fMRI and EEG signal to text.
+Some SOTA brain-to-text decoding models are re-evaluated correctly with the proposed criterion for further research.
+11footnotetext: Work done during internship in JD.com.22footnotetext: Corresponding author.
 1 Introduction
@@
 Recent studies (Makin et al., 2020; Wang and Ji, 2022; Xi et al., 2023; Tang et al., 2023; Duan et al., 2024) have successfully decoded non-invasive brain signals (e.g.
-fMRI, EEG) to text by applying deep computational neural networks.
-However, no consensus has reached on how to split the cognitive dataset for training, validating, and testing.
-Most of the prior work (Ye et al., 2023; Tang et al., 2023) performed subject-specific data splitting for training and evaluating decoding models.
-Under this splitting rule, data for training, validating, and testing all comes from one specific subject from the cognitive dataset.
-For example, Tang et al.
-(2023) picked three subjects out of seven and conducted model training and evaluation on the three subjects respectively.
+fMRI, EEG) to text by applying deep neural networks.
+Most of these works perform within-subject data splitting for training and evaluating decoding models.
 This subject-specific splitting method causes two main problems.
-First, it only utilizes a tiny part of the whole dataset.
-Since the collection of brain signals is costly and time-consuming, such splitting method results in significant waste of data resources.
-Second, it leads to the poor generalization of decoding models.
-As every subject’s brain has unique functional and anatomical structures, subject-specific models may exhibit considerable variability across individuals and fail to generalize to other subjects.
+First, it only uses data from one subject of the whole dataset for training and testing.
+Since brain signal collection is costly and time-consuming, such splitting method results in a significant waste of data resources.
+Second, it leads to poor model generalization.
+As every brain has unique functional and anatomical structures, subject-specific models may exhibit considerable variability across individuals and fail to generalize to other subjects (Liu et al., 2024).
 Moreover, decoding models trained from scratch on limited data are prone to facing the overfitting problem.
-Figure 1: Illustration of naturalistic language comprehension dataset for brain-to-text decoding. Path with the same color indicates one sample for training/validating/testing.
-Some studies (Wang and Ji, 2022; Xi et al., 2023) began to shed light on cross-subject data splitting, which views all the subjects’ data as a whole and performs splitting according to a given ratio (e.g.
-8:1:1 for the training set, validation set, and test set).
-Cross-subject data splitting effectively compensates for the shortcomings of subject-specific splitting and has been widely applied in brain-to-image decoding (Wang et al., 2024; Liu et al., 2024).
-However, unlike datasets for brain-to-image decoding, where subjects are guided to see different and unrepeated pictures, different subjects will hear the same story in the naturalistic language comprehension dataset for brain-to-text decoding, which challenges cross-subject data splitting.
-As shown in Figure 1, such dataset is usually formatted in subject-task-text-signal ($S$-$M$-$T$-$F$) pair, indicating the brain signal $F$ of subject $S$ stimulated by hearing text $T$ from task $M$.
-Current cross-subject data splitting methods (Wang and Ji, 2022; Xi et al., 2023) can be summarized as five categories: (1) split by subjects $S$, (2) split by tasks $M$, (3) split by randomly picking signal frames $F$, (4) split by randomly picking signal frames under certain task $F$-$M$, (5) split by randomly picking consecutive signal frames under certain task $F$-$M$.
-However, based on our observations, all these splitting methods suffer from data leakage problem, namely part of the test data is mixed into the training set, which leads to overfitting in model training and overestimation in model evaluation.
-Specifically, modern brain-to-text decoding models follow an encoder-decoder manner.
-We pick two representative models: EEG2Text (Wang and Ji, 2022) and UniCoRN (Xi et al., 2023) for investigating the damage of data leakage.
-The former is an end-to-end encoder-decoder framework, while the latter first pre-trains the encoder and then applies it in the decoder training.
-Experiments support that data leakage affects model training on both the encoder side and decoder side.
-For the encoder component, if subjects’ brain signals in the test set are mixed into the training set, the encoder will become overfitting and fail to well represent unseen subjects’ brain signals.
-As to the decoder, the situation gets worse if text stimuli are leaked.
-Since the decoder follows an auto-regressive manner and generates token by token, data leakage will cause the decoder to memorize seen paragraphs during the teacher-forcing training stage, which leads to poor generalization to unseen text.
-To avoid data leakage and fairly evaluate the performance of brain-to-text decoding models, we propose a cross-subject data splitting criterion.
+Human brain responds similarly to the same stimuli, despite the individual discrepancy (Hasson et al., 2004; Pereira et al., 2018).
+Therefore, some studies (Wang and Ji, 2022; Xi et al., 2023; Duan et al., 2024) begin to shed light on cross-subject brain-to-text decoding, which performs data splitting based on all the subjects, trains and evaluates decoding model once.
+Cross-subject data splitting effectively compensates for the shortcomings of subject-specific splitting, and has been widely applied in brain-to-image decoding (Wang et al., 2024; Liu et al., 2024).
+However, unlike datasets for brain-to-image decoding (Allen et al., 2022; Chang et al., 2019) where subjects are guided to see different and unrepeated pictures, different subjects will be stimulated by the same story in common naturalistic language comprehension dataset, which challenges cross-subject data splitting.
+Based on our observations, current cross-subject data splitting methods for brain-to-text decoding are wrong because data for validation and test leaks into the training set, rendering the evaluation of the decoding process meaningless.
+Specifically, we find two types of data leakage: brain signal leakage and text stimuli leakage.
+Brain signal leakage refers to test subject’s brain signal appears in training set.
+Text stimuli leakage refers to text in test set appears in the training set.
+Modern brain-to-text decoding models follow an encoder-decoder manner.
+We pick two representative models: EEG2Text (Wang and Ji, 2022) and UniCoRN (Xi et al., 2023) to reveal data leakage and its damage.
+Experiments support that data leakage affects model training on both encoder and decoder side.
+For the encoder, it will become overfitting and fail to well represent brain signals if brain signal leakage exists.
+For the decoder, the situation gets worse if text stimuli leakage happens.
+Any data leakage would cause the auto-regressive decoder to memorize previously seen paragraphs during training stage, resulting in poor generalization to unseen text.
+To avoid data leakage and fairly evaluate the performance of cross-subject brain-to-text decoding models, we propose a right data splitting method.
 We focus on fMRI and EEG signals in this study, although the proposed criterion could be applied to any datasets satisfying the prescribed format.
-In the proposed method, the dataset is split according to subject-text ($S$-$T$) pairs with the following rules: (1) Brain signals collected from specific subject in validation set and test set will not appear in training set, which means the trained encoder cannot get access to any brain information belonging to subjects in test set.
+In the proposed method, we follow two basic rules: (1) Brain signals collected from specific subject in validation set and test set will not appear in training set, which means the trained encoder cannot get access to any brain information belonging to subjects in test set.
 (2) Text stimuli in validation set and test set will not appear in training set.
@@
 Our contributions can be summarized as follows:
-  • To the best of our knowledge, we propose the first cross-subject data splitting criterion for brain-to-text decoding.
-  • We comprehensively analyze current cross-subject data splitting methods and find that all existing methods suffer from data leakage problem, which severely affects the training and evaluation of decoding models.
-  • Some SOTA brain-to-text decoding models are re-evaluated under the proposed cross-subject data splitting method as baselines for further research.
-Figure 2: Different splitting methods for cognitive dataset. (Color printing is preferred.)
-2 Related Work
+  • To the best of our knowledge, we are the first to identify the issue of data leakage in current cross-subject data splitting methods for brain-to-text decoding.
+  • We define the splitting criterion for cross-subject brain-to-text decoding, and propose a right dataset splitting method.
+  • Some SOTA brain-to-text decoding models are re-evaluated using the proposed cross-subject data splitting method to ensure a fair assessment of their performance.
+2 Preliminary
+2.1 Dataset Description
+A naturalistic language comprehension dataset $\mathcal{D}$ contains brain signals of $N$ subjects when they passively listen to $K$ spoken stories.
+Suppose that not all subjects are stimulated by all stories, and different subjects may hear the same story.
+Formally, $S_{1},S_{2},\ldots,S_{N}$ denotes to the $N$ subjects and $M_{1},M_{2},\ldots,M_{K}$ denotes to the $K$ stories in dataset.
+The $k$-th story $M_{k}$ consists of $l_{k}$ text segments $T_{k1},T_{k2},\ldots,T_{kl_{k}}$.
+If the $i$-th subject $S_{i}$ hears the $j$-th text segment $T_{kj}$, then his brain signal is denoted as $F_{ijk}$.
+Figure 1: Illustration of how to build graph to describe dataset step by step.
+2.2 Use Graph to Describe Dataset
+We use multigraph and $k$-partite graph (detailed in D.1 & D.2) to describe the intricate structure of naturalistic language comprehension dataset.
+Definition 2.1.
+A naturalistic language comprehension dataset $\mathcal{D}$ can be represented via a directed $4$-partite multigraph $\mathcal{G}_{\mathcal{D}}$.
+How to build the directed $4$-partite multigraph $\mathcal{G}_{\mathcal{D}}$ step by step is shown in Figure 1.
+Graph 1 is a $2$-partite graph indicating subject $S_{i}$ listening to story $M_{k}$.
+Subject $S_{i}$ and story $M_{k}$ are viewed as vertices, and edges connecting them indicate certain type of relationship (e.g.
+$S_{i}$ “listen to” $M_{k}$ in this case).
+Graph 2 illustrates that story $M_{k}$ consists of text segments $T_{kj}$.
+Graph 3 shows the brain signals $F_{ijk}$ of subject $S_{i}$ stimulated by text segment $T_{kj}$.
+Graph 4 is an example of combining the three $2$-partite graphs Graph 1-3: $F_{122},F_{222},F_{322},F_{422}$ are brain signals of $S_{1},S_{2},S_{3},S_{4}$ stimulated by text segment $T_{22}$ from story $M_{2}$.
+In this example, four edges between $M_{2}$ and $T_{22}$ correspond to the different responses of four subjects to the same text segment.
+There are three edges between $S_{2}$ and $M_{2}$ because $M_{2}$ contains three text segments.
+Edges of the same color indicate one sample in dataset.
+Graph 5 shows the complete directed $4$-partite multigraph $\mathcal{G}_{\mathcal{D}}$ for representing whole dataset.
+Every sample in dataset can be represented through ordered subject-story-text-brain ($S_{i},M_{k},T_{kj},F_{ijk}$) pair.
+We introduce the formal notation of $\mathcal{G}_{\mathcal{D}}$:
+Figure 2: Different splitting methods for cross-subject brain-to-text decoding. (Color printing is preferred.)
+Notation 2.2.
+$\mathcal{G}_{\mathcal{D}}=(\mathcal{V},\mathcal{E},f)$, where $\mathcal{V}=\mathcal{S}\cup\mathcal{M}\cup\mathcal{T}\cup\mathcal{F}$, $\mathcal{S}=\{S_{i}\}_{i=1}^{N}$, $\mathcal{M}=\{M_{k}\}_{k=1}^{K}$, $\mathcal{T}=\{T_{kj}\}_{k,j=1}^{K,l_{k}}$, $\mathcal{F}=\{F_{ijk}\}_{i,j,k=1}^{N,l_{k},K}$ denote subject set, story set, text segment set, and brain signal set.
+$f:\mathcal{E}\rightarrow\mathcal{V}\otimes\mathcal{V}$ is an incidence function that maps each edge to a pair of vertices.
+2.3 Brain-to-Text Decoding Task
+The brain-to-text decoding task seeks to build a decoding model that reconstructs natural language text from brain signals, with the goal of accurately decoding what the subject hears.
+Take fMRI and EEG signal for example.
+fMRI captures brain responses at second level whereas EEG samples brain activity at millisecond level.
+So the pre-processing for fMRI and EEG input varies.
+Previous practice in fMRI-to-text decoding (Tang et al., 2023; Xi et al., 2023) concatenated $L$ future fMRI frames with text segments to form one sample:
+$T_{k,j}^{*}$ $=concat(T_{k,j},T_{k,j+1},\ldots,T_{k,j+L}),$ (1)
+$F_{i,j,k}^{*}$ $=concat(F_{i,j,k},F_{i,j+1,k},\ldots,F_{i,j+L,k}).$ (2)
+In this case, one $(S_{i},M_{k},T_{kj},F_{ijk})$ pair in graph $\mathcal{G}_{\mathcal{D}}$ only represents the start point of one sample, while $(S_{i},M_{k},T_{kj}^{*},F_{ijk}^{*})$ indicates the whole sample.
+In EEG-to-text decoding, previous methods sampled continuous EEG signal $F_{ijk}$ that corresponds to text $T_{kj}$.
+So one $(S_{i},M_{k},T_{kj},F_{ijk})$ pair is viewed as one sample in our definition.
+3 Methodology
+3.1 Cross-Subject Data Splitting Criterion
+Consistent with cross-subject brain-to-image decoding (Wang et al., 2024; Liu et al., 2024), the dataset splitting should obey two basic principles: (1) If brain signal $F_{ijk}$ appears in test set, then any brain signal $F_{i*k}$ belonging to this subject $i$ should not appear in training set.
+(2) If text segment $T_{kj}$ appears in test set, then it should not appear in training set.
+Consistent with the definitions in Section 2, graph $\mathcal{G}_{\mathcal{D}}$ is applied to describe data splitting.
+Since the validation samples are split in the same manner as the test samples, we focus solely on the test set.
+Therefore, we have $\mathcal{G}_{\mathcal{D}}=\mathcal{G}_{train}\cup\mathcal{G}_{test}$.
+3.2 Analysis of Current Splitting Methods
+Edges with different colors are used to represent their classification as either part of the training set or the test set.
+As shown in Figure 2, $(S_{i},M_{k},T_{kj},F_{ijk})$ pairs with green edges indicate training samples, and those with orange edges are test samples.
+Current cross-subject data splitting methods (Wang and Ji, 2022; Xi et al., 2023) can be summarized as five categories:
+  • Method (a): Split subjects $\mathcal{S}$ randomly with given ratio.
+    $\mathcal{G}_{train}=\{(S_{i},M_{k},T_{kj},F_{ijk})|$ (3)
+    $\forall(S_{i}^{\prime},M_{k}^{\prime},T_{kj}^{\prime},F_{ijk}^{\prime})\in\mathcal{G}_{test},S_{i}\neq S_{i}^{\prime}\}$
+  • Method (b): Split stories $\mathcal{M}$ randomly with given ratio.
+    $\mathcal{G}_{train}=\{(S_{i},M_{k},T_{kj},F_{ijk})|$ (4)
+    $\forall(S_{i}^{\prime},M_{k}^{\prime},T_{kj}^{\prime},F_{ijk}^{\prime})\in\mathcal{G}_{test},M_{k}\neq M_{k}^{\prime}\}$
+  • Method (c): Split all the brain signals $\mathcal{F}$ randomly with given ratio.
+    $\mathcal{G}_{train}=\{(S_{i},M_{k},T_{kj},F_{ijk})|$ (5)
+    $\forall(S_{i}^{\prime},M_{k}^{\prime},T_{kj}^{\prime},F_{ijk}^{\prime})\in\mathcal{G}_{test},F_{ijk}\neq F_{ijk}^{\prime}\}$
+  • Method (d): Different from Method (c), it splits brain signals under each story randomly with given ratio, and union them to form the whole training and test set.
+  • Method (e): Different from Method (d), it splits continuous brain signals under each story with given ratio, and union them to form the whole training and test set.
+To facilitate a thorough analysis of data leakage, we introduce the concept of brain signal leakage and text stimuli leakage.
+Brain signal leakage refers to test subject’s brain signal appears in training set.
+Text stimuli leakage refers to text segment in test set appears in the training set.
+Formal definitions of two types of data leakage are given.
+Definition 3.1.
+Brain signal leakage happens when
+$\forall(S_{i},M_{k},T_{kj},F_{ijk})\in\mathcal{G}_{train},$ (6)
+$\exists(S_{i}^{\prime},M_{k}^{\prime},T_{kj}^{\prime},F_{ijk}^{\prime})\in\mathcal{G}_{test},S_{i}^{\prime}=S_{i}.$
+Definition 3.2.
+Text stimuli leakage happens when
+$\forall(S_{i},M_{k},T_{kj},F_{ijk})\in\mathcal{G}_{train},$ (7)
+$\exists(S_{i}^{\prime},M_{k}^{\prime},T_{kj}^{\prime},F_{ijk}^{\prime})\in\mathcal{G}_{test},T_{kj}^{\prime}=T_{kj}.$
+Data leakage can be directly identified in graph $\mathcal{G}_{\mathcal{D}}$.
+As shown in Figure 2, if edges connected to $S_{i}$ are of different colors, it indicates that brain signals of $S_{i}$ appears in both training set and test set, which leads to brain signal leakage.
+Similarly, if edges connected to $T_{kj}$ are of different colors, it suggests that text segment $T_{kj}$ appears in both training set and test set, which leads to text stimuli leakage.
+fMRI / EEG Method(a) Method(b) Method(c) Method(d) Method(e)
+Brain Signal Leakage ✗ / ✗ ✓/ ✓ ✓/ ✓ ✓/ ✓ ✓/ NA
+Text Stimuli Leakage ✓/ ✓ ✗ / ✗ ✓/ ✓ ✓/ ✓ ✓/ NA
+Table 1: Data leakage in five different splitting methods applied to fMRI and EEG to text decoding separately.
+As a result, in the scenario of EEG signals where $(S_{i},M_{k},T_{kj},F_{ijk})$ is viewed as a sample: Method (a) suffers from text stimuli leakage.
+Method (b) faces brain signal leakage.
+Method (c) is affected by leakage of both text stimuli and brain signals.
+Method (d) and (e) do not show any differences compared to method (c) in EEG-to-text decoding.
+In fMRI-to-text decoding, continuous fMRI frames and text stimuli are concatenated to form one sample.
+$(S_{i},M_{k},T_{kj},F_{ijk})$ indicates the start point of one sample instead of the whole sample (recall Section 2.3).
+In this case, method (d) and (e) are different.
+Similar to method (c), method (d) and (e) face both brain signal leakage and text stimuli leakage.
+But for method (e) the text stimuli is slight.
+It only happens in the overlapping part between training samples and test samples.
+The situations of data leakage in different splitting methods are detailed in Table 1.
+Figure 3: General frameworks of current brain-to-text decoding models and how data leakage affect them.
+Figure 4: The detailed steps of our proposed cross-subject data splitting method. (Color printing is preferred.)
+3.3 Frameworks of Current Decoding Models and How Data Leakage Affect Them
+Current brain-to-text decoding models adopt an encoder-decoder framework, where the encoder is responsible for converting brain signals into low-dimensional representations and the decoder (usually Transformer-based) learns to map these representations to natural language.
+Two representative models EEG2Text (Wang and Ji, 2022) and UniCoRN (Xi et al., 2023) are selected for investigating the affection of data leakage.
+As shown in Figure 3(a), EEG2Text applies an end-to-end training manner.
+EEG feature sequence is first extracted by a multi-layer Transformer encoder and then converted to natural language with a pretrained BART (Lewis et al., 2020).
+UniCoRN provides a unified framework for EEG and fMRI to text decoding.
+It follows a two-stage training manner as shown in Figure 3(b).
+Take fMRI-to-text decoding for example, the encoder is first pre-trained with a brain signal reconstruction task to capture spatial and temporal feature via a 3D-CNN and multi-layer Transformer encoder module.
+Then BART (Lewis et al., 2020) is fine-tuned to translate fMRI representation into natural language.
+Figure 3 illustrates how brain signal leakage and text stimuli leakage affect current encoder-decoder based brain-to-text decoding framework.
+For the encoder component, if subjects’ brain signals in test set are mixed into training set, the encoder will become overfitted and fail to well represent unseen subjects’ brain signals.
+For the decoder component, since it generates token by token in an auto-regressive manner, data leakage will cause the decoder to memorize seen text during the teacher-forcing training stage.
+The decoder will predict next token regardless of encoded brain signals.
+3.4 A Right Cross-Subject Splitting Method
+A right cross-subject splitting method is proposed to eliminate both brain signal leakage and text stimuli leakage.
+$\mathcal{G}_{train}=\{(S_{i},M_{k},T_{kj},F_{ijk})|\forall(S_{i}^{\prime},M_{k}^{\prime},$ (8)
+$T_{kj}^{\prime},F_{ijk}^{\prime})\in\mathcal{G}_{test},S_{i}\neq S_{i}^{\prime},T_{kj}\neq T_{kj}^{\prime}\}.$
+Given the differences of EEG and fMRI dataset, we address them separately and propose two data splitting methods.
+In EEG dataset, $(S_{i},M_{k},T_{kj},F_{ijk})$ forms one sample.
+As shown in Figure 4, our proposed splitting method consists of three steps:
+  • Step 1: Select $\sum_{k=1}^{K}l_{k}$ samples from $\mathcal{G}_{\mathcal{D}}$ and form a new graph $\mathcal{G}_{\mathcal{D}}^{\prime}$ that satisfies
+    $\forall(S_{i}^{\prime},M_{k}^{\prime},T_{kj}^{\prime},F_{ijk}^{\prime}),(S_{i}^{\prime\prime},M_{k}^{\prime\prime},T_{kj}^{\prime\prime},F_{ijk}^{\prime\prime})$ (9)
+    $\in\mathcal{G}_{\mathcal{D}}^{\prime},T_{kj}^{\prime}\neq T_{kj}^{\prime\prime}.$
+  • Step 2: Split $\mathcal{G}_{\mathcal{D}}^{\prime}$ to $\mathcal{G}_{train}^{\prime}$ and $\mathcal{G}_{test}^{\prime}$ with a given ratio. The splitting should follow
+    $\mathcal{G}_{train}^{\prime}=\{(S_{i},M_{k},T_{kj},F_{ijk})|$ (10)
+    $\forall(S_{i}^{\prime},M_{k}^{\prime},T_{kj}^{\prime},F_{ijk}^{\prime})\in\mathcal{G}_{test}^{\prime},S_{i}\neq S_{i}^{\prime}\},$
+    $\mathcal{G}_{test}^{\prime}=\{(S_{i},M_{k},T_{kj},F_{ijk})|$ (11)
+    $\forall(S_{i}^{\prime},M_{k}^{\prime},T_{kj}^{\prime},F_{ijk}^{\prime})\in\mathcal{G}_{train}^{\prime},S_{i}\neq S_{i}^{\prime}\}.$
+  • Step 3: Expand $\mathcal{G}_{train}^{\prime}$ and $\mathcal{G}_{test}^{\prime}$ with $\mathcal{G}_{train\_exp}^{\prime}$ and $\mathcal{G}_{test\_exp}^{\prime}$ separately.
+    $\mathcal{G}_{train}^{\prime}\leftarrow\mathcal{G}_{train}^{\prime}\cup\mathcal{G}_{train\_exp}^{\prime}$ (12)
+    $\mathcal{G}_{test}^{\prime}\leftarrow\mathcal{G}_{test}^{\prime}\cup\mathcal{G}_{test\_exp}^{\prime}$
+    where $\mathcal{G}_{train\_exp}^{\prime}$ and $\mathcal{G}_{test\_exp}^{\prime}$ are
+    $\mathcal{G}_{train\_exp}^{\prime}$ $=\{(S_{i},M_{k},T_{kj},F_{ijk})\in\mathcal{G}_{\mathcal{D}}|$ (13)
+    $S_{i}\in\mathcal{S}_{train}^{\prime},T_{kj}\in\mathcal{T}_{train}^{\prime}\},$
+    $\mathcal{G}_{test\_exp}^{\prime}$ $=\{(S_{i},M_{k},T_{kj},F_{ijk})\in\mathcal{G}_{\mathcal{D}}|$ (14)
+    $S_{i}\in\mathcal{S}_{test}^{\prime},T_{kj}\in\mathcal{T}_{test}^{\prime}\}.$
+$\mathcal{S}_{train}^{\prime}$, $\mathcal{T}_{train}^{\prime}$, $\mathcal{S}_{test}^{\prime}$, $\mathcal{T}_{test}^{\prime}$ indicate subject set, text segment set in $\mathcal{G}_{train}^{\prime}$ and subject set, text segment set in $\mathcal{G}_{test}^{\prime}$ respectively.
+Some samples are discarded in our proposed splitting method, i.e.
+$\mathcal{G}_{\mathcal{D}}\neq\mathcal{G}_{train}^{\prime}\cup\mathcal{G}_{test}^{\prime}$.
+In Appendix D, we demonstrate that it is unavoidable for some samples to be discarded in order to satisfy the cross-subject data splitting criterion.
+To fMRI dataset, continuous text segments and brain signals are concatenated to form one sample $(S_{i},M_{k},T_{kj}^{*},F_{ijk}^{*})$.
+If we follow the same splitting method as to EEG dataset, text stimuli leakage will happen in the overlapping part of two samples, when one sample is assigned to training set and the other is assigned to validation or test set.
+We propose a simple solution that achieves the balance between discarding as little data as possible while ensuring zero data leakage: Step 1 and Step 3 remain the same as splitting method for EEG dataset.
+In Step 2, $\mathcal{G}_{train}^{\prime}$ and $\mathcal{G}_{test}^{\prime}$ should follow
+$\mathcal{G}_{train}^{\prime}=\{(S_{i},M_{k},T_{kj},F_{ijk})|\forall(S_{i}^{\prime},M_{k}^{\prime},$ (15)
+$T_{kj}^{\prime},F_{ijk}^{\prime})\in\mathcal{G}_{test}^{\prime},S_{i}\neq S_{i}^{\prime},M_{k}\neq M_{k}^{\prime}\},$
+$\mathcal{G}_{test}^{\prime}=\{(S_{i},M_{k},T_{kj},F_{ijk})|\forall(S_{i}^{\prime},M_{k}^{\prime},$ (16)
+$T_{kj}^{\prime},F_{ijk}^{\prime})\in\mathcal{G}_{train}^{\prime},S_{i}\neq S_{i}^{\prime},M_{k}\neq M_{k}^{\prime}\}.$
+4 Experimental Settings
+4.1 Implementation Detail
+We test two SOTA cross-subject brain-to-text decoding models UniCoRN (Xi et al., 2023) and EEG2Text (Wang and Ji, 2022) on fMRI dataset Narratives (Nastase et al., 2021) and EEG dataset ZuCo (Hollenstein et al., 2018).
+Because the number of stories in ZuCo dataset is too small, and method (e) makes no difference to EEG as method (d), we only consider splitting method (a), (c), (d) for EEG.
+We follow the same settings of UniCoRN and EEG2Text, except all the datasets are split to the ratio of 8:1:1 for fair comparison.
+Details are shown in Appendix B.
+4.2 Evaluation Metrics
+Data Leakage Metrics
+We design two novel evaluation metrics Brain Signal Leakage Rate (BSLR) and Text Stimuli Leakage Rate (TSLR) to quantify two types of data leakage.
+Note that the situation for validation set is the same as test set, so we only consider test set in experiments.
+BSLR indicates the average percentage of each subject’s brain signals in test set appearing in training set, which could be formulated as
+$\frac{1}{N_{test}}\sum_{i=1}^{N_{test}}\min(1,\frac{|\{F_{ijk}|F_{ijk}\in(\mathcal{G}_{test}\cap\mathcal{G}_{train})\}|}{|\{F_{ijk}|F_{ijk}\in\mathcal{G}_{train}\}|})$ (17)
+where $N_{test}$ stands for the total number of subjects in test set.
+$|\cdot|$ stands for the cardinality of a set.
+Function $\min(\cdot,\cdot)$ is applied to make sure for each subject the data leakage rate is less than one.
+The definition of TSLR is different for EEG signal and fMRI signal.
+Since $(S_{i},M_{k},T_{kj},F_{ijk})$ indicates one sample in EEG dataset, definition of TSLR for EEG dataset is similar to BSLR, which measures the average percentage of certain text in test set appearing in training set.
+$\frac{1}{M_{test}}\sum_{j=1}^{M_{test}}\min(1,\frac{|\{T_{kj}|T_{kj}\in(\mathcal{G}_{test}\cap\mathcal{G}_{train})\}|}{|\{T_{kj}|T_{kj}\in\mathcal{G}_{train}\}|})$ (18)
+where $M_{test}$ stands for the total number of text segments in test set.
+Type Method Narratives ZuCo
+BSLR(%)($\%$) (a) 0.00±0.00 0.00±0.00
+(b) 9.67±4.80 /
+(c) 12.50±0.04 12.50±0.03
+(d) 12.80±0.01 12.59±0.02
+(e) 12.27±0.01 /
+(f) 0.00±0.00 0.00±0.00
+TSLR(%)($\%$) (a) 100.00±0.00 22.50±1.31
+(b) 0.00±0.00 /
+(c) 100.00±0.00 13.07±0.11
+(d) 99.82±0.17 12.88±0.04
+(e) 9.29±0.06 /
+(f) 0.00±0.00 0.00±0.00
+Table 2: Results of Brain Signal Leakage Rate (BSLR) and Text Stimuli Leakage Rate (TSLR). Lower is better.
+Dataset Model Method Original Test Set / Additional Test Set
+BLEU-1 BLEU-2 BLEU-3 ROUGE1-F
+Narratives UniCoRN (a) 49.56 / 18.43 30.49 / 1.25 21.07 / 0.00 40.65 / 16.38
+(b) 26.37 / 23.31 7.50 / 5.79 2.48 / 1.44 19.62 / 18.74
+(c) 50.24 / 16.96 30.83 / 0.09 21.23 / 0.00 41.01 / 15.12
+(d) 49.63 / 17.20 30.29 / 1.15 20.85 / 0.00 41.03 / 15.83
+(e) 28.94 / 21.79 9.39 / 4.62 4.07 / 1.19 19.49 / 18.78
+(f) 22.83 / 21.64 5.69 / 4.97 1.43 / 1.28 19.04 / 18.45
+ZuCo UniCoRN (a) 58.09 / 18.54 49.23 / 1.31 43.23 / 0.00 67.50 / 15.39
+(c) 52.30 / 18.38 42.89 / 1.03 36.80 / 0.00 67.29 / 15.25
+(d) 50.02 / 19.84 43.53 / 1.20 32.71 / 0.03 67.33 / 15.12
+(f) 23.32 / 22.89 7.78 / 7.46 3.01 / 2.75 17.92 / 17.63
+EEG2Text (a) 51.22 / 17.41 33.83 / 1.04 22.99 / 0.00 46.58 / 15.92
+(c) 53.83 / 17.38 38.99 / 0.84 29.57 / 0.00 53.56 / 16.07
+(d) 53.92/ 16.86 41.06 / 1.32 23.12 / 0.00 49.38 / 15.83
+(f) 24.49 / 23.71 7.49 / 7.42 2.28 / 2.33 25.74 / 23.30
+Table 3: Performance of brain-to-text decoding models under different splitting methods on original test set and an additional test set.
+The green mark and red mark denotes a method without and with text stimuli leakage.
+To fMRI dataset, continuous fMRI frames with corresponding text segments are concatenated as one sample.
+As a result, TSLR for fMRI signal is considered as the average percentage of the same text segments in test set appearing in training set, which is
+$\frac{1}{M_{test}}\sum_{j=1}^{M_{test}}\tau\frac{|\{T_{kj}|T_{kj}\in(\mathcal{G}_{test}\cap\mathcal{G}_{train})\}|}{|\mathcal{G}_{test}|\times L}$ (19)
+where $\tau=0$ if $\{T_{kj}|T_{kj}\in\mathcal{G}_{test}\cap\mathcal{G}_{train}\}=\emptyset$ else
+$\tau=\min(1,\frac{|\{T_{kj}|T_{kj}\in\mathcal{G}_{train}\}|}{|\{T_{kj}|T_{kj}\in(\mathcal{G}_{test}\cap\mathcal{G}_{train}\})|}).$ (20)
+Decoding Performance Metrics
+BLEU (Papineni et al., 2002) and ROUGE (Lin, 2004) are applied to measure the decoding performance.
+BLEU measures the n-gram overlap between decoded content and ground truth.
+ROUGE-N comparing the consistency of N-grams between the decoded content and the ground truth.
+5 Experiments and Analysis
+We first quantify the data leakage condition of different methods with BSLR and TSLR metrics.
+Then we demonstrate the damage of data leakage on encoder side and decoder side.
+For model encoder, we analyze its validation loss under different splitting methods.
+For model decoder, three experiment settings are applied: (1) An additional test set that ensures zero data leakage is left out as comparison to original test set.
+(2) The input brain signals are randomly shuffled.
+(3) Training original models with more epochs and smaller learning rate.
+5.1 Verification for Data Leakage
+Experiments on BSLR and TSLR are conducted four times with different seeds.
+The results in Table 2 are consistent with theoretical analysis.
+A value of zero in BSLR and TSLR demonstrate no brain signal leakage and text stimuli leakage, while higher values suggest more significant data leakage issues.
+Notably, only our method (f) prevents both brain signal leakage and text stimuli leakage.
+Figure 5: Validation loss of encoder under different dataset splitting methods in two datasets.
+5.2 Damage of Data Leakage to Encoder
+Dataset Model Method Ordered Input / Shuffled Input
+BLEU-1 BLEU-2 BLEU-3 ROUGE1-F
+Narratives UniCoRN (a) 49.56 / 47.39 30.49 / 28.95 21.07 / 18.40 40.65 / 35.12
+(b) 26.37 / 20.18 7.50 / 3.52 2.48 / 0.51 19.62 / 15.58
+(c) 50.24 / 48.48 30.83 / 30.21 21.23 / 19.39 41.01 / 38.43
+(d) 49.63 / 50.21 30.29 / 32.18 20.85 / 21.46 41.03 / 41.69
+(e) 28.94 / 24.84 9.39 / 6.56 4.07 / 2.04 19.49 / 17.90
+(f) 22.83 / 18.21 5.69 / 2.47 1.43 / 0.22 19.04 / 16.83
+ZuCo UniCoRN (a) 58.09 / 59.23 49.23 / 51.35 43.23 / 44.27 67.50 / 68.93
+(c) 52.30 / 50.24 42.89 / 37.96 36.80 / 30.21 67.29 / 63.43
+(d) 50.02 / 51.12 43.53 / 40.85 32.71 / 28.24 67.33 / 64.88
+(f) 23.32 / 19.38 7.78 / 2.51 3.01 / 0.00 17.92 / 15.21
+EEG2Text (a) 51.22 / 50.63 33.83 / 32.19 22.99 / 20.63 46.58 / 44.70
+(c) 53.83 / 50.33 38.99 / 33.42 29.57 / 23.19 53.56 / 48.78
+(d) 53.92 / 51.46 41.06 / 35.87 23.12 / 24.75 49.38 / 47.42
+(f) 24.49 / 18.72 7.49 / 2.01 2.28 / 0.00 25.74 / 15.36
+Table 4: Performance of brain-to-text decoding models under different splitting methods with ordered brain signals and randomly shuffled brain signals as model input respectively.
+The green mark and red mark denotes a method without and with text stimuli leakage correspondingly.
+Dataset Model BLEU-N ($\%$) ROUGE-1 ($\%$)
+$N=1$ $N=2$ $N=3$ $N=4$ $R$ $P$ $F$
+Narratives UniCoRN 22.83 5.69 1.43 0.48 15.55 24.80 19.04
+ZuCo UniCoRN 23.32 7.78 3.01 1.09 18.47 20.00 17.92
+EEG2Text 24.49 7.49 2.28 0.62 23.98 23.95 25.74
+Table 5: A fair benchmark for evaluating the performance of cross-subject brain-to-text decoding models.
+Evaluating the encoder independently can be challenging in an end-to-end training scenario.
+Therefore, we primarily focus on a pre-trained encoder.
+Validation loss is applied to measure data leakage, as a proper evaluation index of encoder’s representation ability is missing.
+The validation loss of encoder under different data splitting methods is shown in Figure 5.
+For fMRI data, the presence of brain signal leakage causes the validation loss of methods (b), (c), (d), and (e) to continuously decrease even over extended training epochs, which indicates the encoder is actually overfitting and its representation ability is degrading.
+In contrast, with methods (a) and (f) that are not affected by brain signal leakage, the validation loss quickly increases after reaching its minimum within a few epochs.
+For EEG, we find validation loss keeps dropping for all methods even with very long training epochs, regardless of brain signal leakage or not.
+We think the poor spatial resolution of EEG signal might lead to this phenomenon.
+5.3 Damage of Data Leakage to Decoder
+Evaluation on Additional Test Set
+An additional test set that ensures zero data leakage is left out to evaluate the actual performance of brain-to-text decoding models.
+If the original test set is correctly split, its decoding result should be similar to that of the additional test set.
+From Table 3, we observe that the decoding model tends to overfit when text stimuli leakage occurs, as seen in methods (a), (c), (d), and (e) in Narratives, and methods (a) and (c) in ZuCo.
+The BLEU and ROUGE score is significantly lower in the additional test set.
+While in our proposed splitting method (f), the decoding performance of original and additional test set are similar.
+We also notice that methods with a high Text Stimuli Leakage Rate (TSLR), such as method (a) in Narratives, exhibit more overfitting compared to methods with a low TSLR, like method (e).
+Shuffle Input Brain Signals
+We conduct a chance-level experiment to investigate whether decoding models learn language reconstruction from brain signals.
+Specifically, the input brain signals are randomly shuffled.
+Decoding performance in test set is expected to be very poor if text stimuli leakage does not happen, as the shuffled input is considered as noise.
+However, if text stimuli in test set leaks into training set, the model will simply memorize seen text and the decoding performance is not supposed to be affected.
+Results are presented in Table 4.
+For fMRI, we find the decoding performance of models under splitting method (a), (c), and (d) remain the same no matter the input is ordered or shuffled.
+Similar phenomenon is also observed in EEG dataset when it comes to splitting method (a), (c), (d).
+But in splitting method without text stimuli leakage, model performance with shuffled input drops significantly.
+This experiment demonstrates that the brain-to-text decoding task become meaningless when text stimuli leakage exists, as the Transformer block is capable of generating text that was previously encountered during the training phase.
+Longer Training Epochs with Smaller Learning Rate
+According to fundamental machine learning principle, model performance in test set will first increase and then drop as the training proceeds.
+In this experiment, we try training models under different splitting methods with longer training epochs and smaller learning rate.
+If text stimuli leakage happens, the model is overfitting and its performance is supposed to keep increasing.
+Results and detailed analysis are presented in Appendix E.
+In conclusion, the model’s performance on test set continues to improve when text stimuli leakage happens, confirming that such leakage results in significant overfitting in decoding models.
+5.4 A Fair Benchmark
+We re-evaluate two SOTA models for brain-to-text decoding under our cross-subject data splitting method and release a fair benchmark.
+UniCoRN is tested for both fMRI and EEG decoding, EEG2Text model is tested for EEG decoding.
+The results are listed in Table 5.
+For EEG dataset, UniCoRN achieves higher results in BLEU-2,3,4 while EEG2Text is better in BLEU-1 and ROUGE-1.
+6 Conclusion and Discussion
+In this paper, we evidence that all current dataset splitting methods for cross-subject brain-to-text decoding have data leakage problem through theoretical analysis and experiments.
+Such data leakage leads to model overfitting and largely exaggerates model performance, rendering model evaluation meaningless.
+To fix this issue, we propose a right cross-subject data splitting method.
+Current SOTA models are re-evaluated for further researches.
+It’s essential to realize the false promise of current SOTA methods.
+This inspires future researches to design more general and accurate models for brain-to-text decoding, under the right cross-subject dataset splitting method.
+It should be noted that we don’t seek to propose any modifications to improve current models in the scope of this paper.
+We mainly focus on revealing the false dataset splitting method and its detrimental impact on cross-subject brain-to-text decoding research.
+Limitations
+The limitations of this work include three aspects: (1) Although our splitting method can be applied to any natural language comprehension cognitive dataset, we only analyze cross-subject data splitting methods in fMRI and EEG dataset.
+We leave the investigation of other cognitive signals (e.g.
+ECoG, MEG, etc.) to future work.
+(2) Our proposed dataset splitting method meets the above requirements at the expense of discarding some data in the dataset.
+We recommend future datasets in this domain follow these guidelines.
+The division of the training set, validation set, and test set should be provided when the dataset is released.
+Besides, we suggest hiring new subjects with unique stimuli for the validation set and test set, which is good for testing the generalization ability of models without loss of data.
+(3) During experiments we find existing models rely more on a strong auto-regressive decoder to achieve good generation quality.
+The encoder is of limited use in all SOTA models.
+And we also notice in experiments that the encoder of EEG2Text keeps overfitting whether with or without brain signal leakage.
+We leave it as future research.
+Ethics Statement
+In this paper, we introduce a new dataset splitting method to avoid data leakage for decoding brain signals to text task.
+Experiments are conducted on the publicly accessible cognitive datasets “Narratives” and ZuCo1.0 with the authorization from their respective maintainers.
+Both datasets have been de-identified by dataset providers and used for researches only.
+Acknowledgements
+This research is supported by the National Natural Science Foundation of China (No.62476127), the Natural Science Foundation of Jiangsu Province (No.BK20242039), the Basic Research Program of the Bureau of Science and Technology (ILF24001), the Meituan Research Fund (No.APAP202507180115), the Scientific Research Starting Foundation of Nanjing University of Aeronautics and Astronautics (No.YQR21022), and the High Performance Computing Platform of Nanjing University of Aeronautics and Astronautics.
+Appendix A Related Work
 Brain Signal
@@
 (2023) improved the model design and proposed a unified framework for decoding both fMRI and EEG signals.
-3 Methodology
-In this section, we will first introduce the formal definition of brain-to-text decoding and the general description of dataset format.
-Then we systematically analyze current cross-subject data splitting methods and point out that all existing methods suffer from two kinds of data leakage issues: brain signal leakage and text stimuli leakage.
-Finally, a cross-subject splitting criterion is proposed which avoids the above-mentioned data leakage problems.
-3.1 Task Definition
-Given the brain signal $F_{ij}$ stimulated by $i$-th subject $S_{i}$ hearing or reading certain text $T_{j}$, brain-to-text decoding aims to decode $F_{ij}$ back to text $T_{j}^{\prime}$ and make $T_{j}^{\prime}$ as similar as possible to $T_{j}$.
-The composition of $F_{ij}$ and $T_{j}$ is different as to fMRI and EEG.
-The former samples brain information discretely with a fixed time interval TR, while the latter samples continuously.
-To fMRI, consistent text segments $s_{j}$ with corresponding fMRI frames $f_{ij}$ are concatenated to form a sample pair $\left\langle F_{ij},T_{j}\right\rangle$, where $T_{j}=concat(s_{j},s_{j+1},\ldots,s_{j+L-1})$ and $F_{ij}=concat(f_{ij},f_{i,j+1},\ldots,f_{i,j+L-1})$, and $|T_{j}|=|F_{ij}|=L$.
-To EEG, since signals corresponding to complete text stimuli are available and they are continuous, we bond text $T_{j}$ (i.e.
-text stimuli) and EEG signal $F_{ij}$ together to form a sample pair $\left\langle F_{ij},T_{j}\right\rangle$.
-Under most scenarios, each text stimulus $T_{j}$ belongs to one certain task $M_{k}$.
-So the signal-text pair $\left\langle F_{ij},T_{j}\right\rangle$ can be further split into $\left\langle F_{ij},M_{k}\right\rangle$ and $\left\langle M_{k},T_{kj}\right\rangle$.
-Figure 3: The process of our proposed cross-subject data splitting method. (Color printing is preferred.)
-One of the purposes of cross-subject splitting is to endow models the ability to decode unseen subject’s brain signal.
-As a result, if brain signal $F_{ij}$ appears in test set $S_{test}$, any signal $F_{i*}$ belonging to subject $i$ should not appear in training set $S_{train}$.
-Similarly, text stimuli $T_{kj}$ in $S_{test}$ should not appear in $S_{train}$.
-The above splitting rules for training set can be formulated by Cartesian product:
-$S_{train}=F_{train}\times T_{train},$ (1)
-where
-$F_{train}=\{F_{ij}|i\in I\},$ (2)
-$I=\{i|F_{ij}\notin S_{val},S_{test},\forall j\},$ (3)
-and
-$T_{train}=\{T_{kj}|T_{kj}\notin S_{val},S_{test}\}.$ (4)
-Such rules are also applicable to validation and test set. We omit their displays here for simplicity.
-3.2 Current Splitting Methods
-Current cross-subject data splitting methods can be summarized as five categories according to classifying objectives $S_{i},M_{k},T_{kj},F_{ij}$.
-More specifically, the five dataset splitting methods are characterised as (1) split by subjects $S_{i}$, (2) split by tasks $M_{k}$, (3) split by randomly picking signal frames $F_{ij}$, (4) split by randomly picking signal frames under certain task $F_{ij}$-$M_{k}$, (5) split by randomly picking consecutive signal frames under certain task $F_{ij}$-$M_{k}$, corresponding to (a), (b), (c), (d), (e) in Figure 2.
-Figure 2 vividly displays the differences between current dataset splitting methods.
-In this example, we choose 4 subjects ($S_{1}$ to $S_{4}$) with 3 tasks ($M_{1}$ to $M_{3}$) each containing 4 ($T_{11}$ to $T_{14}$), 3 ($T_{21}$ to $T_{23}$), 4 ($T_{31}$ to $T_{34}$) text stimuli respectively.
-$F_{ij}$ indicates the brain signal of $i$-th subject stimulated by $j$-th text under task $M_{k}$, e.g.
-$F_{21}$ means brain signal of $S_{2}$ hearing $T_{11}$.
-The line connecting two symbols indicates they are related to one sample in dataset.
-Take path $S_{1},M_{1},T_{11},F_{11}$ for example, it indicates that subject $S_{1}$ listens to text stimuli $T_{11}$ belonging to task $M_{1}$ and $S_{1}$’s corresponding brain signal is recorded as $F_{11}$.
-Some symbols are connected with several lines.
-For example, the four lines between $S_{1}$ and $M_{1}$ correspond to $\left\langle M_{1},T_{11}\right\rangle$, $\left\langle M_{1},T_{12}\right\rangle$, $\left\langle M_{1},T_{13}\right\rangle$, $\left\langle M_{1},T_{14}\right\rangle$ counting from left to right.
-Similarly, the three lines between $M_{1}$ and $T_{11}$ correspond to $\left\langle S_{1},M_{1}\right\rangle$, $\left\langle S_{2},M_{1}\right\rangle$, $\left\langle S_{3},M_{1}\right\rangle$ respectively.
-The same rules can be extended to other lines and symbols.
-The green lines and orange lines stand for training samples and testing samples.
-The grey dotted line means the sample is abandoned, which will be introduced in our data splitting method.
-As the splitting of validation set is the same as test set, we only consider training set and test set in this section for simplicity.
-We will use method (a), (b), (c), (d), (e) to represent five current dataset splitting methods in the rest of the paper. Method (a) splits the dataset according to subjects, which can be described as
-$S_{train}=\{\left\langle F_{ij},T_{kj}\right\rangle|S_{i}\notin S_{val},S_{test}\}$ (5)
-for training set. Method (b) splits the dataset according to tasks, which is described as
-$S_{train}=\{\left\langle F_{ij},T_{kj}\right\rangle|M_{k}\notin S_{val},S_{test}\}$ (6)
-for training set. Method (c), (d), and (e) all split the dataset according to brain signal frames
-$S_{train}=\{\left\langle F_{ij},T_{kj}\right\rangle|F_{ij}\notin S_{val},S_{test}\}.$ (7)
-However, there are slight differences between these three methods.
-Method (c) views all the brain signal frames in dataset as a whole and splits according to the default proportion (e.g.
-8:1:1).
-Method (d) views signal frames under certain task $M_{k}$ as a whole and splits proportionally, and then unions all training sets under different tasks to form a complete set for training.
-Method (e) is similar to method (d).
-They both first split training, validation, and test set under certain task proportionally and then union them.
-The difference lies in that method (d) randomly picks signal frames, while method (e) picks consecutive signal frames.
-Type Method Narratives / ZuCo Average
-seed1111 seed2222 seed3333 seed4444
+Appendix B Implementation Details
 […]
-(b) 6.73 / - 6.32 / - 7.7 / - 17.93 / - 9.67 / -
-(c) 12.55 / 12.52 12.52 / 12.55 12.48 / 12.48 12.44 /12.46 12.50 / 12.50
-(d) 12.81 / 12.60 12.8 / 12.58 12.78 / 12.56 12.79 / 12.61 12.795 / 12.59
-(e) 12.28 / - 12.27 / - 12.26 / - 12.27 / - 12.27 / -
-(f) 0.00 / 0.00 0.00 / 0.00 0.00 / 0.00 0.00 / 0.00 0.00 / 0.00
-[…]
-(b) 0.00 / - 0.00 / - 0.00 / - 0.00 / - 0.00 / -
-(c) 100.00 / 13.21 100.00 / 13.06 100.00 / 12.91 100.00 / 13.1 100.00 / 13.07
-(d) 99.93 / 0.00 99.81 / 0.00 99.54 / 0.00 99.99 / 0.00 99.82 / 0.00
-(e) 9.19 / - 9.31 / - 9.36 / - 9.29 / - 9.29 / -
-(f) 0.00 / 0.00 0.00 / 0.00 0.00 / 0.00 0.00 / 0.00 0.00 / 0.00
-Table 1: Results of Brain Signal Leakage Rate (BSLR) and Text Stimuli Leakage Rate (TSLR).
-We first point out the data leakage problems in current splitting methods through the analysis of training set and test set composition.
-Specifically, following the definition in subsection 3.1, two kinds of data leakage, brain signal leakage and text stimuli leakage, are defined.
-The data leakage situation of different methods can be reflected through visualization in Figure 2.
-Lines between $S_{i}$ and $M_{k}$ indicate brain signal leakage situation and lines between $T_{kj}$ and $M_{k}$ indicate text stimuli leakage situation.
-If lines associated with $S_{i}$ or $T_{kj}$ are of different colours, data in test set leaks into training set.
-Remind the composition of samples differs as to fMRI signal and EEG signal, so the dataset splitting methods are different for two kinds of brain signal too.
-Since fMRI signals need to be sampled continuously with a certain length $L$, one sample shown in Figure 2 (e.g.
-$S_{1}$-$M_{1}$-$T_{11}$-$F_{11}$) is actually the first part of one fMRI sample, with $L-1$ continuous brain signal frames following (e.g.
-$F_{12},F_{13},\ldots$).
-In this sense, brain signal leakage doesn’t exist in method (a) for EEG, but method (a) suffers from text stimuli leakage.
-Text stimuli leakage does not exist in method (b) but brain signal leakage exists.
-Method (c) suffers from both brain signal leakage and text stimuli leakage.
-Method (d) and (e) are the same to EEG, with brain signal leakage.
-For fMRI, the situation of data leakage for different methods is similar to EEG, except for method (d) and (e), which are the same for EEG but actually different for fMRI.
-Method (d) suffers from both brain signal leakage and text stimuli leakage while in method (e) text stimuli leakage happens in the overlap between training samples and test samples.
-3.3 Cross-Subject Splitting Criterion
-To eliminate data leakage from both brain signal leakage and text stimuli leakage, we split the dataset by $\left\langle S_{i},T_{j}\right\rangle$ pairs as shown in (f) of Figure 2.
-Since EEG and fMRI are different in the composition of dataset, we treat them separately and propose two data splitting methods.
-As to EEG dataset where $F_{ij}$ and $T_{j}$ form a sample, we consider a bipartite graph $\mathcal{G}_{1}=(\mathcal{U},\mathcal{V},\mathcal{E})$ where $\mathcal{U}=\{S_{i}\}_{i=1}^{N}$, $\mathcal{V}=\{T_{j}\}_{j=1}^{M}$.
-$\mathcal{E}$ is the edge between node in $\mathcal{U}$ and node in $\mathcal{V}$, indicating $\left\langle S_{i},T_{j}\right\rangle$ pair in the dataset.
-$N$ is the total number of subjects and $M$ is the total number of unique text stimuli.
-We assert $M>N$, so $e=(u,v)\in\mathcal{E}$ exists for every $v\in\mathcal{V}$, as each text stimuli is listened or read by at least one subject.
-As shown in step 2 of Figure 3, first we pick one edge for each node $v\in\mathcal{V}$ and build a new bipartite graph $\mathcal{G}_{2}=(\mathcal{U},\mathcal{V},\mathcal{E}^{\prime})$.
-Then following step 3, we split graph $\mathcal{G}_{2}$ by subject $\mathcal{U}$ with the given splitting ratio and form three disjoint graphs $\mathcal{G}_{train},\mathcal{G}_{val},\mathcal{G}_{test}$.
-In step 4, some edges satisfying zero data leakage condition are not included in the graph.
-We add these edges to corresponding graphs, extending each graph $\mathcal{G}_{train},\mathcal{G}_{val},\mathcal{G}_{test}$ to its maximally scalable state and finishing the dataset splitting process.
-The sample in fMRI dataset is formatted in $F_{ij}=concat(f_{ij},f_{i,j+1},\ldots,f_{i,j+L-1})$ and $T_{j}=concat(s_{j},s_{j+1},\ldots,s_{j+L-1})$.
-If we follow the same process as EEG, text stimuli leakage will occur in the overlapping part of two samples, when one sample is assigned to training set and the other is assigned to validation or test set.
-We propose a simple solution that achieves the balance between abandoning as little data as possible and ensuring zero data leakage.
-Instead of $\left\langle S_{i},T_{j}\right\rangle$ pair, we consider $\left\langle S_{i},M_{k}\right\rangle$ pair and apply the above-mentioned algorithm.
-More details and pseudo-code are available in Appendix B.
-[…]
-$N=1$ $N=2$ $N=3$ $N=4$ $R$ $P$ $F$
+We perform the same filtering steps to “Narratives” dataset as UniCoRN paper (Xi et al., 2023) and the same filtering steps to ZuCo1.0 as EEG2Text paper (Wang and Ji, 2022).
+In BSLR and TSLR calculation, the number of four different seeds are set as $1,2,3,4$ respectively.
+In signal reconstruction task for encoder of UniCoRN, the batch size of EEG and fMRI data is 512 and 320 respectively.
+The learning rate is set as 1e-4 and 1e-3 separately as the author claimed in the original paper.
+In the fair benchmark, for fMRI data, encoder of UniCoRN is trained through 1e-4 learning rate and decaying to 1e-6 finally for 30 training epochs.
+Decoder is trained through 1e-4 learning rate and decaying to 1e-6 finally for 10 training epochs with 90 batch size.
+Sample length $L$ is set as 10 for all experiments related to fMRI.
+For EEG data, EEG2Text model is trained with 1e-6 learning rate for 80 epochs.
+UniCoRN model is trained with the same settings as fMRI data.
+Appendix C Cross-Subject Data Splitting in Practice
+We present the pseudo-code of two dataset splitting methods for EEG and fMRI signal.
+We only consider a bipartite graph $\mathcal{G}_{1}=(\mathcal{U},\mathcal{V},\mathcal{E})$ instead of a $4$-partite graph in real practice.
+For EEG signal, $\mathcal{U}=\{S_{i}\}_{i=1}^{N}$, $\mathcal{V}=\{T_{j}\}_{j=1}^{M}$.
+While for fMRI signal, $\mathcal{U}=\{S_{i}\}_{i=1}^{N}$, $\mathcal{V}=\{M_{k}\}_{k=1}^{K}$.
+$\mathcal{E}$ is the edge between node in $\mathcal{U}$ and node in $\mathcal{V}$.
+$N,M,K$ indicate the total number of subjects, text segments and stories.
+We assert $M>N$ for EEG dataset and $K<N$ for fMRI dataset, so $e=(u,v)\in\mathcal{E}$ exists for every $v\in\mathcal{V}$, as each text segment or story is listened by at least one subject.
+As shown in step 1 of Figure 4, first we pick one edge for each node $v\in\mathcal{V}$ and build a new bipartite graph $\mathcal{G}_{2}=(\mathcal{U},\mathcal{V},\mathcal{E}^{\prime})$.
+Then following step 2, we split graph $\mathcal{G}_{2}$ by subject $\mathcal{U}$ with the given splitting ratio and form three disjoint graphs $\mathcal{G}_{train},\mathcal{G}_{val},\mathcal{G}_{test}$.
+In step 3, we extend each graph $\mathcal{G}_{train},\mathcal{G}_{val},\mathcal{G}_{test}$ by adding edges without data leakage.
+The main difference of splitting methods for EEG and fMRI lies in how $\mathcal{G}_{2}$ is generated.
+We always choose the side with fewer nodes in bipartite graph $\mathcal{G}_{1}$ to generate $\mathcal{G}_{2}$.
+Specifically, in Algorithm 1 where we assert $|\mathcal{U}|<|\mathcal{V}|$, the adjacency matrix is initialized as $M\times N$.
+In Algorithm 2 where $|\mathcal{V}|<|\mathcal{U}|$, the adjacency matrix is initialized as $N\times K$.
+All assertions are based on real cognitive datasets.
+One more thing to notice is that in Line 14 of both pseudo-code, the loop indicates extending training set, validation set, and test set respectively.
+So the names of variable should be alternated in the repeat loop and the displayed part in pseudo-code is a case example of extending training set.
+We write it in this way for simplicity of expression.
+Appendix D Supplementary Proof
+Definition D.1.
+An directed multigraph $\mathcal{G}$ is a type of graph which is permitted to have multiple edges between two vertices.
+When the edges own identity, $\mathcal{G}$ can be written as $\mathcal{G}=(\mathcal{V},\mathcal{E},f)$, where $f:\mathcal{E}\rightarrow\mathcal{V}\times\mathcal{V}$ is an incidence function that maps each edge to a pair of vertices.
+Definition D.2.
+A $k$-partite graph $\mathcal{G}$ is a type of graph that can be divided into $k$ distinct independent sets such that no two vertices in the same set are connected.
+$\mathcal{G}=(\mathcal{V},\mathcal{E})$, where $\mathcal{V}=\mathcal{V}_{1}\cup\mathcal{V}_{2}\cup\cdots\cup\mathcal{V}_{k}$ and $\forall i\neq j,\mathcal{V}_{i}\cap\mathcal{V}_{j}=\emptyset$.
+Notation D.3.
+$\otimes$ is a Cartesian product-like operator.
+$X\otimes Y=\{(x,y)|x\in X,y\in Y,$ there exists relationship between $x$ and $y$ in dataset}\}.
+It’s designed to describe the connectivity among $\mathcal{S},\mathcal{M},\mathcal{T},\mathcal{F}$.
+For example, edges in $\mathcal{S}\otimes\mathcal{M}$ indicates certain subjects are stimulated by certain stories as described in dataset.
+Definition D.4.
+The training set for cross-subject brain-to-text decoding should be formatted in $\mathcal{G}_{train}=\mathcal{S}_{train}\otimes\mathcal{M}\otimes\mathcal{T}_{train}\otimes\mathcal{F}_{train}$, where $\mathcal{S}_{train}=\{S_{i}|\forall S_{i}^{\prime}\in\mathcal{S}_{test},S_{i}\neq S_{i}^{\prime}\}$; $\mathcal{F}_{train}=\{F_{ijk}|i\in I\}$, $I=\{i|\forall j,\forall k,F_{ijk}\notin\mathcal{F}_{test}\}$; $\mathcal{T}_{train}=\{T_{kj}|\forall T_{kj}^{\prime}\in\mathcal{T}_{test},T_{kj}\neq T_{kj}^{\prime}\}$.
+Why a method without brain signal leakage and text stimuli leakage must satisfy cross-subject brain-to-text decoding criterion
+Training set $\mathcal{G}_{train}$ without brain signal leakage and text stimuli leakage is formatted in
+$\mathcal{G}_{train}$ $=\{(S_{i},M_{k},T_{kj},F_{ijk})|$ (21)
+$\forall(S_{i}^{\prime},M_{k}^{\prime},T_{kj}^{\prime},F_{ijk}^{\prime})\in\mathcal{G}_{test},$
+$S_{i}\neq S_{i}^{\prime},T_{kj}\neq T_{kj}^{\prime}\}$
+$=\mathcal{S}_{train}\otimes\mathcal{M}\otimes\mathcal{T}_{train}\otimes\mathcal{F}$
+where $\mathcal{S}_{train}=\{S_{i}|\forall S_{i}^{\prime}\in\mathcal{S}_{test},S_{i}\neq S_{i}^{\prime}\},\mathcal{T}_{train}=\{T_{kj}|\forall T_{kj}^{\prime}\in\mathcal{T}_{test},T_{kj}\neq T_{kj}^{\prime}\}$.
+Since $F_{ijk}\in\mathcal{F}$ indicates brain signal of subject $S_{i}$ stimulated by text segment $T_{kj}$, and given the definition of operator $\otimes$, $\mathcal{F}$ is determined when $\mathcal{S}$ and $\mathcal{T}$ are specified, which is
+$\mathcal{F}$ $=\{F_{ijk}|i\in I,kj\in J\},$ (22)
+$I$ $=\{i|S_{i}\in\mathcal{S}_{train}\},$
+$J$ $=\{kj|T_{kj}\in\mathcal{T}_{train}\}.$
+$\mathcal{F}$ can also be written as $\mathcal{F}=\{F_{ijk}|i\in I\}$, $I=\{i|\forall j,\forall k,F_{ijk}\notin\mathcal{F}_{test}\}$, which is equal to Definition D.4.
+Why the proposed splitting method satisfy zero data leakage
+Take the splitting method for EEG signal as example, the training set and test set after step 1 and step 2 already satisfy
+$\mathcal{G}_{train}=\{(S_{i},M_{k},T_{kj},F_{ijk})|\forall(S_{i}^{\prime},M_{k}^{\prime},$ (23)
+$T_{kj}^{\prime},F_{ijk}^{\prime})\in\mathcal{G}_{test},S_{i}\neq S_{i}^{\prime},T_{kj}\neq T_{kj}^{\prime}\}$
+$\mathcal{G}_{test}=\{(S_{i},M_{k},T_{kj},F_{ijk})|\forall(S_{i}^{\prime},M_{k}^{\prime},$ (24)
+$T_{kj}^{\prime},F_{ijk}^{\prime})\in\mathcal{G}_{train},S_{i}\neq S_{i}^{\prime},T_{kj}\neq T_{kj}^{\prime}\}$
+So we only need to prove expanded graph $\mathcal{G}_{train\_exp}^{\prime}$ and $\mathcal{G}_{test\_exp}^{\prime}$ satisfy zero data leakage, which is obvious from Equation 13 and 14.
+Why we must discard samples to ensure no data leakage
+If $\mathcal{G}_{train}\cup\mathcal{G}_{test}=\mathcal{G}_{\mathcal{D}}$, suppose $\forall(S_{i},M_{k},T_{kj},F_{ijk})\in\mathcal{G}_{train}$, $(S_{i}^{\prime},M_{k}^{\prime},T_{kj}^{\prime},F_{ijk}^{\prime})\in\mathcal{G}_{test},S_{i}\neq S_{i}^{\prime},T_{kj}\neq T_{kj}^{\prime}$.
+For $f(\mathcal{E})=(M_{k},T_{kj})$, $f(\mathcal{E}^{\prime})=(M_{k}^{\prime},T_{kj}^{\prime})$, $T_{kj}\neq T_{kj}^{\prime}$:
+  • If $M_{k}=M_{k}^{\prime}$, then there must exist a subject $S_{i}=S_{i}^{\prime}$ such that he is stimulated by the whole stories.
+  • If $M_{k}\neq M_{k}^{\prime}$, then there must exist a subject $S_{i}=S_{i}^{\prime}$ such that he is stimulated by two different stories.
+As a result, if $\mathcal{G}_{train}\cup\mathcal{G}_{test}=\mathcal{G}_{\mathcal{D}}$, then $\exists(S_{i},M_{k},T_{kj},F_{ijk})\in\mathcal{G}_{train}$, $(S_{i}^{\prime},M_{k}^{\prime},T_{kj}^{\prime},F_{ijk}^{\prime})\in\mathcal{G}_{test}$, $s.t.$ $S_{i}=S_{i}^{\prime}$ or $T_{kj}=T_{kj}^{\prime}$.
+Some samples must be discarded to ensure no data leakage.
+Dataset Model Method With Teacher-Forcing / Without Teacher-Forcing
+BLEU-1 BLEU-2 BLEU-3 ROUGE1-F
+Narratives UniCoRN (a) 49.56 / 33.82 30.49 / 9.44 21.07 / 4.89 40.65 / 21.87
+(b) 26.37 / 19.68 7.50 / 5.50 2.48 / 1.76 19.62 / 16.86
+(c) 50.24 / 29.87 30.83 / 8.35 21.23 / 3.21 41.01 / 19.03
+(d) 49.63 / 30.20 30.29 / 8.78 20.85 / 3.65 41.03 / 19.42
+(e) 28.94 / 21.46 9.39 / 6.24 4.07 / 1.83 19.49 / 17.39
+(f) 22.83 / 16.85 5.69 / 4.24 1.43 / 0.65 19.04 / 15.34
+ZuCo UniCoRN (a) 58.09 / 19.47 49.23 / 7.69 43.23 / 2.97 67.50 / 17.25
+(c) 52.30 / 19.70 42.89 / 7.54 36.80 / 2.93 67.29 / 17.37
+(d) 50.02 / 22.02 43.53 / 8.28 32.71 / 3.15 67.33 / 18.33
+(f) 23.32 / 14.02 7.78 / 2.57 3.01 / 0.82 17.92 / 11.95
+EEG2Text (a) 51.22 / 21.99 33.83 / 7.42 22.99 / 2.94 46.58 / 17.78
+(c) 53.83 / 20.41 38.99 / 7.25 29.57 / 2.48 53.56 / 17.32
+(d) 53.92 / 19.80 41.06 / 7.46 23.12 / 3.06 49.38 / 17.29
+(f) 24.49 / 13.52 7.49 / 2.86 2.28 / 0.78 25.74 / 11.20
+Table 6: Performance of brain-to-text decoding models under different splitting methods with teacher-forcing and without teacher-forcing.
+The green mark and red mark denotes a method without and with text stimuli leakage correspondingly.
+Model Epoch+lr+Method BLEU-N ($\%$) ROUGE-1 ($\%$)
+$N=1$ $N=2$ $N=3$ $N=4$ $F$ $P$ $R$
 UniCoRN 10+1e-3+(a) 49.56 30.49 21.07 15.49 44.83 50.41 40.65
@@
 Here UniCoRN indicates the encoder of UniCoRN is randomly initialized instead of pre-trained through signal reconstruction task.
-4 Experimental Settings
-We test some SOTA brain-to-text decoding models on two popular cognitive datasets Narratives (Nastase et al., 2021) and ZuCo (Hollenstein et al., 2018).
-Dataset details are introduced in Appendix A.
-Comprehensive experiments are conducted to evidence the existence of the following phenomena: (1) Brain signals and text stimuli in test set leak into training set in all current dataset splitting methods.
-(2) The model’s generalization ability, particularly that of the auto-regressive decoder, has been overestimated due to data leakage.
-Because the number of tasks in EEG dataset is too small and method (e) makes no difference to EEG as method (d), we only consider method (a), (c), (d) for EEG.
-4.1 Implementation
-We follow the same settings of UniCoRN (Xi et al., 2023) and EEG2Text (Wang and Ji, 2022), except all the datasets are split to the ratio of 8:1:1 for fair comparison.
-All experiments are conducted on NVIDIA A100-SXM4-40GB GPUs.
-More details are shown in Appendix A.
-[…]
-$N=1$ $N=2$ $N=3$ $N=4$ $R$ $P$ $F$
+Model Epoch+lr+Method BLEU-N ($\%$) ROUGE-1 ($\%$)
+$N=1$ $N=2$ $N=3$ $N=4$ $F$ $P$ $R$
 UniCoRN 50+1e-4+(a) 58.09 49.23 43.23 38.43 63.88 61.12 67.50
@@
 Table 8: Generation quality of UniCoRN and EEG2Text model for EEG under different training settings.
-4.2 Data Leakage Metrics
-We have analyzed two kinds of data leakage: brain signal leakage and text stimuli leakage in section 3. In this part, we will quantify two kinds of data leakage through experiments.
-To better illustrate the extent of data leakage of different data splitting methods, we design two novel evaluation metrics named Brain Signal Leakage Rate (BSLR) and Text Stimuli Leakage Rate (TSLR) for detecting brain signal leakage and text stimuli leakage.
-Note that the situation for validation set is the same as test set, so we only consider test set in experiments.
-BSLR indicates the average percentage of each subject’s brain signals in test set appearing in training set, which could be formulated as
-$\frac{1}{N}\sum_{i=1}^{N}\min(1,\frac{|\{F_{ij}|F_{ij}\in S_{test}\cap S_{train}\}|}{|\{F_{ij}|F_{ij}\in S_{train}\}|})$ (8)
-where $N$ stands for the total number of subjects in test set.
-||$\cdot$|| ⋅ | stands for the cardinality of a set.
-Function $\min(\cdot,\cdot)$ is applied to make sure for each subject the data leakage rate is less than one.
-The definition of TSLR is somewhat different for EEG signal and fMRI signal.
-As to EEG signal where brain signals are sampled continuously, it’s easy to match certain text stimuli with corresponding signals.
-Its TSLR is similar to BSLR, which indicates the average percentage of certain text in test set appearing in training set.
-TSLR for EEG data can be calculated through
-$\frac{1}{M}\sum_{j=1}^{M}\min(1,\frac{|\{T_{ij}|T_{ij}\in S_{test}\cap S_{train}\}|}{|\{T_{ij}|T_{ij}\in S_{train}\}|})$ (9)
-where $M$ stands for the total number of unique text periods in test set and $T_{ij}$ stands for $j$-th period of text stimuli received by $i$-th subject.
-The fMRI signal is sampled discretely with a deterministic interval TR, making it hard to acquire signals corresponding to text.
-Previous methods instead concatenated continuous fMRI frames of certain length with their corresponding text segments as training samples.
-As a result, we consider the average percentage of the same text segments in test set appearing in training set as TSLR for fMRI signal.
-It can be formulated as
-$\frac{1}{M}\sum_{j=1}^{M}\tau\frac{|\{T_{ij}|T_{ij}\in S_{test}\cap S_{train}\}|}{|S_{test}|\times L}$ (10)
-where $\tau=0$ if $\{T_{ij}|T_{ij}\in S_{test}\cap S_{train}\}=\emptyset$ else
-$\tau=\min(1,\frac{|\{T_{ij}|T_{ij}\in S_{train}\}|}{|\{T_{ij}|T_{ij}\in S_{test}\cap S_{train}\}|}).$ (11)
-[…]
-$N=1$ $N=2$ $N=3$ $N=4$ $R$ $P$ $F$
-Narratives UniCoRN 22.83 5.69 1.43 0.48 15.55 24.80 19.04
-ZuCo UniCoRN 23.32 7.78 3.01 1.09 18.47 20.00 17.92
-EEG2Text 24.49 7.49 2.28 0.62 23.98 23.95 25.74
-Table 4: A fair benchmark for evaluating brain-to-text decoding.
-5 Results and Analysis
-5.1 Verification for Data Leakage
-We test current data splitting methods and our method on fMRI dataset “Narratives” and EEG dataset ZuCo.
-Considering the influence of randomness in splitting, we randomly select four seeds for experiments.
-The results are shown in Table 1, and are consistent with theoretical analysis.
-For fMRI, current methods apart from method (a) suffer from brain signal leakage, while method (a) has serious text stimuli leakage.
-Method (b) gets no text stimuli leakage but has slight brain signal leakage.
-The situation for EEG is similar to that of fMRI.
-Apart from our proposed method (f), there is no way to achieve zero brain signal leakage and text stimuli leakage at the same time.
-5.2 Damage of Data Leakage
-Brain signal leakage and text stimuli leakage will damage brain-to-text decoding models from both the encoder side and decoder side.
-Effect on Encoder
-The encoder of current brain-to-text decoding models can be trained in two ways: either jointly with the decoder or solely pre-trained through a reconstruction task.
-In the former end-to-end training scenario, it is hard to evaluate encoder performance separately.
-So we mainly focus on the latter, in which case the encoder is pre-trained through an encoder-decoder framework to reconstruct input brain signals.
-The decoder here does not refer to the decoder for text generation.
-It is similar to the structure of the encoder and will be abandoned once the encoder is pre-trained.
-Since a proper evaluation index of the encoder’s representation ability is missing, validation loss is applied to measure the effect of data leakage.
-We test different splitting methods on two cognitive datasets.
-The validation loss of encoder is shown in Figure 4.
-For fMRI, influenced by leakage of brain signals, the validation loss of method (b), (c), (d), (e) keeps dropping even with long training epochs.
-The encoder is actually overfitting and degrading.
-For method (a) and (f) without brain signal leakage, the validation loss quickly rises after reaching the lowest point with a few epochs, satisfying the basic rule of machine learning.
-For EEG, we find validation loss keeps dropping for all methods even with very long training epochs, regardless of brain signal leakage or not.
-We think the poor spatial resolution of EEG signal might lead to this phenomenon.
-Figure 4: Validation loss of encoder under different dataset splitting methods in two datasets.
-Effect on Decoder
-All SOTA models choose the pre-trained language model BART (Lewis et al., 2020) as decoder.
-The powerful model is able to achieve fluent open vocabulary text generation.
-However, if data leakage occurs, due to the auto-regressive generation that calculates the probability of current token based on all previous tokens, the decoder will generate seen text given the first few words, and fail to generalize to unseen text.
-The influence of text stimuli leakage on decoder is detected through BLEU scores (Papineni et al., 2002) and ROUGE-1 scores (Lin, 2004), which measure text similarity between generated text and ground truth.
+Appendix E Supplementary Experiment
+The influence of teacher-forcing
+Jo et al.
+(2024) pointed out that previous methods applied teacher-forcing during generation, which has the tendency to cause overfitting problems.
+Therefore, we conduct experiments on models without teacher-forcing to ensure that the affection of data leakage is not influenced by teacher-forcing.
+Results are shown in Table 6.
+Without the influence of teacher-forcing decoding, splitting methods with data leakage will still lead to overestimation of model performance.
+Experiments on longer training epochs with smaller learning rate
 If evaluation indicators keep improving as training epochs increase, we believe part of the test set is leaked into training set and the model is overfitting.
@@
 As shown in Table 7, we test two kinds of UniCoRN models.
-One is UniCoRN with finely tuned hyper-parameters claimed in the original paper, and the other is UniCoRN with a randomly initialized encoder.
-Empirically, the former will perform much better than the latter.
-However, in method (a), (c), (d), due to text stimuli leakage, if we reduce the learning rate and extend training epochs, UniCoRN performs much better than UniCoRN and its performance keeps rising with longer training epochs.
+One is UniCoRN with hyper-parameters claimed in the original paper, and the other is UniCoRN whose encoder is randomly initialized.
+Besides, UniCoRN is trained with longer epochs and smaller learning rate.
+In method (a), (c), (d), due to text stimuli leakage, if we reduce the learning rate and extend training epochs, UniCoRN performs much better than UniCoRN and its performance keeps rising with longer training epochs.
 As to method (b) and (e) with no text stimuli leakage, changing training epochs or learning rates makes no obvious difference to model performance.
@@
 For method (d) without text stimuli leakage, both models reach optimal performance after the first few rounds of training epochs.
-5.3 A Fair Benchmark
-We re-evaluate current SOTA models for brain-to-text decoding under our cross-subject data splitting method and release a fair benchmark.
-UniCoRN is tested for both fMRI and EEG decoding, EEG2Text model is tested for EEG decoding.
-The results are listed in Table 4.
-For EEG dataset, UniCoRN achieves higher results in BLEU-2,3,4 while EEG2Text is better in BLEU-1 and ROUGE-1.
-6 Conclusion
-In this paper, we conduct a comprehensive study on existing cross-subject data splitting methods, and evidence that all these methods suffer from data leakage problem.
-Such data leakage largely exaggerates model performance and leads to poor generalization.
-To fix this issue, we propose a cross-subject data splitting criterion for brain-to-text decoding, aiming to improve the utilization efficiency of cognitive dataset and the generalization ability of decoding models.
-Experiments are conducted on fMRI and EEG dataset respectively.
-Current SOTA models are re-evaluated under this proposed splitting method and a fair benchmark is released for further research in this domain.
-Limitations
-The “Narratives” dataset and the ZuCo dataset provide researchers with precise brain signal resources stimulated by text or voice.
-However, in the brain-to-text decoding task, both subject’s brain signals and text stimuli in the validation and test set need to be invisible to the training set, which makes splitting these public datasets difficult.
-Our proposed dataset splitting method meets the above requirements at the expense of discarding some data in the dataset.
-We recommend future datasets in this domain follow these guidelines.
-The division of the training set, validation set, and test set should be provided when the dataset is released.
-Besides, we suggest hiring new subjects with unique stimuli for the validation set and test set, which is good for testing the generalization ability of models without loss of data.
-What’s more, we find existing models rely more on a strong auto-regressive decoder to achieve good generation quality.
-The encoder is of limited use in all SOTA models.
-And we also notice in experiments that the encoder of EEG2Text keeps overfitting whether with or without brain signal leakage.
-We leave it to research in the future.
-Ethics Statement
-In this paper, we introduce a new dataset splitting method to avoid data leakage for decoding brain signals to text task.
-Experiments are conducted on the publicly accessible cognitive datasets “Narratives” and ZuCo1.0 with the authorization from their respective maintainers.
-Both datasets have been de-identified by dataset providers and used for researches only.
-Appendix A Implementation Details
-[…]
-More details in experiments are supplemented in this section.
-We perform the same filtering steps to “Narratives” dataset as UniCoRN paper (Xi et al., 2023) and the same filtering steps to ZuCo1.0 as EEG2Text paper (Wang and Ji, 2022).
-In BSLR and TSLR calculation, the number of four different seeds are set as $1,2,3,4$ respectively.
-In signal reconstruction task for encoder of UniCoRN, the batch size of EEG and fMRI data is 512 and 320 respectively.
-The learning rate is set as 1e-4 and 1e-3 separately as the author claimed in the original paper.
-In the fair benchmark, for fMRI data, encoder of UniCoRN is trained through 1e-4 learning rate and decaying to 1e-6 finally for 30 training epochs.
-Decoder is trained through 1e-4 learning rate and decaying to 1e-6 finally for 10 training epochs with 90 batch size.
-Sample length $L$ is set as 10 for all experiments related to fMRI.
-For EEG data, EEG2Text model is trained with 1e-6 learning rate for 80 epochs.
-UniCoRN model is trained with the same settings as fMRI data.
-Appendix B Details of Cross-Subject Splitting
-For fMRI dataset, we consider a bipartite graph $\mathcal{G}_{1}=(\mathcal{U},\mathcal{V},\mathcal{E})$ where $\mathcal{U}=\{S_{i}\}_{i=1}^{N}$, $\mathcal{V}=\{M_{k}\}_{k=1}^{M}$.
-$\mathcal{E}$ is the edge between node in $\mathcal{U}$ and node in $\mathcal{V}$, indicating $\left\langle S_{i},M_{k}\right\rangle$ pair in the dataset.
-$N$ is the total number of subjects and $M$ is the total number of tasks.
-We assert $M<N$, so $e=(u,v)\in\mathcal{E}$ exists for every $v\in\mathcal{V}$, as each text stimuli is listened or read by at least one subject.
-As shown in step 2 of Figure 3, first we pick one edge for each node $v\in\mathcal{V}$ and build a new bipartite graph $\mathcal{G}_{2}=(\mathcal{U},\mathcal{V},\mathcal{E}^{\prime})$.
-Then following step 3, we split graph $\mathcal{G}_{2}$ by subject $\mathcal{U}$ with the given splitting ratio and form three disjoint graphs $\mathcal{G}_{train},\mathcal{G}_{val},\mathcal{G}_{test}$.
-In step 4, some edges satisfying zero data leakage condition are not included in the graph.
-We add these edges to corresponding graphs, extending each graph $\mathcal{G}_{train},\mathcal{G}_{val},\mathcal{G}_{test}$ to its maximally scalable state and finishing the dataset splitting process.
-We also release the pseudo-code of two dataset splitting methods for EEG and fMRI signal.
-As shown in Figure 3, our proposed dataset splitting method consists of four steps.
-The blue lines stand for the situation of original dataset.
-The main difference between two methods lies in the how $\mathcal{G}_{2}$ is generated.
-We always choose the side with fewer nodes in bipartite graph $\mathcal{G}_{1}$ to perform $\mathcal{G}_{2}$ generation.
-For example, in Algorithm 1 where we assert $|\mathcal{U}|<|\mathcal{V}|$, the adjacency matrix is initialized as $M\times N$.
-In Algorithm 2 where $|\mathcal{V}|<|\mathcal{U}|$, the adjacency matrix is initialized as $N\times K$.
-All assertions are based on analysis of cognitive datasets.
-One more thing to notice is that in Line 14 of both pseudo-code, the loop indicates extending training set, validation set, and test set respectively.
-So the names of variable should be alternated in the repeat loop and the displayed part in pseudo-cod is a case example of extending training set.
-We write it in this way for simplicity of expression.
 1 Initialize: Bipartite graph $\mathcal{G}_{1}=(\mathcal{U},\mathcal{V},\mathcal{E})$, $\mathcal{G}_{2}=(\mathcal{U},\mathcal{V},\mathcal{E}^{\prime})$ where $\mathcal{U}=\{S_{i}\}_{i=1}^{N}$ and $\mathcal{V}=\{T_{j}\}_{j=1}^{M}$, Adjacency matrix $A_{1}$ of $\mathcal{G}_{1}$ where $A_{1}[i][j]=1$ if node $i$ and node $j$ is connected else $A_{1}[i][j]=0$, Adjacency matrix $A_{2}$ of $\mathcal{G}_{2}$ where $A_{2}[i][j]=0$, Array $C$ where $len(C)=len(\mathcal{U})$ and $C[i]=0$;
@@
 7
-8      Minimum $=\min(C_{copy})$;
+8 Minimum $\min(C_{copy})$;
 […]
-[…]
+10 $C[.index]\leftarrow C[Minimum$.index]+1.index]+1; // Make degree of nodes balanced
 11Split by subjects $\mathcal{U}$ according to default ratio;
@@
 29
-30      Minimum $=\min(C_{copy})$;
+30 Minimum $\min(C_{copy})$;
 […]
-[…]
+32 $C[.index]\leftarrow C[Minimum$.index]+1.index]+1; // Make degree of nodes balanced
 33Split by tasks $\mathcal{V}$ according to default ratio;